### v1

# Threshold Graphs are Globally Synchronizing

Abstract.
Place $n$ nodes on the unit circle and connect some pairs with springs that pull connected nodes toward each other. Which graphs always drive all nodes to converge to a single point, no matter where they start? This is known as the global synchronization problem of the Kuramoto model and is closely tied to the benign nonconvexity of its synchronization landscape. In this work, we show how this dynamical and continuous question can be reduced to a static and combinatorial one. We prove that threshold graphs, the graph class closed under adding isolated nodes and adding universal nodes, are globally synchronizing, based purely on a combinatorial analysis of unit-vector configurations in the plane. It is particularly striking that threshold graphs range from extremely sparse structures such as stars to the complete graph, suggesting that global synchronization in this class cannot be attributed to density. Our proof shows that global synchronization can instead emerge from a form of geometric symmetry.
1. Introduction
The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks. Since then, synchronization phenomena have been widely studied across both natural and technological systems, from the flashing of fireflies [BB68, SS93] and the rhythmic beating of heart cells [Win67, MS06] to synchronization in power grids [DB13], and the Transformer architecture [GLPR23a, CRMB24]. The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kur75], strikes a balance between mathematical tractability and sufficient complexity, has become one of the key mathematical models for understanding the collective dynamics of coupled oscillators.
The Kuramoto model describes a system of $n$ oscillators continuously evolving on the unit circle. Each vertex $i\in V$ corresponds to an oscillator with phase (angle) $\theta_{i}:\mathbb{R}\rightarrow\mathbb{S}^{1}$. The dynamics of phase $\theta_{i}$ of each oscillator evolves according to:
$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),$ (1)
This means that each oscillator’s phase dynamics is governed by its intrinsic frequency $\omega_{i}$ while also being influenced by the phase differences with its neighbors in a nonlinear manner. There are two lines of research regarding this model focusing heterogenous and homogenouse case respectively.
Studies on the former case primarily consider a systems of oscillators with heterogenous intrinsic frequences $\{\omega_{i}\}_{i=1}^{N}$ coupled through a complete graph with identical weight $\boldsymbol{A}_{ij}=\frac{K}{N}$. Here, $K$ determines how strongly oscillators influence each other. A fundamental question in the heterogeneous setting is to determine the critical coupling strength $K_{c}$ such that, for $K>K_{c}$, the oscillators achieve frequency synchronization, i.e., $\dot{\theta}_{i}-\dot{\theta}_{j}\to 0$, meaning that all oscillators rotate with the same angular velocity. If this indeed holds, it is natural to further ask whether the oscillators not only rotate with the same angular velocity, but also converge to the same phase, known as phase synchronization, that is, whether $\theta_{i}-\theta_{j}\rightarrow 0$, as $t\rightarrow\infty$. The study of synchronization in the heterogeneous Kuramoto model is particularly relevant to real-world systems. In many natural and engineered networks such as power grids, neuronal circuits, and circadian rhythms, the individual oscillators typically have distinct natural frequencies. Within this heterogeneous setting, a rich variety of nonlinear dynamical phenomena can emerge, including partial synchronization, multistability, and even chaotic behavior. For further background and details, we refer the reader to [Str00, ABV+05, DB14].
Recent works on the Kuramoto model has moved beyond the fully connected case to more general interaction graphs. Such settings are not only mathematically appealing but also more realistic, since interactions in natural and engineered systems are rarely all-to-all. In this work, we focus on the homogeneous case with identical intrinsic frequencies ($\omega_{i}=\omega$ for all $i$) and aim to understand how graph structure governs synchronization. To begin with, we first remark that in this case, frequency synchronization always holds and the remaining question is whether phasr synchronization can be achieved. This is not obvious if we only look at the equation $\dot{\theta}_{i}=\omega+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})$, for all $i$, since it tells that the instantaneous angular velocity of each node depends both on this common intrinsic term and on the coupling with its neighbors. However, applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, the equations becomes
$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),\quad i\in\{1,\cdots,n\},$ (2)
and this reduced system is in fact a gradient flow of the Kuramoto energy function
$E(\theta)=\frac{1}{2}\sum_{1\leq i<j\leq n}\boldsymbol{A}_{ij}(1-\cos(\theta_{i}-\theta_{j}))$ (3)
as its gradient satisfies $\dot{\theta_{i}}{dt}=-\frac{\partial E}{\partial\theta_{i}}$. By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$. Hence, the angle differences do not lead to frequency discrepancies in the asymptotic sense. In other words, frequency synchronization always holds, as they share common frequency $\omega$. Now the real challenge in the homogeneous case lies in understanding whether phase synchronization can always be achieved, regardless of the initial phase configurations. It is evident that this property is entirely determined by the graph structure, since the adjacency matrix $\mathbf{A}$ of the graph is the only external parameter in the homogeneous Kuramoto model (2) and equivalently in its energy formulation (3). Thus, we define:
Definition 1.1 (Globally synchronizing graph).
We say that a graph $G$ with adjacency matrix $\mathbf{A}$ is globally synchronizing if, all oscillators converge to the same phase, i.e.,
$\lim_{t\to\infty}\theta_{i}(t)-\theta_{j}(t)=0\quad\text{for all }i,j\in V.$
for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero. In other words, almost all trajectory converges to the synchronous state $\theta_{i}=\theta_{j}$, for all $i,j\in V(G)$.
The key question is now stated as:
On which graphs, the Kuramoto model (2) exhibits global synchronization?
This seemingly simple description conceals a mathematically challenging task. To address this problem, we note that the Kuramoto model (2) can be formulated as a gradient flow of the energy function (3), as mentioned before. Hence, understanding the geometry of (3) provides direct insight into the long-term dynamics of (2). In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous equilibria. If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing. In contrast, if the energy function admits no spurious local minima, then the graph $G$ is globally synchronizing, since this rules out convergence to any non-synchronous equilibrium except possibly for a set of initial conditions of measure zero. A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality together with the center manifold theorem; see Lemma A.1 in [GLPR23b] for an excellent exposition.
Contributions and prior work. Aiming to understand which graphs and how them foster global synchronization, previous studies on this problem can be broadly divided into two lines of research: (i) investigating whether random graphs globally synchronize, and (ii) determining the minimum edge density required to ensure global synchronization. Both approaches are, in a certain sense, grounded in the fact that the complete graph is globally synchronizing, and that both random graphs and very dense graphs tend to behave like complete graphs. Intuitively, a graph of sufficiently high density becomes homogeneous enough to approximate a complete graph. In this sense, full connectivity is not strictly required, that is, global synchronization can still emerge with a small gap from the fully connected case.
In this work, we prove that threshold graphs, as a graph class that by no means closed to complete, but still globally synchronize.
(i) What are threshold graphs? Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices. This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one. Thus our goal is to show all threshold graphs with sequence ending by $1$ must globally synchronizes, because otherwise it is not connected.
Figure 1. Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$). Node labels show the construction order: 11 is the initial vertex, later numbers follow the binary code. White and black circles denote isolated and dominating vertices (bit 0 and bit 11). Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
(ii) Why threshold graphs? Many theorems start from initial observations, from which suitable conjectures are formulated. Our case is no exception. Our investigation into global synchronization on threshold graphs began with a simple observation: complete graphs and star graphs, respectively the densest and sparsest connected graphs in the class of threshold graphs, both exhibit global synchronization. Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize. Indeed, although the two extremes synchronize, this does not necessarily imply that the entire spectrum of threshold graphs does.
However, two further observations strengthened our confidence in this conjecture. Firstly, in threshold graphs, the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N[u]\subseteq N(v)$ or $N[v]\subseteq N(u)$. This total preorder of vicinities suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization. Secondly, no counter-example was found in our numerical experiments up to 50 size graphs [^1] First, we transform the equilibrium equations of the Kuramoto model into a system of polynomial equations. We then use Bertini software package [BHSW13] to compute the zeros of this system, and subsequently evaluate the Jacobian at each solution to determine their stability.. These evidence encourages us to investigate whether all connected threshold graphs global synchronize. In particular, we established the following result:
Theorem 1.2.
Any connected threshold graph globally synchronizes.
Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma. This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony. Based on this observation, the bootstrap strategy attempts to show that all second-order critical points must lie within a half-circle. If successful, this rules out the existence of non-synchronous equilibria and thereby implies global synchronization.
However, for threshold graphs, it remains unclear how to apply this strategy. It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment. In contrast, threshold graphs span the full spectrum from extremely sparse examples such as stars to dense cliques, and their highly inhomogeneous, layered structure complicates the application of this strategy. As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometry condition. This condition naturally leads us to study structural twins (pairs of nodes sharing the same neighborhood in the graph) as well as a new notion we introduce, that of geometric twins (a broader concept depending jointly on the graph structure and the equilibrium configuration $\theta$). These notions clarify how the relative positions of phasors behave at second order stationary point of the energy landscape $\mathbb{E}(\boldsymbol{\theta)}$. Building on this understanding, we employ an inductive strategy: synchronization is first established locally on small substructures, and then shown to propagate across the whole graph along the neighborhood-inclusion hierarchy of threshold graphs.
(iii) What we learned from global synchrony of threshold graphs? While identifying previously unknown classes of globally synchronizing graphs contributes to the expansion of known results, a more fundamental objective of this line of research is to elucidate the mechanisms that give rise to global synchronization itself. It is already known that for certain classes of graphs, global synchronization arises through mechanisms that, in a certain sense, are close to those of complete graphs. Threshold graphs, however, lie outside this regime. Nevertheless, we prove in this work that they also exhibit global synchronization. What deeper insight can we draw from this result?
Our analysis shows that the synchronization of threshold graphs is instead driven by another mechanism, namely their intrinsic symmetries. These symmetries are not merely combinatorial, such as sets of nodes sharing identical neighborhoods, but are geometric in nature. We will introduce this notion in detail later, but to give an intuition, these geometric symmetries emerge inductively as the graph is reduced node by node. Through this process, we uncover certain sets of nodes that, while not symmetric in the structural sense of graph theory, exhibit geometric symmetry at the second-order critical points of (3), which allows us to show that these nodes form a synchronized group. As the reduction continues, more geometric symmetries are revealed, leading to the formation of progressively larger synchronized groups, until the entire node set forms a single synchronized group, thereby guaranteeing that all second order critical points are synchronized states, thus the graph is global synchronizating.
1.1. Notation
Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges. Furthermore, since the homogeneous Kuramoto model (2) on $G$ is fully determined by the graph topology, we sometimes use system $G$ to denote the Kuramoto model on this graph for simplicity. Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$.
2. Preliminaries of threshold graphs
Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines. They were in fact discovered independently and reported in different journals by researchers working in distinct fields. The term “threshold graph” was introduced by Chvátal and Hammer in their study of set-packing problems. Around the same time, the same class of graphs was independently discovered by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs. Since then, threshold graphs have appeared in various contexts, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences.
Many equivalent conditions have been found for threshold graphs including constructive forms and forbidden configurations. See for example [MP95] for nine different characterizations.
Definition 2.1.
A graph is threshold if it contains no induced $P_{4}$, $C_{4}$ and $2K_{2}$.
One of the many characterizations of threshold graphs is constructive. The construction is in terms of two basic operations; adding a disjoint vertex (adjacent to no existing vertex), or a dominating vertex (adjacent to all existing vertices).
The construction process can be encoded into a binary sequence, called threshold sequence, that is $(b_{1},\dots,b_{n})\in\{0,1\}^{n}$ such that starting from a single vertex, for each $b_{i}=0$ we add an isolated vertex, and for each $b_{i}=1$ we add a dominating vertex; see Figure 1.
11223344
(a) $P_{4}$
11223344
(b) $C_{4}$
11223344
(c) $2K_{2}$
Figure 2. Forbidden induced subgraphs for threshold graphs: graphs are $\{P_{4},C_{4},2K_{2}\}$-free. Only edges within each induced pattern are drawn.
Remark 2.2.
Interestingly, threshold graphs have also been studied in synchronization contexts beyond oscillator models. In particular, [PS95] surveyed their use in the design of synchronization protocols for cooperating processes, focusing on mutual exclusion via semaphore-based control mechanisms and the so-called PV-chunk semaphores. In some papers, these are also referred to as Pvc-definable graphs.
3. Preliminaries on Kuramoto synchronization
In this section, we aim to establish the correspondence between dynamical and optimization concepts in the homogeneous Kuramoto model. Since the homogeneous Kuramoto model (2) can be formulated as a gradient dynamical system with the energy function (3) and the real analycity of (3), the dynamical notions of equilibrium points and stable equilibria naturally correspond to the optimization notions of first-order and second-order stationary points of the energy function. Before showing this correspondence, we first provide the formal definitions of (stable) equilibria in dynamical systems and of first- and second-order stationary points in optimization. It is worth noting that such a one-to-one correspondence between the two sets of concepts is rather uncommon. This special structure is precisely what allows the analysis of global synchronization in the Kuramoto model to be directly connected to the geometry of the nonconvex optimization landscape.
In general, the evolution of a dynamical system can often be viewed as a descent process of a potential, known as a Lyapunov function, which monotonically decreases along the system trajectories. However, Lyapunov functions are not unique in general, let alone the existence of a clear correspondence between the two classes of concepts discussed above. Moreover, even if the system is of gradient type so that its energy function serves as a natural candidate for a Lyapunov function, such a correspondence is still not guaranteed in general. The remarkable property of the homogeneous Kuramoto model considered here is that its energy function is real analytic, and it is precisely this property that ensures the correspondence between dynamical and optimization notions. For further discussion and proofs, we refer the reader to [AK06].
3.1. Equilibrium and local stability
In this section, we introduce the notion of equilibrium and stable equilibrium of the Kuramoto model (2).
By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02]. Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
Definition 3.1 (Equilibrium of the Kuramoto model (2)).
We call $\boldsymbol{\theta}\in\mathbb{R}^{n}$ an equilibrium of the Kuramoto model (2) if it satisfies
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (4)
An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
Proposition 3.2 (Characterization of stable equilibrium).
An equilibrium point $\boldsymbol{\theta}$ is stable if the Jacobian matrix $\boldsymbol{\mathbf{J}}$, defined as:
$\boldsymbol{\mathbf{J}}_{ij}=\begin{cases}-\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i=j,\\[8.0pt] \boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i\neq j,\end{cases}$
is negative semi-definite.
3.2. Stationary point and local optimality
Definition 3.3 (Stationary point of the energy function (3)).
We call $\theta\in\mathbb{R}^{n}$ a stationary point of the energy function (3) if it satisfies $\nabla E(\theta)=0$, namely
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (5)
Obviously, from Proposition 3.1 and Definition 3.3, a configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a stationary point of the energy function (3) if and only if it is an equilibrium of the Kuramoto model (2).
Definition 3.4 (Local minimum of the energy function).
A configuration $\theta\in\mathbb{R}^{n}$ is called a local minimum of the energy function $E$ if there exists $\varepsilon>0$ such that
$E(\theta)\leq E(\theta^{\prime}),\quad\forall\theta^{\prime}\in\mathbb{R}^{n}\ \text{with}\ \|\theta^{\prime}-\theta\|<\varepsilon.$
Proposition 3.5 (First- and second-order necessary conditions).
If a configuration $\theta$ is a local minimum of $E$, then
  1. (1)
    $\nabla E(\theta)=0$ (first-order condition), and
  2. (2)
    the Hessian $\nabla^{2}E(\theta)$ is positive semidefinite.
A point satisfying the above two conditions is also called a second-order stationary point (SOSP) of (3).
3.3. Correspondence between stability and geometry
In this section, we briefly outline the relationship between equilibrium stability in the dynamical system (2) and the geometric structure of the energy function (3). Understanding this correspondence provides valuable intuition. In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order critical point.
At the first-order level, the two notions coincide: a configuration $\boldsymbol{\theta}$ is a critical point of the energy function $E(\boldsymbol{\theta})$ if and only if it is an equilibrium of the dynamical system (2), since both are characterized by
$\sum_{j=1}^{n}A_{ij}\sin(\theta_{i}-\theta_{j})=0\text{ for all }i\in[n]$
One can further relate the stability of an equilibrium to second-order conditions of the energy landscape. It is precisely because of the real analycity of the energy function (3), the local asymptotic stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (3). This is a remarkable property since it does not hold in general for dynamical systems. In general, the dynamics of a system may admit more complex behaviors near equilibria such as limit cycles, or chaotic dynamics, etc. The relation mentioned here is summarized in Figure 3. One can find the proof in the celebrated work [AK06].
SOSP of 3 Local minima of 3 = stable equilibria of 2
Figure 3. Relation between different concepts. Here equivalence between stable equilibrium and local minimum is established in [AK06].
3.4. Synchronous and aligned states
In this section, we prove that all aligned states are equilibria. Among them, however, the non-synchronous ones are not second-order stationary points (SOSPs). Before that, we first define the notions of synchronous states and aligned states.
Definition 3.6.
A state $(\theta_{1},\cdots,\theta_{n})$ of $G=(V,E)$ is called a synchronous state if for all $i,j\in V$,
$\theta_{i}\equiv\theta_{j}\pmod{2\pi}$.
More generally, it is called an aligned state if for all $i,j\in V$,
$\theta_{i}\equiv\theta_{j}\pmod{\pi}$.
Clearly, every synchronous state is also an aligned state. The stability of all aligned states is:
Theorem 3.7.
Any aligned state is an equilibrium of the Kuramoto model (2) on the graph $G$ or equivalently stationary point of $E(\boldsymbol{\theta})$. Among them, synchronous states are second-order critical point of (3) [^2] In fact, they are asymptotically stable equilibria up to rotational invariance, while non-synchronous aligned states are not second-order critical point of (3).
Proof.
Firstly, it is straightforward to see that aligned states satisfy the condition (4), therefore they are equilibria of $G$, or equivalently stationary point of $E(\boldsymbol{\theta}$.
Secondly, we turn to the second-order criticality of synchronous states. The Hessian matrix at synchronous states is given by
$\boldsymbol{\mathbf{H}}_{ij}=\begin{cases}\sum_{j=1}^{n}\boldsymbol{A}_{ij},&\text{if }i=j,\\[8.0pt] -\boldsymbol{A}_{ij},&\text{if }i\neq j,\end{cases}$ (6)
[…]
Thirdly, regarding the second-order criticality of non-synchronous aligned states. Without loss of generality [^3] Any aligned state can be rotated uniformly on all angles $\theta_{1},\cdots,\theta_{n}$, such that $\theta_{i}=0$ or $\pi$, for any $i\in[n]$., assume $\theta_{i}=0$ or $\pi$, for any $i\in[n]$. Let $\boldsymbol{x}=(x_{1},\cdots,x_{n})$ be an $n$-dimensional vector that indicates which group (either at angle 0 or $\pi$) each node belongs to
[…]
Then the Hessian matrix at any antipodal state is not negative semidefiniteness. Indeed, this can be inferred by the quadratic form
$\boldsymbol{x}^{T}\boldsymbol{\mathbf{H}}\boldsymbol{x}=\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=-\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}<0$
since $\boldsymbol{A}_{ij}\geq 0$, and there must exist at least one pair of nodes $i,j$ such that $\theta_{x}\equiv\theta_{y}\pmod{\pi}$, $k\in\mathbb{Z}$, given the connectness of $G$. Therefore, any non-synchronous aligned state is not a SOSP. ∎
Remark 3.8.
In fact, synchronous states are asymptotically stable equilibria up to a global rotation. Here, asymptotic stability is a stronger notion than mere stability: it requires that small perturbations lead to trajectories returning to the equilibrium. On the other hand, non-synchronous aligned states are not SOSPs and therefore cannot be local minima of 3 or say not stable equilibrium of 2.
The instability of non-synchronous states can be intuitively understood as follows: a set of antipodally positioned oscillators must evolve in a way that decreases the system’s energy, as there always exists an interaction force between these antipodal pairs that drives them closer together.
$A_{12}1122$
Figure 4. Synchronous (left) and non-synchronous aligned state (right). Arrows indicate the nodes moving closer along the circle, implying this state is not stable.
4. Combinatorial planar geometry of equilibrium
As we will introduce later, the key idea of our proof is to assume that a configuration $\boldsymbol{\theta}$ is a second-order critical point and show that this assumption forces $\boldsymbol{\theta}$ to be synchronized, meaning that all node vectors coincide. This requires us to understand the geometric configuration of second-order critical points on the unit circle. In the previous section, we presented the algebraic conditions characterizing first- and second-order critical points.
In this section, we employ a complex-number formulation to transform these algebraic conditions into geometric ones, i.e., combinatorial relations between the vectors associated with the nodes. This seemingly simple transformation significantly simplifies the original landscape geometry problem, reducing it to a combinatorial-geometric problem involving vectors on the two-dimensional plane. With this, one just needs to understand whether, under the geometric form of the first- and second-order conditions, it can be deduced that all vectors must be identical.
4.1. Vector-labelled graph
Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$, we associate to each node $i\in V$ its phasor
$\boldsymbol{v}_{i}:=(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}.$
This yields a vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ which encodes the phase configuration $\boldsymbol{\theta}$ geometrically.
$v_{i}$$\theta_{i}$
Figure 5. Representation of $v_{i}$ with angle $\theta_{i}$.
Although straightforward to construct, this object captures both the equilibrium condition and its stability in the Kuramoto model (2) on the graph $G$.
More precisely, to determine whether a given $\boldsymbol{\theta}$ is an equilibrium and whether it is stable (or more general, second-order critical point), the combinatorial geometry of the vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ provides a geometric view to algebraic one (solving nonlinear equations and analyzing Hessian at stationary point), as detailed in Section 4.2.
4.2. Alignment and equilibrium
Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms. For each node $i\in V$, we aggregate the phasors of its neighbors into a single vector:
$\sum_{j\in N(i)}\boldsymbol{v}_{j}.$
This vector represents the aggregated state of node $i$’s neighbors. If it is not aligned with $\boldsymbol{v}_{i}$, then intuitively, due to the intrinsic synchronization mechanism in the Kuramoto model, the nodes will tend to move toward alignment instead of staying there at rest, therefore not in equilibrium. The equivalence between the alignment $\sum_{j\in N(i)}\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}$ for every node $i\in V$ and the equilibrium condition can be rigorously established by representing the phasors as complex numbers.
Lemma 4.1 (Equilibrium condition rephrased geometrically).
A state $\boldsymbol{\theta}$ is an equilibrium of the Kuramoto model (2) on $G=(V,E)$ with adjacency $\boldsymbol{A}$ if and only if, for each $i\in V$,
$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}\quad\text{where}\quad\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}).$ (8)
Proof of Lemma 4.1.
Let us firstly derive an equivalent form of (8). Rewrite all vectors $\boldsymbol{v}_{i}$ in (8) in terms of their Cartesian components $(\cos\theta_{i},\sin\theta_{i})$, yielding
$\sum_{j\in N(i)}(\cos\theta_{j},\sin\theta_{j})=\mu_{i}(\cos\theta_{i},\sin\theta_{i}).$
Equivalently, this can be written as
$\left\{\begin{aligned} \sum_{j\in N(i)}\cos\theta_{j}&=\mu_{i}\cos\theta_{i},\\ \sum_{j\in N(i)}\sin\theta_{j}&=\mu_{i}\sin\theta_{i}.\end{aligned}\right$. (9)
This is equivalent to verifying the equality between two complex numbers:
$\sum_{j\in N(i)}\cos\theta_{j}+I\sum_{j\in N(i)}\sin\theta_{j}=\mu_{i}\cos\theta_{i}+I\mu_{i}\sin\theta_{i},$
which can be compactly rewritten as
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}.$ (10)
using the Euler formula $e^{I\theta_{i}}=\cos\theta_{i}+I\sin\theta_{i}$. Now that we have derived the equivalence between conditions (8) and (10), our goal now reduces to proving the equivalence between (10) and (3), namely
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}\Longleftrightarrow\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0.$
This indeed holds.
To prove the implication from the right-hand side to the left-hand side, we construct such a $\mu_{i}$ that satisfies the left-hand side. Dividing both sides of (10) by $e^{I\theta_{i}}$ (noting that $e^{I\theta_{i}}\neq 0$), we obtain
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\mu_{i}.$
Such a $\mu_{i}$ exists and is in fact given by
$\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}),$
at an equilibrium point $\boldsymbol{\theta}$, as we will show below. Expanding the fraction, we obtain
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$ $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$ (11)
$=\sum_{j\in N(i)}(\cos(\theta_{j}-\theta_{i})+I\sin(\theta_{j}-\theta_{i})).$
Since $\boldsymbol{\theta}$ is an equilibrium, the imaginary part must vanish, leading to
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
Thus, we conclude
$\sum_{j\in N(i)}e^{I\theta_{j}}=\sum_{j\in N(i)}{\cos(\theta_{j}-\theta_{i})}\cdot e^{I\theta_{i}}.$
Conversely, to prove the left-hand side implies the right-hand side, we show that if there exists a constant $\mu_{i}$ such that
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}},$
then $\boldsymbol{\theta}$ is an equilibrium point. This is equivalent to showing that
$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
Indeed, we compute
$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$ (12)
$=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$
$=\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$
$=\mu_{i}.$
Thus, the proof is complete. ∎
Remark 4.2 (Normalization and local coherence).
One can rewrite (4.1) into
$\boldsymbol{v}_{i}=\frac{1}{\mu_{i}}\sum_{j\in N(i)}\boldsymbol{v}_{j},\quad\text{where }\mu_{i}=\left\|\sum_{j\in N(i)}\boldsymbol{v}_{j}\right\|=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}).$
The expression describes a geometric condition of equilibrium: At equilibrium, each phasor aligns with the direction of the sum of its neighbors’ vectors and is scaled to unit length. Although introduced as a normalization factor, the quantity $\mu_{i}$ carries physical meanings. If the neighbors are highly coherent, the sum vector is large, requiring a larger $\mu_{i}$ to normalize $\boldsymbol{v}_{i}$ to unit length; if the directions are dispersed, $\mu_{i}$ is small. In this sense, $\mu_{i}$ measures how strongly the phasors of node $i$’s neighbors align directionally. Indeed, defining the local order parameter
$R_{i}:=\frac{1}{\deg(i)}\left|\sum_{j\in N(i)}e^{i\theta_{j}}\right|,$
we have the relation $\mu_{i}=\deg(i)\cdot R_{i}$, showing that $\mu_{i}$ is proportional to the standard notion of local phase coherence.
Remark 4.3 (Equilibrium and force).
This equation admits a natural physical interpretation using the concept of force. Write each phasor $\boldsymbol{v}_{i}$ in complex form as $e^{i\theta_{i}}$. Then, intuitively, each neighboring node $j\in N(i)$ pulls node $i$ with a force given by $e^{I(\theta_{j}-\theta_{i})}$. Under this interpretation, equation (4.1) states that: At equilibrium, the total force acting on node $i$ sums to a real scalar. In other words, the net force is purely radial and contributes only to scaling. Due to the unit circle geometry $\left|\boldsymbol{v}_{i}\right|=1$ of the Kuramoto model, such a scaling force does not induce any motion while a tangential component can cause rotation.
Notation. (1) We will call $\boldsymbol{\theta}$ is an equilibrium of $G$ receives the force $\boldsymbol{\mu}=(\mu_{1},\cdots,\mu_{n})$, if the vetor-labaled graph representation of $\boldsymbol{\theta}$ satisfies (4.1). Note that, in such case, $\mu_{i}=\sum_{j}A_{ij}\cos(\theta_{j}-\theta_{i})$, for all $i\in V$; (2) At state $\boldsymbol{\theta}$, we say that node $i\in V$ is at equilibrium at $\boldsymbol{\theta}$ with respect to $G$ if there exists a scalar $\mu_{i}$ such that (8) holds. Clearly, if $\boldsymbol{\theta}$ is an equilibrium, then every node $i\in V$ is at equilibrium. In particular, we call a node $i$ is positively (or negatively, or zero) forced, if $\mu_{i}>0$ (or <0<0, or =0=0).
4.3. Force and stability
Having established the correspondence between vector alignment at each node and the equilibrium condition, we now turn to the question of stability.
Alignment alone does not guarantee stability. Intuitively, if there exists a node $i$ such that the angle between $\boldsymbol{v}_{i}$ and the resultant vector $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ is zero, a small perturbation will cause the system to return to equilibrium, much like a ball resting at the bottom of a potential valley. In contrast, if the angle is $\pi$, the system is analogous to a ball positioned at the peak of a hill: although it is momentarily at rest, a small perturbation will cause the ball to roll away.
In this section, this intuition is formally confirmed by using the relationship between the negative semi-definiteness of the Jacobian and its diagonal entries, as well as the connection between the stability of equilibrium in gradient dynamical systems and the local geometry of the corresponding energy function.
Figure 6. Illustration of different stability conditions for node $i$. Here, the solid arrow represents $v_{i}$, while the dashed arrow represents $\sum_{j\in N(i)}v_{j}$. The gray region indicates that, in order for node $i$ to attain equilibrium under non-negative forcing, the endpoint of $\sum_{j\in N(i)}v_{j}$ must lie within this sector. (A) Node $i$ is not at equilibrium since (8) does not hold; (B) Node $i$ is at equilibrium but negatively forced, and thus unstable. Therefore, $\boldsymbol{\theta}$ cannot be a second-order critical point. In cases (C) and (D), node $i$ is at equilibrium, and is zero forced and positively forced respectively. Thus, satisfies the necessary condition for $\boldsymbol{\theta}$ to be a second-order critical point.
Lemma 4.4.
Assume $\boldsymbol{\theta}$ is an equilibrium state of system $G=(V,E)$ or equivalently first order stationary point of (3). In other words, there exist real numbers $\mu_{1},\dots,\mu_{n}$ such that for all $i\in V$,
$\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
If $\boldsymbol{\theta}$ is a second order stationary point of (3), then all $\mu_{i}$ must be non-negative.
Remark 4.5 (Mechanical insight into node at stable equilibrium).
The equilibrium condition in Lemma 4.4 for node $i\in V$ has a mechanical interpretation. The sum of the phasors representing its neighbors can be visualized as a set of unit-length rods with directions. We place one phasor’s starting point at the origin, then connect the remaining phasors head-to-tail. The head-to-tail connected rods can freely adjust their orientations, but the endpoint of the last phasor must remain on the right-end unrestricted sliding track, to ensure node $i$ being stable there.
Figure 7. $\boldsymbol{v}_{i}$ is pinned down at zero angle, which determines the region where $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}$ should be (the right-end unrestricted region). The dashed vectors represent the phasors of nodes connected to ii.
Proof of Lemma 4.4.
We begin by noting that the $i$-th diagonal entry of the Hessian matrix $\mathbf{H}$ at $(\theta_{1},\cdots,\theta_{n})$ can be rewritten as
$\mathbf{H}_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
where the second equality uses the equilibrium condition (4.1), and the last equality is deduced from re-writing the unit vectors in condition $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}$ into the form using complex numbers. If there exists $\mu_{i}<0$, then equivalently $\mathbf{H}_{ii}<0$, which contradicts the positive definiteness of $\mathbf{H}$, see [Bha07] [HJ85]. Hence, $\boldsymbol{\theta}$ cannot be a second-order critical point. Consequently, it is neither a local minimum of (3) nor a stable equilibrium of (2). Therefore, $\mathbf{H}_{ii}=\mu_{i}>0$ holds for all $i\in[n]$, where $n$ is the size of $G$. ∎
To summarize, the lemma above shows that at any second-order critical point $\boldsymbol{\theta}$, the phasor of each node must be positively aligned with the sum of the phasors of its neighbors (whenever the latter is nonzero). That is,
$\boldsymbol{v}_{i}\;\Uparrow\;\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\boldsymbol{0},$
for all $i\in[n]$.
Notation. Here, We write $a\Uparrow b$, if both are nonzero and point in the same direction (i.e., the angle between them is zero). In the remainder of this manuscript, we will use this notation, as we find it more intuitive.
Remark 4.6 (This is only a necessary condition).
Conversely, at a critical point (equilibrium), all $\mu_{i}>0$ for all $i\in V$ is not sufficient to guarantee that $\boldsymbol{\theta}$ is a second-order critical point. As a counter-example, the equilibrium $\boldsymbol{\theta}^{*}=(0,0,0,0,\pi,\pi,\pi,\pi)$ of the Kuramoto model on the graph in Figure 8 is a SOSP.
12345678
Figure 8. At the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$, each node experiences a force from its neighbors that is aligned with its own direction. However, this configuration is not stable, as can be verified by examining the eigenvalues of the corresponding Hessian matrix.
Indeed, the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$ is not a local minimum of the energy function. When nodes $1,2,3,4$ are jointly perturbed slightly away from angle 0 in the same direction, they do not return to the original configuration. Instead, they continue to move toward the $\pi$-oriented nodes $5,6,7,8$, indicating a decrease in energy and confirming that the configuration is unstable. This can be rigorously proven by quadratic form $\boldsymbol{x}^{T}\boldsymbol{H}\boldsymbol{x}$, we omit it here.
4.4. Notations
The below definitions are nodewise concept since they are always made with respect to a node in certain state $\boldsymbol{\theta}$. For example, when we say that node $i$ attains equilibrium at $\boldsymbol{\theta}$, we mean that
$\mu_{i}\boldsymbol{v}_{i}\;=\;\sum_{j\in N(i)}\boldsymbol{v}_{j},\qquad\mu_{i}\in\mathbb{R}.$
Similarly, when we say that node $i$ attains a stable equilibrium at $\theta$, we mean the same balance equation (8) holds with $\mu_{i}\geq 0$. In the proof of global synchrony for threshold graphs, we frequently use this necessary condition. To make the reasoning more intuitive for the reader, we assign different colors to nodes representing different cases.
Condition on $i$ Equation Illustration
Equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\in\mathbb{R}$ $i$
Stable equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\geq 0$ $i$
Asymptotically stable (positively forced) $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}>0$ $i$
Zero-forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}=0$ $i$
Negatively forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}<0$ $i$
Table 1. Stability classification of node $i$ at equilibrium $\theta$. The sign of $\mu_{i}$ determines the type of stability, visualized by node color.
It is important to note that these definitions do not require the state $\boldsymbol{\theta}$ itself to be an equilibrium or stable equilibrium of the system. Indeed, at a given state $\boldsymbol{\theta}$ it is possible that some nodes satisfy the condition that their own phasor is parallel to the aggregated phasor formed by their neighbors. Such nodes are in equilibrium at this particular configuration $\boldsymbol{\theta}$; however, this does not imply that the configuration itself is an equilibrium state, since other nodes may not satisfy the same condition. Nevertheless, if $\boldsymbol{\theta}$ is indeed such an equilibrium, then the above conditions are of course implied, according to Lemmas 4.4 and 4.1.
5. Geometric twins and their phasor geometry
In this section, we introduce the concept of geometric twins (g-twins for abbreviation), that takes into account the unit vectors $\boldsymbol{v}_{i}=(\cos\theta_{i},\sin\theta_{i})$ associated with each node $i$ in a Kuramoto state $\boldsymbol{\theta}$. Two nodes are said to be g-twins if they satisfy certain geometric relationships at this configuration. It is worth noting that, due to the conceptual similarity, we borrow the term twins from graph theory. However, we precede it with the word geometric to emphasize that our definition depends on the positions of the node phasors in the specific configuration. This concept captures a form of geometric symmetry at $\boldsymbol{\theta}$ that will be crucial in our proof for global synchrony of threshold graphs.
5.1. Closed geometric twins
Definition 5.1 (Closed g-twins).
Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$. Two nodes $i,j\in V$ are called closed g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
$\bigl(\boldsymbol{v}_{i}\parallel\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
or equivalently
$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}$
where $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
If moreover both strengths are nonnegative, meaning that each vector $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ points in the same direction as the sum of the other and the common vector $\boldsymbol{q}$ (or that summation is the zero vector), then we call them stable closed g-twins.
Theorem 5.2 (Phasor geometry of closed g-twins).
Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of closed g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector falls into one of the following three cases:
  1. (1)
    $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$ and $\mu_{a}=\mu_{b}$;
  2. (2)
    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$ and $\mu_{a}+\mu_{b}=-2$ and $\mu_{a}\neq-1$;
  3. (3)
    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=-1$ and $\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=0$.
$\mu_{a}$$\mu_{b}$
Figure 9. Feasible region of strengths $(\mu_{a},\mu_{b})$ for closed g-twins. It lies on the lines $\mu_{a}=\mu_{b}$ (black, with $a$ and $b$ synchronized) and $\mu_{a}+\mu_{b}=-2$ (red, with $a$ and $b$ in antipodal position). At their intersection (1,1)(-1,-1), marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
Proof.
According to the definition of closed g-twins
$\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\quad\text{and}\quad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$
we have
$\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}-\boldsymbol{v}_{b}=\mu_{b}\boldsymbol{v}_{b}-\boldsymbol{v}_{a}$
Thus
$(\mu_{a}+1)\boldsymbol{v}_{a}=(\mu_{b}+1)\boldsymbol{v}_{b}.$ (13)
Case 1: $\mu_{a}=\mu_{b}=-1$. In this case, we have
$\boldsymbol{q}=\boldsymbol{v}_{a}+\boldsymbol{v}_{b}$
Case 2: $\mu_{a}\neq-1$ and $\mu_{b}\neq-1$. In this case, from (13), it must hold that
$\boldsymbol{v}_{a}\parallel\boldsymbol{v}_{b}.$
This leads to two subcases: (1) If $\mu_{a}=\mu_{b}\neq-1$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) If $\mu_{a}+1=-(\mu_{b}+1)$, then equivalently, we have $\mu_{a}+\mu_{b}=-2$. This implies $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$. ∎
Definition 5.3 (Stable closed g-twins).
Under the setting of Definition 5.1, we say that nodes $i$ and $j$ are stable closed g-twins if furthermore
$\bigl(\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\Uparrow\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
or equivalently
$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
Similarly, one can define open g-twins and stable open g-twins at state $\boldsymbol{\theta}$, though this concept will not be used in our proofs in this chapter. It may nevertheless be of independent interest in understanding the geometry of non-adjacent nodes with certain geometric symmetries at equilibria.
5.2. Open geometric twins
Definition 5.4 (Open g-twins).
Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$. Two nodes $i,j\in V$ are called open g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
$\left(\boldsymbol{v}_{i}\parallel\boldsymbol{q}\quad\text{and}\quad\boldsymbol{v}_{j}\parallel\boldsymbol{q}\right)\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
or equivalently,
$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}.$
Here $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
Definition 5.5 (Stable open g-twins).
Under the setting of Definition 5.4, we say that nodes $i$ and $j$ form geometric stable open twins if furthermore
$\boldsymbol{v}_{i}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
and
$\boldsymbol{v}_{j}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
or equivalently,
$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
Remark 5.6.
For brevity, whenever the conditions of Definition 5.4 are satisfied, we shall simply say that $i$ and $j$ form open g-twins with strengths $(\mu_{i},\mu_{j})$, with respect to the common vector $\boldsymbol{q}$. The terminology for closed g-twins, stable g-twins, and stable closed g-twins is used analogously.
Under the special geometric symmetries described above, the phasors of g-twins exhibit a particular symmetric configuration on the unit circle. This in turn allows us to obtain a precise characterization of the relative positions of the phases $\theta_{i}$ of nodes with such hidden geometric symmetry, both at equilibrium and at stable equilibrium of the Kuramoto model. We refer to the notation on node at (stable) equilibrium in Table 1 for clarity.
Theorem 5.7 (Phasor geometry of open g-twins).
Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of open g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector $\boldsymbol{q}$ falls into one of the following three cases:
  1. (1)
    $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$ and $\mu_{a}=\mu_{b}\neq 0$;
  2. (2)
    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$ and $\mu_{a}=-\mu_{b}\neq 0$;
  3. (3)
    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=0$.
$\mu_{a}$$\mu_{b}$
Figure 10. Feasible region of strengths $(\mu_{a},\mu_{b})$ for open g-twins. It lies on the lines $\mu_{a}=\mu_{b}$ (black, corresponding to $a$ and $b$ synchronized) and $\mu_{a}=-\mu_{b}$ (red, corresponding to $a$ and $b$ antipodal). At their intersection $(0,0)$, marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
Proof.
Since
$\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b},$ (14)
the following cases arise.
Case 1. $\mu_{a}=\mu_{b}=0$, then we have $\boldsymbol{q}=0$.
Case 2. If $\mu_{a},\mu_{b}\neq 0$, then, from 14 we have
$\boldsymbol{v}_{a}=\frac{1}{\mu_{a}}\boldsymbol{q}=\frac{\mu_{b}}{\mu_{a}}\boldsymbol{v}_{b}.$
Since $\|\boldsymbol{v}_{a}\|=\|\boldsymbol{v}_{b}\|=1$, it follows that $|\mu_{a}|=|\mu_{b}|$. We distinguish two subcases. If $\mu_{a}=\mu_{b}$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. If $\mu_{a}=-\mu_{b}$, then $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$, and one of $\mu_{a},\mu_{b}$ is negative. ∎
6. Formation of geometric twins
The motivation for introducing g-twins in the previous section comes from the proof of global synchronization on threshold graphs. In that proof, it turns out that synchronization does not rely solely on structural symmetry (i.e., forming structural twins), but can also arise from a more general geometric symmetry, where the corresponding phasors occupy special symmetric positions on the unit circle.
In this section, we first mention that structural twins at equilibrium constitute the simplest case of forming g-twins. We then move on to more general situations where nodes are not structural twins, yet under both specific geometric conditions (on the phasors) and accompanying structural conditions, they still form g-twins. Both cases will play a crucial role in the proof of global synchronization for threshold graphs.
6.1. Structural twins at equilibrium
Definition 6.1 (Structural Twins).
Let $G=(V,E)$ be a graph. Two nodes $i,j\in V$ are called structural closed twins if
$N[i]=N[j],$ (15)
and structural open twins if
$N(i)=N(j),$ (16)
(a) Closed twins ($N[i]=N[j]$).
(b) Open twins ($N(u)=N(v)$).
Figure 11. Illustration of structural twins.
Remark 6.2.
In the discussion above we used the term structural to emphasize that whether two nodes are twins depends purely on the graph structure. In graph theory they are simply called true or adjacent twins, and false or non-adjacent twins.
Clearly, if $\boldsymbol{\theta}$ is an equilibrium point of the Kuramoto model with respect to graph $G$, and two nodes $a$ and $b$ satisfy condition 15, then they must form a pair of closed g-twins. This is because, according to Lemma 4.1, the nodes $a$ and $b$ are structurally symmetric (satisfying (15)) 8, that is
[…]
$\boldsymbol{q}=\sum_{j\in N(a)\setminus\{b\}}\boldsymbol{v}_{j}=\sum_{j\in N(b)\setminus\{a\}}\boldsymbol{v}_{j}$
such that $\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4. The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14.
There, colors indicate node-wise stability: Blue indicates $\mu>0$ (stable), red for $\mu<0$ (unstable), black for $\mu=0$.
Figure 12. Structural closed twins attaining nodewise equilibrium.
Remark 6.3.
Here the case (d) is detailed in the following figure. One can clearly see that
$\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=\boldsymbol{0},$
where $N$ denotes the set of common neighbors of $a$ and $b$, excluding each other.
Figure 13. Geometric illustration of case (d) for false twins at equilibrium. The red arrow represents the vector sum $\boldsymbol{q}$, where $N$ is the common neighborhood of nodes $a$ and $b$, excluding each other. In this case, $\mu_{a}=\mu_{b}=-1$, and the three vectors $\boldsymbol{v}_{a}$, $\boldsymbol{v}_{b}$, and $\boldsymbol{q}$ form a triangle summing to zero.
Similarly, structural open twins $a,b$ at equilibrium in $\boldsymbol{\theta}$ are also geometric open twins, since there exists
$\boldsymbol{q}=\sum_{j\in N(a)}\boldsymbol{v}_{j}=\sum_{j\in N(b)}\boldsymbol{v}_{j}$
such that $\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4. The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14. The same as in the above mentioned closed twins case, that colors indicate their stability.
Figure 14. Structural open twins attaining nodewise equilibrium.
Our strategy for proving that a given graph admits global synchronization is to characterize the positions of all nodes at stable equilibrium (in fact, more general, at SOSP of $E(\boldsymbol{\theta})$). Furthermore, the lemma on structural twins at equilibrium shows that whenever two structural twins both attain stable equilibrium, they must synchronize if they are closed twins. It is worth noting, however, that in the case of open twins, synchronization is not guaranteed: it may happen that both strengths $\mu_{i}$ and $\mu_{j}$ vanish, in which case the phasors of $i$ and $j$ remain unconstrained.
At this point, we are already able to characterize certain new classes of globally synchronizing graphs, according to results in this section.
Theorem 6.4.
Complete split graphs, windmill graphs, complete bipartite graphs are globally synchronizing.
(a) Complete split graphs.
(b) Windmill graphs $W_{5,6}$.
(c) Complete bipartite.
6.2. Synchronous pendant extension
In this section, we present a scenario in which a group of nodes forms closed g-twins without being structural closed twins, that is called synchronous pendant extension. Specifically, consider a set $S$ that satisfies the structural assumption that the additional neighborhood of some subset $S_{1}\subseteq S$ is attached to $S_{1}$ in a leaf-like manner. The geometric assumption, on the other hand, is that this additional neighborhood synchronizes with $S_{1}$.
To prepare for the presentation and proof of Lemma 6.6, we first establish Lemma 6.5, which highlights the following fact: If there exists a synchronized, fully connected set $S$ of nodes with a synchronized pendant attached, then each node in $S$ is aligned with the vector sum of the phasors of its remaining neighbors, namely those not in $S$ nor in the pendant. From the perspective of synchronization propagation, this can be interpreted as synchronization spreading from $S$ together with its pendant to the group of all other neighbors.
Lemma 6.5.
Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
$W=Q\cup S\cup P.$
under the following structural assumptions:
(i) For all $i\in Q$, we have $N(i)\subseteq S$;
(ii) The subgraph $G[S]$ is a complete graph;
(iii)For any $i\in S$, its neighbourhood consists of the set $P$, its neighbours in $Q$, and the rest of the clique $S\setminus\{i\}$.
Assume further that all nodes in $Q\cup S$ are synchronized, i.e., there exists $\boldsymbol{v}$ such that $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}=\boldsymbol{v},\quad\forall i,j\in Q\cup S$. Then
$\sum_{i\in P}\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}\quad\text{or}\quad\sum_{i\in P}\boldsymbol{v}_{i}=\boldsymbol{0}.$
Figure 16. Illustration of Lemma 6.5. Elements in the figure follow the same conventions as in Figure 18. Moreover, a large node represents a set of nodes. An edge from a small node $A_{i}$ to a large node $S$ indicates that $A_{i}$ is connected to all nodes in $S$. The diagram depicts a configuration where a set of synchronized “false twin-like” nodes, with differing but synchronized neighbors, implies that their shared neighbors (the large white node) sum to a vector pointing in the same direction as the synchronized nodes.
Proof.
For any node $i\in S$,
$N(i)=Q\cup(S\setminus\{i\})\cup P$ (17)
and
$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$ (18)
Plug (17) into (18), we obtain
$\sum_{j\in Q\cup(S\setminus\{i\})\cup P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
Since $Q\cup S$ synchronize, we have
$(|Q|+|S|-1)\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
Thus
$\sum_{j\in P}\boldsymbol{v}_{j}=\left(\mu_{i}-\left(|Q|+|S|-1\right)\right)\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
Now, we are left to show $\mu_{i}-\left(|Q|+|S|-1\right)<0$ is not possible. Indeed, let $\boldsymbol{v}$ be an $n$-dimensional vector where the $i$-th element is one if $i\in Q\cup S$, and zero otherwise. Then
$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\cup S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (19)
$=\sum_{i\in S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$
$=\sum_{i\in S}\sum_{j\in P}\boldsymbol{v}_{i}\cdot\boldsymbol{v}_{j}$
$=\sum_{i\in S}\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}$
$=|S|\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}$
where the second equality is due to there is no connection between $Q$ and $P$. Clearly, if $\mu_{i}-\left(|Q|+|S|-1\right)<0$, then the above form is negative, and this contradicts to the stability of $\boldsymbol{\theta}$. The proof is complete. ∎
We are now ready to present the case of a synchronous pendant, which constitutes one possible configuration leading to the formation of geometric stable closed twins. This case will serve as one of the core lemmas in our proof for threshold graphs.
Consider a set of nodes $S$ attaining stable equilibrium at $\boldsymbol{\theta}$ which induces a clique [^4] Forming a clique guarantees the “closed” property required for closed twins.. $S$ can be divided into $S_{1}$ which connects only to a set $P$, and a synchronized group $S_{2}$ which in addition connects to a synchronized pendant set [^5] Specifically, $Q$ can be understood as a generalization of a leaf attached to one node, and it is well known that the presence of a leaf does not affect the stability of the state. $Q$. We will show that although $S_{1}$ and $S_{2}$ are not structurally identical, this difference, arising from the synchronous pendant, does not interfere with synchronization, in other words, the two groups behave as geometric stable closed twins.
The proof of Lemma 6.6 is based on Lemma 6.5. If we can decompose the set $P$ there into $\{A\}\cup P$, where $A$ connects to all nodes in $S$, and all nodes in $P$, then we have the following Lemma.
Lemma 6.6.
Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
$W=Q\cup S_{1}\cup S_{2}\cup P.$
under the following structural assumptions:
(i) For all $i\in Q$, we have $N(i)\subseteq S_{1}$;
(ii) The subgraph $G[S_{1}\cup S_{2}]$ is a complete graph;
(iii) For any $i\in S_{1}\cup S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\cup S_{2}$. Formally, if we denote $Q_{i}=N(i)\cap Q$, then
$N(i)\;=\;P\cup Q_{i}\cup\bigl((S_{1}\cup S_{2})\setminus\{i\}\bigr).$
Equivalently,
$N(i)\setminus\bigl(Q\cup(S_{1}\cup S_{2})\bigr)\;=\;P.$
Assume further that all nodes in $Q\cup S_{1}$ are synchronized, i.e.,
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\cup S_{1}.$
Then the set $S_{1}\cup S_{2}$ forms a group of geometric stable closed twins, therefore they synchronize.
Figure 17. Illustration of Lemma 6.6. Elements follow the same conventions as in previous figures. The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes. Each orange arc separates a small blue node from a group of nodes it connects to, indicating that the vector of the blue node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc. The figure illustrates how the synchronized block $A\cup S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
Proof.
Observe that the set $S_{1}$ induces a fully-connected subgraph, and $Q$ is a synchronous pendant to $S_{1}$, then according to Lemma 6.6, we have
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
Combine this with the fact
$\boldsymbol{v}_{i}=\boldsymbol{v}_{j}\quad\forall i\in S_{1},j\in S_{2},$
we obtain
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
Equivalently, we can write
$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (20)
Another observation is on nodes in $S_{2}$ that
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}$
since every node in $S_{2}$ attains stable equilibrium in $\boldsymbol{\theta}$. Equivalently, we can write
$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (21)
Combined the facts 20 and 21, we get nodes in $S_{1}\cup S_{2}$ are geometric stable closed twins. ∎
7. Main proof: the induction argument
7.1. One more key lemma
To complete the argument, we establish one more lemma prior to the proof of the main theorem.
Lemma 7.1 (Synchronization of closed twins with common neighbour).
Let $S\subseteq V$ be a set of structural closed twins in a graph $G$, and let among their common neighbourhood, there exists a node $A$ such that its neighbourhood is exactly $S$, i.e., $N(A)=S$. Assume that all nodes in $S\cup\{A\}$ attain stable equilibrium at configuration $\boldsymbol{\theta}$, that is,
$\forall i\in S\cup\{A\},\quad\mu_{i}v_{i}=\sum_{j\in N(i)}v_{j}\quad\text{ with }\quad\mu_{i}\geq 0,$
then the nodes in $S\cup\{A\}$ synchronize at $\boldsymbol{\theta}$.
Figure 18. Illustration of Lemma 7.1. All green nodes attain stable equilibrium. The set $S$ (middle row) forms a set of false twins, and node AA (top) has neighborhood exactly $S$. Thin lines represent the edges in the original graph. The conclusion is that all green nodes must synchronize, indicated by the thick edges in the right diagram.
Proof.
Since $S$ forms a set of structural stable closed twins, we have $i,j\in S$, we have $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$ and let us denote this common phasor as $\boldsymbol{v}$. Moreover, since $A$ only connects to the set $S$ which is synchronizing, and it attains stable equilibrium at $\boldsymbol{\theta}$, we have
$\mu_{A}\boldsymbol{v}_{A}=\sum_{j\in N(A)}\boldsymbol{v}_{j}=\sum_{j\in S}\boldsymbol{v}_{j}=|S|\boldsymbol{v}$
Thus $\boldsymbol{v}_{A}=\boldsymbol{v}$. The proof is complete. ∎
7.2. Illustrative example
In this section, we present a simple but general enough example to illustrate how to use Lemma 6.6 and 7.1 to establish global synchronization of connected threshold graphs. Specifically, to show how synchronization starts from a local substructure and eventually propagates to the entire threshold graph. The proof for all connected threshold graphs follows the same idea, with additional work needed for full generalization.
Let us take a look at the threshold graph corresponding to the sequence
$0101$
We use capital letters to represent the nodes corresponding to each bit in the sequence, including the initial starting node.
A0B1C0d1E0f1G0h1I
The corresponding threshold graph is shown as follows.
Figure 19. Threshold graph corresponding to the sequence $0101$, with vertices labeled AA to $I$ in the order of their addition.
Concluding global synchronization step by step.
To show that this graph globally synchronizes, we prove that any second order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ on the threshold graph must be synchronous. Clearly, any second order stationary point $\boldsymbol{\theta}$ is also a local minimum of $E_{G}(\boldsymbol{\theta})$ or stable equilibrium of the Kuramoto model 2. Thus, the Lemmas established in Chapter 4 applies here. That is, every node attains stable equilibrium in $\boldsymbol{\theta}$, according to Lemma 4.4.
We will use Lemmas 6.6 and 7.1 repeatedly in the following illustration.
Step 1. Start with the edge between $H$ and $I$. Synchronization between nodes $H$ and $I$ can be obtained from stability condition of $H$ soely. Indeed, since $H$ attains stable equilibrium at $\boldsymbol{\theta}$, there exists $\mu_{H}\geq 0$ such that $\mu_{H}\boldsymbol{v}_{H}=\sum_{j\in N(H)}\boldsymbol{v}_{j}$ and we know that $\sum_{j\in N(H)}\boldsymbol{v}_{j}=\boldsymbol{v}_{I}$, we obtain $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$. One could also use Lemma to argue $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$, as this is the simplest case of the lemma.
Step 2. Observe that the pair of node $G$ and $I$ forms geometric stable closed twins at $\boldsymbol{\theta}$. Indeed, the only difference between $N[G]$ and $N[I]$ is that $I$ additionally has a synchronizing pendant $H$. Thus $\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$. Now nodes $G,I,H$ are synchronized.
Step 3. Note that node $F$ connects to $G$ and $I$ which are synchronized as established in Step 2, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
Step 4. Observe that nodes $E,G,I$ induce a fully connected subgraph, and nodes $G$ and $I$ have additional synchronizing pendant, therefore, according to Lemma 6.6, three of them form geometric stable closed twins. Thus, $\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step 4. Observe that node $D$ connects to $E,G,I$ which are synchronzed, thus $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
Step 5. Observe that nodes $C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $E,G,I$ admit additional synchronzing pendant $D,F,H$. Consequently $C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
Step 6. Now note that the node $B$ connects to a group of synchronzing nodes $C,E,G,I$. Therefore according to Lemma, node $B$ synchronzes with them as well. That is, $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
Step 7. Observe that nodes $A,C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $C,E,G,I$ admit additional synchronzing pendant $B,D,F,H$. Consequently $A,C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
In conclusion, $\boldsymbol{\theta}$ is a synchronized state, and thus the threshold graph globally synchronizes.
7.3. Main proof
Before proceeding, we introduce some notation. Any connected threshold graph can be specified by a bit sequence consisting of alternating blocks of $1$’s and 0’s. Since the graph is connected, the last bit must be $1$. We group the sequence into maximal contiguous blocks of $1$’s and 0’s. Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block directly above $U_{i}$ (if any). Thus, the sequence takes the form either
$U_{1},I_{2},U_{2},\dots,I_{k-1},U_{k-1},I_{k},U_{k},$
or
$I_{1},U_{1},I_{2},U_{2},\dots,I_{k-1},U_{k-1},I_{k},U_{k},$
depending on whether the sequence begins with a 0-block.
Proof of Theorem 1.2.
We prove the main theorem 1.2 by induction. First, assume $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$. Our goal is to show $\theta_{1}=\cdots=\theta_{n}$. Let the sequence defining the threshold graph consist of maximal contiguous blocks of 1’s and 0’s, arranged as either:
$U_{1},I_{1},U_{2},I_{2},\dots,U_{k}\quad\text{or}\quad I_{0},U_{1},I_{1},U_{2},I_{2},\dots,U_{k}.$
Base case. At a stable equilibrium $\boldsymbol{\theta}$, the last block $U_{k}$ must synchronize. This follows directly from Lemma 5.2, as the nodes in $U_{k}$ are false twins and thus must synchronize.
Inductive hypothesis. Assume that the nodes in the last $m$ blocks of 1’s, $U_{k-m+1},\dots,U_{k}$, are synchronized.
Inductive step. Here, we show that the synchronization of $\bigcup_{i=k-m+1}^{k}U_{i}$ propagates to $\bigcup_{i=k-m}^{k}U_{i}$, since the preceding block $U_{k-m}$ must synchronize with $\bigcup_{i=k-m+1}^{k}U_{i}$. Indeed, since all nodes in $\bigcup_{i=k-m+1}^{k}U_{i}$ synchronize, any node added as isolated nodes whose neighbourhood is a subset of $\bigcup_{i=k-m+1}^{k}U_{i}$ must synchronize with them as well. These nodes are exactly nodes in $\bigcup_{i=k-m+1}^{k}I_{i}$. That is, every nodes $i,j\in\bigcup_{i=k-m+1}^{k}U_{i}\cup I_{i}$ satisfy $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$.
Observe that $\bigcup_{i=k-m+1}^{k}I_{i}$ are in fact synchronizing pendant attaches to $\bigcup_{i=k-m+1}^{k}U_{i}$, and they are the only difference of neighbourhood between the $U_{k-m}$ and $\bigcup_{i=k-m+1}^{k}U_{i}$, thus according to Lemma 6.6, nodes in $\bigcup_{i=k-m}^{k}U_{i}$ form a set of geometric stable closed twins, and they must synchronize.
Conclusion. All nodes in $\cup_{i=1}^{k}\left(U_{i}\cup I_{i}\right)$ or $I_{0}\cup\left(\cup_{i=1}^{k}(U_{i}\cup I_{i})\right)$ synchronize. Namely $\theta_{1}=\cdots=\theta_{n}$, thus any threshold graph is globally synchronizing as long as it is connected. ∎
Acknowledgements
We would like to thank Afonso S. Bandeira for an insightful discussion at the early stage of this work.

### v2

@@
 Abstract.
-Place $n$ nodes on the unit circle and connect some pairs with springs that pull connected nodes toward each other.
-Which graphs always drive all nodes to converge to a single point, no matter where they start?
-This is known as the global synchronization problem of the Kuramoto model and is closely tied to the benign nonconvexity of its synchronization landscape.
-In this work, we show how this dynamical and continuous question can be reduced to a static and combinatorial one.
-We prove that threshold graphs, the graph class closed under adding isolated nodes and adding universal nodes, are globally synchronizing, based purely on a combinatorial analysis of unit-vector configurations in the plane.
-It is particularly striking that threshold graphs range from extremely sparse structures such as stars to the complete graph, suggesting that global synchronization in this class cannot be attributed to density.
-Our proof shows that global synchronization can instead emerge from a form of geometric symmetry.
+The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph.
+Each edge acts like a spring that pulls the two adjacent oscillators toward each other whenever their phases differ.
+A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions, except for a set of measure zero.
+This property is tightly linked to the benign nonconvexity of the model’s energy landscape.
+Existing guarantees for global synchronization rely on minimum-degree thresholds, which require the graph to be highly dense.
+In this work, we show that connected threshold graphs, whose density may vary from $2/n$ to $1$, are globally synchronizing.
+Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
 1. Introduction
 The study of synchronization traces back to the 17th century, when Christiaan Huygens first observed the spontaneous synchronization of pendulum clocks.
-Since then, synchronization phenomena have been widely studied across both natural and technological systems, from the flashing of fireflies [BB68, SS93] and the rhythmic beating of heart cells [Win67, MS06] to synchronization in power grids [DB13], and the Transformer architecture [GLPR23a, CRMB24].
-The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kur75], strikes a balance between mathematical tractability and sufficient complexity, has become one of the key mathematical models for understanding the collective dynamics of coupled oscillators.
-The Kuramoto model describes a system of $n$ oscillators continuously evolving on the unit circle.
-Each vertex $i\in V$ corresponds to an oscillator with phase (angle) $\theta_{i}:\mathbb{R}\rightarrow\mathbb{S}^{1}$.
-The dynamics of phase $\theta_{i}$ of each oscillator evolves according to:
-$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),$ (1)
-This means that each oscillator’s phase dynamics is governed by its intrinsic frequency $\omega_{i}$ while also being influenced by the phase differences with its neighbors in a nonlinear manner.
-There are two lines of research regarding this model focusing heterogenous and homogenouse case respectively.
-Studies on the former case primarily consider a systems of oscillators with heterogenous intrinsic frequences $\{\omega_{i}\}_{i=1}^{N}$ coupled through a complete graph with identical weight $\boldsymbol{A}_{ij}=\frac{K}{N}$.
-Here, $K$ determines how strongly oscillators influence each other.
-A fundamental question in the heterogeneous setting is to determine the critical coupling strength $K_{c}$ such that, for $K>K_{c}$, the oscillators achieve frequency synchronization, i.e., $\dot{\theta}_{i}-\dot{\theta}_{j}\to 0$, meaning that all oscillators rotate with the same angular velocity.
-If this indeed holds, it is natural to further ask whether the oscillators not only rotate with the same angular velocity, but also converge to the same phase, known as phase synchronization, that is, whether $\theta_{i}-\theta_{j}\rightarrow 0$, as $t\rightarrow\infty$.
-The study of synchronization in the heterogeneous Kuramoto model is particularly relevant to real-world systems.
-In many natural and engineered networks such as power grids, neuronal circuits, and circadian rhythms, the individual oscillators typically have distinct natural frequencies.
-Within this heterogeneous setting, a rich variety of nonlinear dynamical phenomena can emerge, including partial synchronization, multistability, and even chaotic behavior.
-For further background and details, we refer the reader to [Str00, ABV+05, DB14].
-Recent works on the Kuramoto model has moved beyond the fully connected case to more general interaction graphs.
-Such settings are not only mathematically appealing but also more realistic, since interactions in natural and engineered systems are rarely all-to-all.
-In this work, we focus on the homogeneous case with identical intrinsic frequencies ($\omega_{i}=\omega$ for all $i$) and aim to understand how graph structure governs synchronization.
-To begin with, we first remark that in this case, frequency synchronization always holds and the remaining question is whether phasr synchronization can be achieved.
-This is not obvious if we only look at the equation $\dot{\theta}_{i}=\omega+\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})$, for all $i$, since it tells that the instantaneous angular velocity of each node depends both on this common intrinsic term and on the coupling with its neighbors.
-However, applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, the equations becomes
-$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),\quad i\in\{1,\cdots,n\},$ (2)
-and this reduced system is in fact a gradient flow of the Kuramoto energy function
-$E(\theta)=\frac{1}{2}\sum_{1\leq i<j\leq n}\boldsymbol{A}_{ij}(1-\cos(\theta_{i}-\theta_{j}))$ (3)
-as its gradient satisfies $\dot{\theta_{i}}{dt}=-\frac{\partial E}{\partial\theta_{i}}$.
+Since then, synchronization phenomena have been widely studied across both natural and technological systems.
+For example, the flashing of fireflies [Buck1968, Strogatz1993], the rhythmic beating of heart cells [Winfree1967, Mikhailov2002], synchronization in power grids [dorfler2012synchronization], coordinated motion in robotic networks [OlfatiSaber2006], modern artificial intelligence models such as neural networks [KuramotoNeurons2024], and Transformer architecture [Geshkovski2025, Criscitiello2024].
+A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness.
+The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kuramoto1975], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
+The Kuramoto model describes $n$ oscillators on the unit circle, coupled according to a graph $G=(V,E)$ with adjacency matrix $A$. Mathematically, it is a set of ordinary differential equations
+$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$ (1)
+where $\theta_{i}:\mathbb{R}\to\mathbb{S}^{1}$ denotes the phase of oscillator $i$.
+Specifically, each oscillator is driven by its intrinsic frequency $\omega_{i}$ and by the nonlinear interactions determined by the phase differences with its neighbors.
+Early studies of the Kuramoto model primarily considered systems of oscillators with heterogeneous intrinsic frequencies $\{\omega_{i}\}_{i=1}^{n}$, coupled through a complete graph with uniform weights $\boldsymbol{A}_{ij}=K/n$.
+Here, $K$ denotes the strength of the coupling between oscillators.
+A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized.
+This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [michaels1987sinoatrial], technological applications such as deep brain stimulation [tass2003desynchronizing, nabi2011singleinput] and power systems [chiang1995direct, dorfler2012synchronization].
+For comprehensive surveys of this line of work and its applications, we refer to [strogatz2000, acebron2005, dorfler2014].
+Another line of research focuses on the homogeneous Kuramoto model where all intrinsic frequencies are identical, i.e., $\omega_{i}=\omega$ for all $i\in\{1,\cdots,n\}$.
+In this case, frequency synchronization is always achieved.
+This becomes immediate after applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, under which the dynamics reduce to the gradient system
+$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in[n].$ (2)
 By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$.
-Hence, the angle differences do not lead to frequency discrepancies in the asymptotic sense.
-In other words, frequency synchronization always holds, as they share common frequency $\omega$.
-Now the real challenge in the homogeneous case lies in understanding whether phase synchronization can always be achieved, regardless of the initial phase configurations.
-It is evident that this property is entirely determined by the graph structure, since the adjacency matrix $\mathbf{A}$ of the graph is the only external parameter in the homogeneous Kuramoto model (2) and equivalently in its energy formulation (3).
-Thus, we define:
+Thus, in the co-rotating frame, the asymptotic frequency of every oscillator is zero.
+Transforming back to the original coordinates, this implies that the oscillators always achieve frequency synchronization.
+In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [Maistrenko2005].
+In the homogeneous setting, the remaining challenge is to determine whether phase synchronization occurs.
+This clearly depends on the initial conditions.
+If all oscillators start within a half circle, they converge to a common phase.
+However, beyond the half-circle region, the global behavior remained poorly understood for a long time.
+Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero?
+This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [bandeira2016low, ling2019landscape, ling2023local] and the dynamics of Transformer architectures [Geshkovski2025, geshkovski2023emergence].
+In this work, we follow this line of research, and beginning with a precise formulation of the problem.
+Problem Setting.
 Definition 1.1 (Globally synchronizing graph).
-We say that a graph $G$ with adjacency matrix $\mathbf{A}$ is globally synchronizing if, all oscillators converge to the same phase, i.e.,
-$\lim_{t\to\infty}\theta_{i}(t)-\theta_{j}(t)=0\quad\text{for all }i,j\in V.$
-for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero.
-In other words, almost all trajectory converges to the synchronous state $\theta_{i}=\theta_{j}$, for all $i,j\in V(G)$.
+We say that a graph $G=(V,E)$ is globally synchronizing if, for almost all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, the solution to (2) satisfies
+$\lim_{t\to\infty}(\theta_{i}(t)-\theta_{j}(t))=0\qquad\text{for all }i,j\in V.$
 The key question is now stated as:
-On which graphs, the Kuramoto model (2) exhibits global synchronization?
-This seemingly simple description conceals a mathematically challenging task.
-To address this problem, we note that the Kuramoto model (2) can be formulated as a gradient flow of the energy function (3), as mentioned before.
-Hence, understanding the geometry of (3) provides direct insight into the long-term dynamics of (2).
-In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous equilibria.
+Which graphs are globally synchronizing?
+To build some intuition for the model (2) and the definition (1.1), one may imagine placing $n$ oscillators on the unit circle and connecting them by edges specified by the adjacency matrix $A$.
+If there is an edge between nodes $i$ and $j$, then whenever their phases differ, the interaction tends to reduce this difference and pull them toward a common phase.
+[…]
+Figure 1. A spring analogy for a coupled oscillator network. The label on each curved arrow marks which node exerts the force.
+From the above illustration of the complete graph $K_{3}$, one may ask whether the three oscillators will eventually converge to a common position when placed as shown in Figure 1.
+Each node is subject to two forces, for instance, node $1$ is pulled toward node $2$ on one side and toward node $3$ on the other.
+So it is not immediately clear in which direction it will move, let alone whether all three nodes will ultimately coincide.
+Even in this simple example, the difficulty of the problem becomes apparent.
+Nevertheless, one may form an intuitive if somewhat heuristic expectation: since in a complete graph each oscillator tends to move closer to every other oscillator, the abundance of interactions may indeed drive the system to synchronize from almost all initial configurations.
+Historically, Verwoerd and Mason [verwoerd2007global] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing.
+This conjecture was resolved by Richard Taylor in 2012 [taylor2012there], and global synchronization of complete graphs has since become a standard benchmark result in this line of research.
+For classical results on the Kuramoto model on complete graphs, we refer to [acebron2005, strogatz2000, PikovskyRosenblum2015, Bick2020].
+In the same work [taylor2012there], Taylor also initiated what is now known as the extremal combinatorics version of the problem [bandeira2025randomstrasse101openproblems2024], which sparked substantial interest and subsequent developments.
+A more detailed discussion of these advances will be given in Prior Work.
+One standard approach to this problem is to study the corresponding energy landscape.
+Note that the Kuramoto model (2) can be formulated as a gradient flow of a certain energy function.
+Precisely, the system of ordinary differential equations (2) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$.
+Its trajectories evolve along the steepest descent direction of $E$.) with energy function $E_{G}(\boldsymbol{\theta})$, satisfying
+$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$ (3)
+where $E_{G}(\boldsymbol{\theta}):(\mathbb{S}^{1})^{|V|}\rightarrow\mathbb{R}$ with graph $G=(V,E)$ is defined as
+$E_{G}(\boldsymbol{\theta})$= $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\cos(\theta_{i}-\theta_{j})\big)$ (4)
+$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle v_{i},v_{j}\rangle\big).$
+Here, $v_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
+Clearly, this function is nonconvex.
+Moreover, understanding the geometry of (4) provides direct insight into the long-term dynamics of (2).
+In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria.
 If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing.
-In contrast, if the energy function admits no spurious local minima, then the graph $G$ is globally synchronizing, since this rules out convergence to any non-synchronous equilibrium except possibly for a set of initial conditions of measure zero.
-A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality together with the center manifold theorem; see Lemma A.1 in [GLPR23b] for an excellent exposition.
-Contributions and prior work.
-Aiming to understand which graphs and how them foster global synchronization, previous studies on this problem can be broadly divided into two lines of research: (i) investigating whether random graphs globally synchronize, and (ii) determining the minimum edge density required to ensure global synchronization.
-Both approaches are, in a certain sense, grounded in the fact that the complete graph is globally synchronizing, and that both random graphs and very dense graphs tend to behave like complete graphs.
-Intuitively, a graph of sufficiently high density becomes homogeneous enough to approximate a complete graph.
-In this sense, full connectivity is not strictly required, that is, global synchronization can still emerge with a small gap from the fully connected case.
-In this work, we prove that threshold graphs, as a graph class that by no means closed to complete, but still globally synchronize.
-(i) What are threshold graphs?
+In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing.
+A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lageman2007] together with the center manifold theorem; see Lemma A.1 in [Geshkovski2025] for an excellent exposition.
+Prior Work.
+Existing guarantees for global synchronization on deterministic graphs focus on the extremal combinatorics version of the problem.
+The aim is to identify minimum-degree thresholds above which a graph is guaranteed to be globally synchronizing.
+Problem 1.
+Let $\delta(G)$ denote the minimum degree of $G$. Find a constant $\mu_{c}\in[0,1]$ such that
+$\delta(G)\ \geq\ \mu_{c}(n-1)\ \Longrightarrow\ G\ \text{is globally synchronizing},$
+whereas for every $\varepsilon>0$ there exist graphs with $\delta(G)\leq(\mu_{c}-\varepsilon)(n-1)$ that are not globally synchronizing.
+The intuition behind this problem is that when every node is connected to a sufficiently large fraction of the network, the graph becomes dense enough for the collective attractive forces to guarantee convergence to synchrony.
+A remarkable property of these degree-based criteria is that, although they are only sufficient, they apply any graph meeting the threshold.
+This means global synchronization is guaranteed regardless of the graph’s finer structural details.
+Several advances have been made.
+The first result, due to Taylor [taylor2012there] in 2012, established a minimum-degree threshold of 0.9375(n1)0.9375(n-1).
+Subsequently, Ling [ling2019landscape] proved that global synchronization holds whenever the minimum degree exceeds 0.7929(n1)0.7929(n-1), and Lu [lu2021synchronization] further improved this bound to 0.7889(n1)0.7889(n-1).
+The conjecture asserts that the minimum-degree threshold for global synchronization is exactly $3/4$, and below this density, even very dense graphs may fail to synchronize.
+We refer to Conjecture 5 (density threshold for global synchrony) in [bandeira2025randomstrasse101openproblems2024] for a formal statement of this problem.
+Global synchrony on random graphs constitutes another important line of inquiry.
+Ling, Xu, and Bandeira [ling2019landscape] proved that for Erdős–Rényi graphs, global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$.
+This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [kassabov2022global].
+Both works raised the conjecture that synchronization should already occur as soon as the graph is connected, that is, for any $\epsilon>0$ and $p\geq(1+\epsilon)\log(n)/n$.
+This conjecture was recently resolved in [abdalla2024expandergraphsgloballysynchronizing].
+Further improvement has been made in [jain2025random, mcrae2025benign].
+For results on global synchrony of random geometric graphs, see [abdalla2024guarantees].
+Interestingly, recent work has pointed out that the idealized transformer dynamics admits a similar optimization form as the Kuramoto energy function [Geshkovski2025].
+This gives rise to the so-called nonlinear setting of the Kuramoto model, obtained by placing a nonlinear function on top of the inner product $E_{f}(\boldsymbol{\theta})=\-\sum_{1\leq i,j\leq n}A_{ij}f\big(\cos(\theta_{i}-\theta_{j})\big).$ The classical Kuramoto case corresponds to the linear function $f(t)=t$, while in the transformer dynamics one takes $f(t)=e^{\beta t}$ where $\beta>0.$ In this setting, the central question is to determine the values of $\beta$ for which global synchronization occurs, that is, the tokens collapse into a single cluster.
+Equivalently, this means that all token are making uniform predictions.
+A well-known conjecture predicted that global synchronization holds for all $\beta\geq 0$.
+The problem was originally formulated in [Geshkovski2025], with an initial result for $\beta\geq\Omega(1/n)$.
+It was subsequently improved in [Criscitiello2024] to cover the range $\beta\in[0,1]$, and was very recently settled by Polyanskiy, Rigollet, and Yao [polyanskiy2025synchronization].
+For more work connecting the Kuramoto model and Transformer (or attention) dynamics, see [geshkovski2023emergence, karagodin2024clustering, chen2025quantitative, polyanskiy2025synchronization, rigollet2025mean, Geshkovski2024] and the references therein.
+Our Contribution.
+Existing sufficient conditions for global synchronization of deterministic graphs require the graph to be sufficiently dense.
+In this work, we present a class of graphs that does not satisfy this minimum-degree requirement and whose densities range from 0 to $1$, yet globally synchronizes nevertheless.
+Through this result.
+Our main result is the following:
+Theorem 1.2.
+Any connected threshold graph globally synchronizes.
+In the following, we briefly introduce the definition of threshold graphs and the motivation for studying their global synchronization.
+We conclude this section with a discussion of the structural mechanism that underlies the global synchronization of this graph class.
+(1) What Are Threshold Graphs?
 Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices.
 This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one.
-Thus our goal is to show all threshold graphs with sequence ending by $1$ must globally synchronizes, because otherwise it is not connected.
+Note that the binary sequence of any connected threshold graph must end with a 1; otherwise, the graph is not connected since there exists at least one isolated node.
+A full formal definition and additional characterizations are given in Section 2.
 Figure 2.
 Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$).
-Node labels show the construction order: 11 is the initial vertex, later numbers follow the binary code.
+Node labels indicate the construction order: vertex $1$ is the initial white vertex, and all subsequent labels follow the binary code of the construction.
 White and black circles denote isolated and dominating vertices (bit 0 and bit 11).
 Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
-(ii) Why threshold graphs?
-Many theorems start from initial observations, from which suitable conjectures are formulated.
-Our case is no exception.
-Our investigation into global synchronization on threshold graphs began with a simple observation: complete graphs and star graphs, respectively the densest and sparsest connected graphs in the class of threshold graphs, both exhibit global synchronization.
+(2) Why Threshold Graphs?
+Many theorems begin with initial observations from which suitable conjectures are formulated, and our case is no exception.
+The investigation into global synchronization on threshold graphs began with several simple observations:
+Observation 1. Star graphs and complete graphs are respectively the densest and sparsest connected threshold graphs, and both exhibit global synchronization.
+$\bigl(\underbrace{\texttt{0}\;\texttt{0}\;\cdots\;\texttt{0}}_{\text{all 0}},\;\texttt{1}\bigr)$ $\bigl(\underbrace{\texttt{1}\;\texttt{1}\;\cdots\;\texttt{1}}_{\text{all 1}}\bigr)$
+Star graph $S_{n}$ Complete graph $K_{n}$
 Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize.
-Indeed, although the two extremes synchronize, this does not necessarily imply that the entire spectrum of threshold graphs does.
+Indeed, although the two extreme cases synchronize, this does not by itself imply that the entire spectrum of threshold graphs does.
 However, two further observations strengthened our confidence in this conjecture.
-Firstly, in threshold graphs, the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N[u]\subseteq N(v)$ or $N[v]\subseteq N(u)$.
-This total preorder of vicinities suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization.
-Secondly, no counter-example was found in our numerical experiments up to 50 size graphs [^1] First, we transform the equilibrium equations of the Kuramoto model into a system of polynomial equations.
-We then use Bertini software package [BHSW13] to compute the zeros of this system, and subsequently evaluate the Jacobian at each solution to determine their stability..
-These evidence encourages us to investigate whether all connected threshold graphs global synchronize.
-In particular, we established the following result:
-Theorem 1.2.
-Any connected threshold graph globally synchronizes.
+Observation 2.
+One of the many equivalent definitions of threshold graphs requires that their neighborhood-inclusion preorder is total, i.e., the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N(u)\subseteq N[v]$ or $N(v)\subseteq N(u)$.
+This total preorder of vicinities [threshold_graphs_book] suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization.
+Observation 3.
+While far from exhaustive, our numerical experiments did not reveal any non-synchronous stable equilibria on connected threshold graphs.
+Our experiments are twofold: (a) To examine global synchronization precisely in small systems, we employed computational algebraic geometry following [mehta2015algebraic].
+More explicitly, we reformulated the equilibrium equations of the Kuramoto model as a system of polynomial equations and used the Bertini software package [BHSW13] to compute all solutions.
+We then evaluated the Jacobian at each solution to determine its stability.
+(b) For larger size systems (up to 200), using ode45 in Matlab, we initialized a large number of random initial conditions and observed that all trajectories converged to the fully synchronized state.
+(3) Differences Between Our Approach and Existing Techniques.
 Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma.
 This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony.
-Based on this observation, the bootstrap strategy attempts to show that all second-order critical points must lie within a half-circle.
-If successful, this rules out the existence of non-synchronous equilibria and thereby implies global synchronization.
+Based on this observation, the bootstrap strategy attempts to show that all second-order stationary points must lie within a half-circle.
+If successful, this rules out the existence of non-synchronous state that the dynamics may converge to and thereby implies global synchronization.
 However, for threshold graphs, it remains unclear how to apply this strategy.
@@
 As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
-In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometry condition.
-This condition naturally leads us to study structural twins (pairs of nodes sharing the same neighborhood in the graph) as well as a new notion we introduce, that of geometric twins (a broader concept depending jointly on the graph structure and the equilibrium configuration $\theta$).
-These notions clarify how the relative positions of phasors behave at second order stationary point of the energy landscape $\mathbb{E}(\boldsymbol{\theta)}$.
+In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometric condition.
+This condition naturally leads us to study twins in the graph-theoretic sense, as well as a broader notion that we introduce, which we call geometric twins.
+The latter depends jointly on the graph structure and on the equilibrium configuration $\boldsymbol{\theta}$.
+These notions clarify how the relative positions of phases behave at second-order stationary point of the energy landscape $E(\boldsymbol{\theta})$.
 Building on this understanding, we employ an inductive strategy: synchronization is first established locally on small substructures, and then shown to propagate across the whole graph along the neighborhood-inclusion hierarchy of threshold graphs.
-(iii) What we learned from global synchrony of threshold graphs?
-While identifying previously unknown classes of globally synchronizing graphs contributes to the expansion of known results, a more fundamental objective of this line of research is to elucidate the mechanisms that give rise to global synchronization itself.
-It is already known that for certain classes of graphs, global synchronization arises through mechanisms that, in a certain sense, are close to those of complete graphs.
+(4) Understanding the Mechanism Behind Global Synchrony in Threshold Graphs.
+Previous global synchronization results focus on density-based conditions.
+For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [lu2020synchronization].
+In other words, these results require the graph to be sufficiently dense.
 Threshold graphs, however, lie outside this regime.
+Depending on the positions of dominating vertices in the generating sequence, their density can vary from $O(2/n)$ (stars) to $1$ (cliques), and many intermediate densities occur.
 Nevertheless, we prove in this work that they also exhibit global synchronization.
 What deeper insight can we draw from this result?
-Our analysis shows that the synchronization of threshold graphs is instead driven by another mechanism, namely their intrinsic symmetries.
-These symmetries are not merely combinatorial, such as sets of nodes sharing identical neighborhoods, but are geometric in nature.
-We will introduce this notion in detail later, but to give an intuition, these geometric symmetries emerge inductively as the graph is reduced node by node.
-Through this process, we uncover certain sets of nodes that, while not symmetric in the structural sense of graph theory, exhibit geometric symmetry at the second-order critical points of (3), which allows us to show that these nodes form a synchronized group.
-As the reduction continues, more geometric symmetries are revealed, leading to the formation of progressively larger synchronized groups, until the entire node set forms a single synchronized group, thereby guaranteeing that all second order critical points are synchronized states, thus the graph is global synchronizating.
-1.1. Notation
+Our analysis shows that the global synchronization of threshold graphs is not a consequence of density.
+Instead, it arises from a different mechanism, namely the presence of local symmetries.
+Some of these symmetries come from the graph structure itself: as we show in Lemma 5.2, closed twins must synchronize at any second-order stationary point.
+There are also symmetries that do not originate purely from combinatorial structure, that is, certain configurations exhibit a form of geometric symmetry.
+As we will show in Section 5.3, in the configurations we call synchronous pendants, pairs of vertices that are not structurally symmetric nonetheless become geometrically equivalent at second-order stationary points and hence synchronize.
+Such symmetric configurations arise inductively along the construction process of threshold graphs.
+Their synchronization propagates backward from later-added vertices to earlier ones, which shows that every second-order stationary point is a synchronous state.
+Consequently, every threshold graph is globally synchronizing.
+Notation.
 Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges.
-Furthermore, since the homogeneous Kuramoto model (2) on $G$ is fully determined by the graph topology, we sometimes use system $G$ to denote the Kuramoto model on this graph for simplicity.
 Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$.
+Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle.
+And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$.
+We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise.
+For a vertex $i$ and a subset $Q$ of vertices, we denote
+$N_{Q}(i)=\{\,q\in Q:q\sim i\,\},$
+the set of neighbors of $i$ that lie in $Q$.
 2. Preliminaries of Threshold Graphs
-Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines.
-They were in fact discovered independently and reported in different journals by researchers working in distinct fields.
-The term “threshold graph” was introduced by Chvátal and Hammer in their study of set-packing problems.
-Around the same time, the same class of graphs was independently discovered by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs.
-Since then, threshold graphs have appeared in various contexts, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences.
-Many equivalent conditions have been found for threshold graphs including constructive forms and forbidden configurations. See for example [MP95] for nine different characterizations.
+Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines such as scheduling theory, psychology, computer science [threshold_graphs_book].
+Threshold graphs have been defined by Chvs and Hammer as follows [ChvatalHammer1973, ChvatalHammer1977].
 Definition 2.1.
-A graph is threshold if it contains no induced $P_{4}$, $C_{4}$ and $2K_{2}$.
-One of the many characterizations of threshold graphs is constructive.
-The construction is in terms of two basic operations; adding a disjoint vertex (adjacent to no existing vertex), or a dominating vertex (adjacent to all existing vertices).
-The construction process can be encoded into a binary sequence, called threshold sequence, that is $(b_{1},\dots,b_{n})\in\{0,1\}^{n}$ such that starting from a single vertex, for each $b_{i}=0$ we add an isolated vertex, and for each $b_{i}=1$ we add a dominating vertex; see Figure 1.
+A graph $G$ is called a threshold graph if it can be constructed from the one-vertex graph by repeatedly applying one of the following two operations:
+  1. (i)
+    add an isolated vertex, adjacent to no existing vertex;
+  2. (ii)
+    add a dominating vertex, adjacent to all existing vertices.
+The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating.
+Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [threshold_graphs_book].
+Theorem 2.2.
+Let $G=(V,E)$ be a graph. The following statements are equivalent:
+  1. (i)
+    $G$ is a threshold graph.
+  2. (ii)
+    $G$ is obtained by iteratively adding isolated or dominating vertices to the one-vertex graph.
+  3. (iii)
+    $G$ admits a weight function $w:V\to\mathbb{R}$ and a threshold $t\in\mathbb{R}$ such that
+    $\{u,v\}\in E\;\Longleftrightarrow\;w(u)+w(v)\geq t,\qquad\forall\,u\neq v.$
+  4. (iv)
+    $G$ is $\{P_{4},C_{4},2K_{2}\}$-free.
+  5. (vi)
+    For any distinct $u,v\in V$, $N(u)\subseteq N[v]\quad\text{or}\quad N(v)\subseteq N[u].$
+Remark 2.3 (Edge densities of threshold graphs).
+The edge density of threshold graphs ranges from $O(1/n)$ to $1$. Indeed, if $(b_{2},\dots,b_{n})$ is the generating sequence, then
+$|E|=\sum_{i:b_{i}=1}(i-1),\qquad\mathrm{dens}(G)=\frac{|E|}{\binom{n}{2}}$.
+By choosing the positions at which $b_{i}=1$, one obtains connected threshold graphs whose densities range from $2/n$ (stars) to $1$ (the complete graph), spanning many intermediate values.
+Moreover, threshold graphs are exactly the graphs whose degree sequences saturate the Erdős–Gallai inequalities in the characteristic Ferrers pattern corresponding to nested neighborhoods [threshold_graphs_book].
 11223344
@@
 (c) $2K_{2}$
-Figure 2. Forbidden induced subgraphs for threshold graphs: graphs are $\{P_{4},C_{4},2K_{2}\}$-free. Only edges within each induced pattern are drawn.
+Figure 3. Forbidden induced subgraphs for threshold graphs.
 Remark 2.4.
-Interestingly, threshold graphs have also been studied in synchronization contexts beyond oscillator models.
-In particular, [PS95] surveyed their use in the design of synchronization protocols for cooperating processes, focusing on mutual exclusion via semaphore-based control mechanisms and the so-called PV-chunk semaphores.
-In some papers, these are also referred to as Pvc-definable graphs.
-3. Preliminaries on Kuramoto synchronization
-In this section, we aim to establish the correspondence between dynamical and optimization concepts in the homogeneous Kuramoto model.
-Since the homogeneous Kuramoto model (2) can be formulated as a gradient dynamical system with the energy function (3) and the real analycity of (3), the dynamical notions of equilibrium points and stable equilibria naturally correspond to the optimization notions of first-order and second-order stationary points of the energy function.
-Before showing this correspondence, we first provide the formal definitions of (stable) equilibria in dynamical systems and of first- and second-order stationary points in optimization.
-It is worth noting that such a one-to-one correspondence between the two sets of concepts is rather uncommon.
-This special structure is precisely what allows the analysis of global synchronization in the Kuramoto model to be directly connected to the geometry of the nonconvex optimization landscape.
-In general, the evolution of a dynamical system can often be viewed as a descent process of a potential, known as a Lyapunov function, which monotonically decreases along the system trajectories.
-However, Lyapunov functions are not unique in general, let alone the existence of a clear correspondence between the two classes of concepts discussed above.
-Moreover, even if the system is of gradient type so that its energy function serves as a natural candidate for a Lyapunov function, such a correspondence is still not guaranteed in general.
-The remarkable property of the homogeneous Kuramoto model considered here is that its energy function is real analytic, and it is precisely this property that ensures the correspondence between dynamical and optimization notions.
-For further discussion and proofs, we refer the reader to [AK06].
+Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered.
+To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [Ausiello1977], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs.
+This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [ChvatalHammer1977].
+3. Preliminaries on Kuramoto Model
+In this section, we introduce two families of concepts for the Kuramoto model: one arising from dynamical systems and the other from nonconvex optimization.
+Presenting these notions side by side is important, as many of the local synchronization lemmas developed later rely on interpreting stability to infer the relative positions of phasors at second-order stationary points of the energy landscape.
 3.1. Equilibrium and Local Stability
 In this section, we introduce the notion of equilibrium and stable equilibrium of the Kuramoto model (2).
-By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02].
+By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [khalil2002nonlinear].
 Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
 $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (5)
-An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02].
+An equilibrium point is considered stable if small perturbations around it remain small over time [khalil2002nonlinear].
 The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@
 A point satisfying the above two conditions is also called a second-order stationary point (SOSP) of (4).
-3.3. Correspondence between stability and geometry
-In this section, we briefly outline the relationship between equilibrium stability in the dynamical system (2) and the geometric structure of the energy function (3).
+3.3. Correspondence between Stability and Optimality
+In this section, we briefly outline the relationship between the equilibrium stability of the dynamical system (2) and the local optimality of the energy function (4).
 Understanding this correspondence provides valuable intuition.
-In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order critical point.
-At the first-order level, the two notions coincide: a configuration $\boldsymbol{\theta}$ is a critical point of the energy function $E(\boldsymbol{\theta})$ if and only if it is an equilibrium of the dynamical system (2), since both are characterized by
-$\sum_{j=1}^{n}A_{ij}\sin(\theta_{i}-\theta_{j})=0\text{ for all }i\in[n]$
-One can further relate the stability of an equilibrium to second-order conditions of the energy landscape.
-It is precisely because of the real analycity of the energy function (3), the local asymptotic stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (3).
-This is a remarkable property since it does not hold in general for dynamical systems.
-In general, the dynamics of a system may admit more complex behaviors near equilibria such as limit cycles, or chaotic dynamics, etc.
+In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order stationary point.
+At the first-order level, the two notions coincide.
+Indeed, a configuration $\boldsymbol{\theta}$ is a first-order stationary point of the energy function (4) if and only if it is an equilibrium of the dynamical system (2) on the graph $G$ with adjacency matrix $A$, since both are characterized by
+$\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i})=0\qquad\text{for all }i\in V.$
+One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape.
+It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4).
 The relation mentioned here is summarized in Figure 4.
-One can find the proof in the celebrated work [AK06].
-SOSP of 3 Local minima of 3 = stable equilibria of 2
-Figure 3. Relation between different concepts. Here equivalence between stable equilibrium and local minimum is established in [AK06].
-3.4. Synchronous and aligned states
-In this section, we prove that all aligned states are equilibria.
-Among them, however, the non-synchronous ones are not second-order stationary points (SOSPs).
-Before that, we first define the notions of synchronous states and aligned states.
-Definition 3.6.
-A state $(\theta_{1},\cdots,\theta_{n})$ of $G=(V,E)$ is called a synchronous state if for all $i,j\in V$,
-$\theta_{i}\equiv\theta_{j}\pmod{2\pi}$.
-More generally, it is called an aligned state if for all $i,j\in V$,
-$\theta_{i}\equiv\theta_{j}\pmod{\pi}$.
-Clearly, every synchronous state is also an aligned state. The stability of all aligned states is:
-Theorem 3.7.
-Any aligned state is an equilibrium of the Kuramoto model (2) on the graph $G$ or equivalently stationary point of $E(\boldsymbol{\theta})$.
-Among them, synchronous states are second-order critical point of (3) [^2] In fact, they are asymptotically stable equilibria up to rotational invariance, while non-synchronous aligned states are not second-order critical point of (3).
-Proof.
-Firstly, it is straightforward to see that aligned states satisfy the condition (4), therefore they are equilibria of $G$, or equivalently stationary point of $E(\boldsymbol{\theta}$.
-Secondly, we turn to the second-order criticality of synchronous states. The Hessian matrix at synchronous states is given by
-$\boldsymbol{\mathbf{H}}_{ij}=\begin{cases}\sum_{j=1}^{n}\boldsymbol{A}_{ij},&\text{if }i=j,\\[8.0pt] -\boldsymbol{A}_{ij},&\text{if }i\neq j,\end{cases}$ (6)
-[…]
-Thirdly, regarding the second-order criticality of non-synchronous aligned states.
-Without loss of generality [^3] Any aligned state can be rotated uniformly on all angles $\theta_{1},\cdots,\theta_{n}$, such that $\theta_{i}=0$ or $\pi$, for any $i\in[n]$., assume $\theta_{i}=0$ or $\pi$, for any $i\in[n]$.
-Let $\boldsymbol{x}=(x_{1},\cdots,x_{n})$ be an $n$-dimensional vector that indicates which group (either at angle 0 or $\pi$) each node belongs to
-[…]
-Then the Hessian matrix at any antipodal state is not negative semidefiniteness. Indeed, this can be inferred by the quadratic form
-$\boldsymbol{x}^{T}\boldsymbol{\mathbf{H}}\boldsymbol{x}=\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=-\sum_{i|{\theta_{i}=0}}\sum_{j|{\theta_{j}=\pi}}\boldsymbol{A}_{ij}<0$
-since $\boldsymbol{A}_{ij}\geq 0$, and there must exist at least one pair of nodes $i,j$ such that $\theta_{x}\equiv\theta_{y}\pmod{\pi}$, $k\in\mathbb{Z}$, given the connectness of $G$.
-Therefore, any non-synchronous aligned state is not a SOSP.
-∎
-Remark 3.8.
-In fact, synchronous states are asymptotically stable equilibria up to a global rotation.
-Here, asymptotic stability is a stronger notion than mere stability: it requires that small perturbations lead to trajectories returning to the equilibrium.
-On the other hand, non-synchronous aligned states are not SOSPs and therefore cannot be local minima of 3 or say not stable equilibrium of 2.
-The instability of non-synchronous states can be intuitively understood as follows: a set of antipodally positioned oscillators must evolve in a way that decreases the system’s energy, as there always exists an interaction force between these antipodal pairs that drives them closer together.
-$A_{12}1122$
-Figure 4. Synchronous (left) and non-synchronous aligned state (right). Arrows indicate the nodes moving closer along the circle, implying this state is not stable.
-4. Combinatorial planar geometry of equilibrium
-As we will introduce later, the key idea of our proof is to assume that a configuration $\boldsymbol{\theta}$ is a second-order critical point and show that this assumption forces $\boldsymbol{\theta}$ to be synchronized, meaning that all node vectors coincide.
-This requires us to understand the geometric configuration of second-order critical points on the unit circle.
-In the previous section, we presented the algebraic conditions characterizing first- and second-order critical points.
-In this section, we employ a complex-number formulation to transform these algebraic conditions into geometric ones, i.e., combinatorial relations between the vectors associated with the nodes.
-This seemingly simple transformation significantly simplifies the original landscape geometry problem, reducing it to a combinatorial-geometric problem involving vectors on the two-dimensional plane.
-With this, one just needs to understand whether, under the geometric form of the first- and second-order conditions, it can be deduced that all vectors must be identical.
-4.1. Vector-labelled graph
-Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$, we associate to each node $i\in V$ its phasor
+Such an equivalence was established in the celebrated work [Absil2006StableEquilibrium].
+SOSP of 4 Local minima of 4 = stable equilibria of 2
+Figure 4. Relation between different concepts.
+4. Vector Geometry of Equilibria
+In this section, we recast the algebraic first- and second-order stationary conditions in Section 3 into geometric ones.
+This geometric viewpoint replaces the nonlinear equations (6) and the Hessian constraints by combinatorial relations among the vectors associated with the nodes.
+They form the starting point of our analysis.
+Additionally, stronger geometric constraints will be derived in later sections and combined with the results here to rule out all non-synchronous second-order stationary points of the energy function $E_{G}(\boldsymbol{\theta})$ where $G$ is any connected threshold graph.
+4.1. Vector-labeled Graph
+The basic object underlying this geometric perspective is the vector-labeled graph corresponding to angle configuration of $\boldsymbol{\theta}$.
+Defining such object allows us reasoning about whether a given state $\boldsymbol{\theta}$ is an equilibrium, and whether an equilibrium is stable, i.e., can we move the angles so as to decrease the energy.
+Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
 $\boldsymbol{v}_{i}:=(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}.$
-This yields a vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ which encodes the phase configuration $\boldsymbol{\theta}$ geometrically.
-$v_{i}$$\theta_{i}$
-Figure 5. Representation of $v_{i}$ with angle $\theta_{i}$.
-Although straightforward to construct, this object captures both the equilibrium condition and its stability in the Kuramoto model (2) on the graph $G$.
-More precisely, to determine whether a given $\boldsymbol{\theta}$ is an equilibrium and whether it is stable (or more general, second-order critical point), the combinatorial geometry of the vector-labeled graph $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$ provides a geometric view to algebraic one (solving nonlinear equations and analyzing Hessian at stationary point), as detailed in Section 4.2.
+This is called a phasor, following the standard convention in the Kuramoto model.
+The object $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$, called vector-labeled graph encoding $\boldsymbol{\theta}$ combinatorially and geometrically.
+$\boldsymbol{v}_{i}$$\theta_{i}$
+Figure 5. Representation of $\boldsymbol{v}_{i}$ with angle $\theta_{i}$.
 4.2. Alignment and Equilibrium
 Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms.
-For each node $i\in V$, we aggregate the phasors of its neighbors into a single vector:
-$\sum_{j\in N(i)}\boldsymbol{v}_{j}.$
-This vector represents the aggregated state of node $i$’s neighbors.
-If it is not aligned with $\boldsymbol{v}_{i}$, then intuitively, due to the intrinsic synchronization mechanism in the Kuramoto model, the nodes will tend to move toward alignment instead of staying there at rest, therefore not in equilibrium.
-The equivalence between the alignment $\sum_{j\in N(i)}\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}$ for every node $i\in V$ and the equilibrium condition can be rigorously established by representing the phasors as complex numbers.
 Lemma 4.1 (Equilibrium condition rephrased geometrically).
@@
 $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}\quad\text{where}\quad\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}).$ (7)
+The proof can be found in [monzon2005global], we present here for completeness.
 Proof of Lemma 4.1.
@@
 In this sense, $\mu_{i}$ measures how strongly the phasors of node $i$’s neighbors align directionally.
-Indeed, defining the local order parameter
-$R_{i}:=\frac{1}{\deg(i)}\left|\sum_{j\in N(i)}e^{i\theta_{j}}\right|,$
-we have the relation $\mu_{i}=\deg(i)\cdot R_{i}$, showing that $\mu_{i}$ is proportional to the standard notion of local phase coherence.
+Indeed, defining the local order parameter $R_{i}:=\frac{1}{\deg(i)}\left|\sum_{j\in N(i)}e^{i\theta_{j}}\right|,$ we have the relation $\mu_{i}=\deg(i)\cdot R_{i}$, showing that $\mu_{i}$ is proportional to the standard notion of local phase coherence.
 Remark 4.3 (Equilibrium and force).
@@
 Due to the unit circle geometry $\left|\boldsymbol{v}_{i}\right|=1$ of the Kuramoto model, such a scaling force does not induce any motion while a tangential component can cause rotation.
-Notation.
-(1) We will call $\boldsymbol{\theta}$ is an equilibrium of $G$ receives the force $\boldsymbol{\mu}=(\mu_{1},\cdots,\mu_{n})$, if the vetor-labaled graph representation of $\boldsymbol{\theta}$ satisfies (4.1).
-Note that, in such case, $\mu_{i}=\sum_{j}A_{ij}\cos(\theta_{j}-\theta_{i})$, for all $i\in V$; (2) At state $\boldsymbol{\theta}$, we say that node $i\in V$ is at equilibrium at $\boldsymbol{\theta}$ with respect to $G$ if there exists a scalar $\mu_{i}$ such that (8) holds.
-Clearly, if $\boldsymbol{\theta}$ is an equilibrium, then every node $i\in V$ is at equilibrium.
-In particular, we call a node $i$ is positively (or negatively, or zero) forced, if $\mu_{i}>0$ (or <0<0, or =0=0).
-4.3. Force and stability
-Having established the correspondence between vector alignment at each node and the equilibrium condition, we now turn to the question of stability.
-Alignment alone does not guarantee stability.
-Intuitively, if there exists a node $i$ such that the angle between $\boldsymbol{v}_{i}$ and the resultant vector $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ is zero, a small perturbation will cause the system to return to equilibrium, much like a ball resting at the bottom of a potential valley.
-In contrast, if the angle is $\pi$, the system is analogous to a ball positioned at the peak of a hill: although it is momentarily at rest, a small perturbation will cause the ball to roll away.
-In this section, this intuition is formally confirmed by using the relationship between the negative semi-definiteness of the Jacobian and its diagonal entries, as well as the connection between the stability of equilibrium in gradient dynamical systems and the local geometry of the corresponding energy function.
+(A)(B)(C)(D)
 Figure 6.
-Illustration of different stability conditions for node $i$.
-Here, the solid arrow represents $v_{i}$, while the dashed arrow represents $\sum_{j\in N(i)}v_{j}$.
-The gray region indicates that, in order for node $i$ to attain equilibrium under non-negative forcing, the endpoint of $\sum_{j\in N(i)}v_{j}$ must lie within this sector.
-(A) Node $i$ is not at equilibrium since (8) does not hold; (B) Node $i$ is at equilibrium but negatively forced, and thus unstable.
-Therefore, $\boldsymbol{\theta}$ cannot be a second-order critical point.
-In cases (C) and (D), node $i$ is at equilibrium, and is zero forced and positively forced respectively.
-Thus, satisfies the necessary condition for $\boldsymbol{\theta}$ to be a second-order critical point.
+Geometric intuition for the first- and second-order stationary conditions.
+The solid arrow represents $\boldsymbol{v}_{i}$, and the dashed arrow represents $\sum_{j\in N(i)}\boldsymbol{v}_{j}$.
+For each node $i$, the four panels show all possible configurations satisfying the first-order condition.
+Among them, cases (A) and (B) violate the second-order condition because the aggregated phasor lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
+4.3. Angle and Stability
+Given an equilibrium $\boldsymbol{\theta}$, if there exists a node $i\in V$ whose corresponding force satisfies $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point.
+Indeed, a perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
 Lemma 4.4.
-Assume $\boldsymbol{\theta}$ is an equilibrium state of system $G=(V,E)$ or equivalently first order stationary point of (3).
-In other words, there exist real numbers $\mu_{1},\dots,\mu_{n}$ such that for all $i\in V$,
-$\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
-If $\boldsymbol{\theta}$ is a second order stationary point of (3), then all $\mu_{i}$ must be non-negative.
+If $\boldsymbol{\theta}$ is a second-order stationary point of (4), then either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ or
+$\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0.$
+Proof of Lemma 4.4.
+The $i$-th diagonal entry of $\mathrm{Hess}\,E(\theta)$ can be written as
+$\mathrm{Hess}\,E(\boldsymbol{\theta})_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
+where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form.
+If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bhatia2007] [HornJohnson1985].
+∎
 Remark 4.5 (Mechanical insight into node at stable equilibrium).
@@
 Figure 7.
-$\boldsymbol{v}_{i}$ is pinned down at zero angle, which determines the region where $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}$ should be (the right-end unrestricted region).
-The dashed vectors represent the phasors of nodes connected to ii.
-Proof of Lemma 4.4.
-We begin by noting that the $i$-th diagonal entry of the Hessian matrix $\mathbf{H}$ at $(\theta_{1},\cdots,\theta_{n})$ can be rewritten as
-$\mathbf{H}_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
-where the second equality uses the equilibrium condition (4.1), and the last equality is deduced from re-writing the unit vectors in condition $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}$ into the form using complex numbers.
-If there exists $\mu_{i}<0$, then equivalently $\mathbf{H}_{ii}<0$, which contradicts the positive definiteness of $\mathbf{H}$, see [Bha07] [HJ85].
-Hence, $\boldsymbol{\theta}$ cannot be a second-order critical point.
-Consequently, it is neither a local minimum of (3) nor a stable equilibrium of (2).
-Therefore, $\mathbf{H}_{ii}=\mu_{i}>0$ holds for all $i\in[n]$, where $n$ is the size of $G$.
-∎
-To summarize, the lemma above shows that at any second-order critical point $\boldsymbol{\theta}$, the phasor of each node must be positively aligned with the sum of the phasors of its neighbors (whenever the latter is nonzero).
-That is,
-$\boldsymbol{v}_{i}\;\Uparrow\;\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{j=1}^{n}A_{ij}\boldsymbol{v}_{j}=\boldsymbol{0},$
-for all $i\in[n]$.
-Notation.
-Here, We write $a\Uparrow b$, if both are nonzero and point in the same direction (i.e., the angle between them is zero).
-In the remainder of this manuscript, we will use this notation, as we find it more intuitive.
-Remark 4.6 (This is only a necessary condition).
-Conversely, at a critical point (equilibrium), all $\mu_{i}>0$ for all $i\in V$ is not sufficient to guarantee that $\boldsymbol{\theta}$ is a second-order critical point.
-As a counter-example, the equilibrium $\boldsymbol{\theta}^{*}=(0,0,0,0,\pi,\pi,\pi,\pi)$ of the Kuramoto model on the graph in Figure 8 is a SOSP.
-12345678
-Figure 8.
-At the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$, each node experiences a force from its neighbors that is aligned with its own direction.
-However, this configuration is not stable, as can be verified by examining the eigenvalues of the corresponding Hessian matrix.
-Indeed, the configuration $(0,0,0,0,\pi,\pi,\pi,\pi)$ is not a local minimum of the energy function.
-When nodes $1,2,3,4$ are jointly perturbed slightly away from angle 0 in the same direction, they do not return to the original configuration.
-Instead, they continue to move toward the $\pi$-oriented nodes $5,6,7,8$, indicating a decrease in energy and confirming that the configuration is unstable.
-This can be rigorously proven by quadratic form $\boldsymbol{x}^{T}\boldsymbol{H}\boldsymbol{x}$, we omit it here.
-4.4. Notations
-The below definitions are nodewise concept since they are always made with respect to a node in certain state $\boldsymbol{\theta}$.
-For example, when we say that node $i$ attains equilibrium at $\boldsymbol{\theta}$, we mean that
-$\mu_{i}\boldsymbol{v}_{i}\;=\;\sum_{j\in N(i)}\boldsymbol{v}_{j},\qquad\mu_{i}\in\mathbb{R}.$
-Similarly, when we say that node $i$ attains a stable equilibrium at $\theta$, we mean the same balance equation (8) holds with $\mu_{i}\geq 0$.
-In the proof of global synchrony for threshold graphs, we frequently use this necessary condition.
-To make the reasoning more intuitive for the reader, we assign different colors to nodes representing different cases.
-Condition on $i$ Equation Illustration
-Equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\in\mathbb{R}$ $i$
-Stable equilibrium $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}\geq 0$ $i$
-Asymptotically stable (positively forced) $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}>0$ $i$
-Zero-forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}=0$ $i$
-Negatively forced $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j},\;\;\mu_{i}<0$ $i$
-Table 1. Stability classification of node $i$ at equilibrium $\theta$. The sign of $\mu_{i}$ determines the type of stability, visualized by node color.
-It is important to note that these definitions do not require the state $\boldsymbol{\theta}$ itself to be an equilibrium or stable equilibrium of the system.
-Indeed, at a given state $\boldsymbol{\theta}$ it is possible that some nodes satisfy the condition that their own phasor is parallel to the aggregated phasor formed by their neighbors.
-Such nodes are in equilibrium at this particular configuration $\boldsymbol{\theta}$; however, this does not imply that the configuration itself is an equilibrium state, since other nodes may not satisfy the same condition.
-Nevertheless, if $\boldsymbol{\theta}$ is indeed such an equilibrium, then the above conditions are of course implied, according to Lemmas 4.4 and 4.1.
-5. Geometric twins and their phasor geometry
-In this section, we introduce the concept of geometric twins (g-twins for abbreviation), that takes into account the unit vectors $\boldsymbol{v}_{i}=(\cos\theta_{i},\sin\theta_{i})$ associated with each node $i$ in a Kuramoto state $\boldsymbol{\theta}$.
-Two nodes are said to be g-twins if they satisfy certain geometric relationships at this configuration.
-It is worth noting that, due to the conceptual similarity, we borrow the term twins from graph theory.
-However, we precede it with the word geometric to emphasize that our definition depends on the positions of the node phasors in the specific configuration.
-This concept captures a form of geometric symmetry at $\boldsymbol{\theta}$ that will be crucial in our proof for global synchrony of threshold graphs.
-5.1. Closed geometric twins
-Definition 5.1 (Closed g-twins).
-Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$.
-Two nodes $i,j\in V$ are called closed g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
-$\bigl(\boldsymbol{v}_{i}\parallel\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\parallel\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
-or equivalently
-$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}$
-where $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
-If moreover both strengths are nonnegative, meaning that each vector $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ points in the same direction as the sum of the other and the common vector $\boldsymbol{q}$ (or that summation is the zero vector), then we call them stable closed g-twins.
-Theorem 5.2 (Phasor geometry of closed g-twins).
-Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of closed g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector falls into one of the following three cases:
+With $\boldsymbol{v}_{i}$ fixed at angle 0, the endpoint of the aggregated phasor $\sum_{j}A_{ij}\boldsymbol{v}_{j}$ is constrained to the right-end sliding track.
+The rods may freely adjust their orientations, but the endpoint of the last rod must lie in the feasible sliding interval.
+The solid vectors show the unit-length phasors of the neighbors of node ii.
+The geometric conditions of second-order stationary points presented here are only the most basic one.
+Further necessary geometric conditions will be established in subsequent sections, and these will be essential for showing that every connected threshold graph is globally synchronizing.
+5. Local Synchronization Primitives
+5.1. A Fundamental Geometric Fact
+We begin with a geometric lemma that characterizes all possible relative positions of two unit vectors satisfying a pair of linear relations.
+This fact will serve as the local synchronization primitive used in the induction argument in Section 6.
+Lemma 5.1.
+Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}\subset\mathbb{R}^{2}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
+$\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\qquad\text{and}\qquad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}.$ (12)
+Then the positions of $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$ fall into one of the following three cases:
   1. (1)
@@
   2. (2)
-    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$ and $\mu_{a}+\mu_{b}=-2$ and $\mu_{a}\neq-1$;
+    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$, $\mu_{a}+\mu_{b}=-2$, and $(\mu_{a},\mu_{b})\neq(-1,-1)$;
   3. (3)
-    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=-1$ and $\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=0$.
-$\mu_{a}$$\mu_{b}$
-Figure 9.
-Feasible region of strengths $(\mu_{a},\mu_{b})$ for closed g-twins.
-It lies on the lines $\mu_{a}=\mu_{b}$ (black, with $a$ and $b$ synchronized) and $\mu_{a}+\mu_{b}=-2$ (red, with $a$ and $b$ in antipodal position).
-At their intersection (1,1)(-1,-1), marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
+    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$, $\mu_{a}=\mu_{b}=-1$, and $\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=0$.
 Proof.
-According to the definition of closed g-twins
-$\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\quad\text{and}\quad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$
-we have
+According to (12), we have
 $\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}-\boldsymbol{v}_{b}=\mu_{b}\boldsymbol{v}_{b}-\boldsymbol{v}_{a}$
@@
 ∎
-Definition 5.3 (Stable closed g-twins).
-Under the setting of Definition 5.1, we say that nodes $i$ and $j$ are stable closed g-twins if furthermore
-$\bigl(\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}_{j}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{j}+\boldsymbol{q}=\boldsymbol{0}\bigr)\\ \quad\text{and}\quad\bigl(\boldsymbol{v}_{j}\Uparrow\boldsymbol{v}_{i}+\boldsymbol{q}\;\text{or}\;\boldsymbol{v}_{i}+\boldsymbol{q}=\boldsymbol{0}\bigr)$
-or equivalently
-$\boldsymbol{v}_{j}+\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}\quad\text{and}\quad\boldsymbol{v}_{i}+\boldsymbol{q}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
-Similarly, one can define open g-twins and stable open g-twins at state $\boldsymbol{\theta}$, though this concept will not be used in our proofs in this chapter.
-It may nevertheless be of independent interest in understanding the geometry of non-adjacent nodes with certain geometric symmetries at equilibria.
-5.2. Open geometric twins
-Definition 5.4 (Open g-twins).
-Let $G=(V,E)$ be a graph and $\boldsymbol{\theta}$ a state of the Kuramoto model on $G$.
-Two nodes $i,j\in V$ are called open g-twins at state $\boldsymbol{\theta}$ if there exists a two-dimensional vector $\boldsymbol{q}$ such that
-$\left(\boldsymbol{v}_{i}\parallel\boldsymbol{q}\quad\text{and}\quad\boldsymbol{v}_{j}\parallel\boldsymbol{q}\right)\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
-or equivalently,
-$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\in\mathbb{R}.$
-Here $\boldsymbol{v}_{k}=(\cos\theta_{k},\sin\theta_{k})$ is the unit vector associated with node $k$.
-Definition 5.5 (Stable open g-twins).
-Under the setting of Definition 5.4, we say that nodes $i$ and $j$ form geometric stable open twins if furthermore
-$\boldsymbol{v}_{i}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
-and
-$\boldsymbol{v}_{j}\Uparrow\boldsymbol{q}\quad\text{or}\quad\boldsymbol{q}=\boldsymbol{0},$
-or equivalently,
-$\boldsymbol{q}=\mu_{i}\boldsymbol{v}_{i}=\mu_{j}\boldsymbol{v}_{j}\quad\text{for some }\mu_{i},\mu_{j}\geq 0.$
-Remark 5.6.
-For brevity, whenever the conditions of Definition 5.4 are satisfied, we shall simply say that $i$ and $j$ form open g-twins with strengths $(\mu_{i},\mu_{j})$, with respect to the common vector $\boldsymbol{q}$.
-The terminology for closed g-twins, stable g-twins, and stable closed g-twins is used analogously.
-Under the special geometric symmetries described above, the phasors of g-twins exhibit a particular symmetric configuration on the unit circle.
-This in turn allows us to obtain a precise characterization of the relative positions of the phases $\theta_{i}$ of nodes with such hidden geometric symmetry, both at equilibrium and at stable equilibrium of the Kuramoto model.
-We refer to the notation on node at (stable) equilibrium in Table 1 for clarity.
-Theorem 5.7 (Phasor geometry of open g-twins).
-Consider a state $\boldsymbol{\theta}$ of the Kuramoto model on graph $G=(V,E)$, the phasor positions of open g-twins $a,b\in V$ with strength $\mu_{a},\mu_{b}$ with respect to the common vector $\boldsymbol{q}$ falls into one of the following three cases:
-  1. (1)
-    $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$ and $\mu_{a}=\mu_{b}\neq 0$;
-  2. (2)
-    $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$ and $\mu_{a}=-\mu_{b}\neq 0$;
-  3. (3)
-    $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$ and $\mu_{a}=\mu_{b}=0$.
 $\mu_{a}$$\mu_{b}$
 Figure 8.
-Feasible region of strengths $(\mu_{a},\mu_{b})$ for open g-twins.
-It lies on the lines $\mu_{a}=\mu_{b}$ (black, corresponding to $a$ and $b$ synchronized) and $\mu_{a}=-\mu_{b}$ (red, corresponding to $a$ and $b$ antipodal).
-At their intersection $(0,0)$, marked by a hollow dot, there is no additional constraint on the phasors of $a$ and $b$.
+The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; and (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
+Corollary 5.2.
+Under the assumptions of Lemma 5.1, if in addition $\mu_{a},\mu_{b}\geq 0$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$.
 Proof.
-Since
-$\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b},$ (14)
-the following cases arise.
-Case 1. $\mu_{a}=\mu_{b}=0$, then we have $\boldsymbol{q}=0$.
-Case 2. If $\mu_{a},\mu_{b}\neq 0$, then, from 14 we have
-$\boldsymbol{v}_{a}=\frac{1}{\mu_{a}}\boldsymbol{q}=\frac{\mu_{b}}{\mu_{a}}\boldsymbol{v}_{b}.$
-Since $\|\boldsymbol{v}_{a}\|=\|\boldsymbol{v}_{b}\|=1$, it follows that $|\mu_{a}|=|\mu_{b}|$.
-We distinguish two subcases.
-If $\mu_{a}=\mu_{b}$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$.
-If $\mu_{a}=-\mu_{b}$, then $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$, and one of $\mu_{a},\mu_{b}$ is negative.
+With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
+Definition 5.3 (Geometric twins).
+At state $\boldsymbol{\theta}$ of Kuramoto model on $G$, nodes $i$ and $j$ are called geometric twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the compatibility relations of Lemma 5.1.
+5.2. Phasor Geometry of Closed Twins
+Definition 5.4.
+Let $G=(V,E)$ be a graph. Two nodes $i,j\in V$ are called closed twins if
+$N[i]=N[j],$ (14)
+where $N[i]$ is defined as $N(i)\cup\{i\}$.
+Figure 9. The two black nodes form a pair of closed twins.
+Open twins, defined by $N(i)=N(j)$, will not be needed in this work.
+Corollary 5.5.
+Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$. If nodes $a$ and $b$ form a pair of closed twins, then $\theta_{a}=\theta_{b}$.
+Proof.
+From Lemma 4.4, since $\boldsymbol{\theta}$ is a second-order stationary point, there exist $\mu_{a},\mu_{b}\geq 0$ such that
+$\begin{cases}\displaystyle\sum_{j\in N(a)}\boldsymbol{v}_{j}=\mu_{a}\boldsymbol{v}_{a},\\[4.0pt] \displaystyle\sum_{j\in N(b)}\boldsymbol{v}_{j}=\mu_{b}\boldsymbol{v}_{b}.\end{cases}$
+Because $N[a]=N[b]$, we have $N(a)\setminus\{b\}=N(b)\setminus\{a\},$ and we denote this common set by $M$. Thus,
+$\begin{cases}\displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{b}=\mu_{a}\boldsymbol{v}_{a},\\[4.0pt] \displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b}.\end{cases}$
+By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$, and therefore $\theta_{a}=\theta_{b}$. ∎
+Clearly, this is a case where geometric twins defined in (5.3) formed. Moreover, Corollary 5.5 implies immediately that windmill graphs are globally synchronizing.
+Theorem 5.6.
+Every windmill graph is globally synchronizing.
+Proof.
+Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$, and let $G$ be a windmill graph.
+Within each clique of $G$, all non-central nodes have identical closed neighborhoods and therefore form closed twins.
+By Corollary 5.5, all nodes in each clique must share the same phase.
+Contracting each clique into a single node yields a star graph, which is globally synchronizing.
+Hence all nodes of $G$ have the same phase.
 ∎
-6. Formation of geometric twins
-The motivation for introducing g-twins in the previous section comes from the proof of global synchronization on threshold graphs.
-In that proof, it turns out that synchronization does not rely solely on structural symmetry (i.e., forming structural twins), but can also arise from a more general geometric symmetry, where the corresponding phasors occupy special symmetric positions on the unit circle.
-In this section, we first mention that structural twins at equilibrium constitute the simplest case of forming g-twins.
-We then move on to more general situations where nodes are not structural twins, yet under both specific geometric conditions (on the phasors) and accompanying structural conditions, they still form g-twins.
-Both cases will play a crucial role in the proof of global synchronization for threshold graphs.
-6.1. Structural twins at equilibrium
-Definition 6.1 (Structural Twins).
-Let $G=(V,E)$ be a graph. Two nodes $i,j\in V$ are called structural closed twins if
-$N[i]=N[j],$ (15)
-and structural open twins if
-$N(i)=N(j),$ (16)
-(a) Closed twins ($N[i]=N[j]$).
-(b) Open twins ($N(u)=N(v)$).
-Figure 11. Illustration of structural twins.
-Remark 6.2.
-In the discussion above we used the term structural to emphasize that whether two nodes are twins depends purely on the graph structure.
-In graph theory they are simply called true or adjacent twins, and false or non-adjacent twins.
-Clearly, if $\boldsymbol{\theta}$ is an equilibrium point of the Kuramoto model with respect to graph $G$, and two nodes $a$ and $b$ satisfy condition 15, then they must form a pair of closed g-twins.
-This is because, according to Lemma 4.1, the nodes $a$ and $b$ are structurally symmetric (satisfying (15)) 8, that is
-[…]
-$\boldsymbol{q}=\sum_{j\in N(a)\setminus\{b\}}\boldsymbol{v}_{j}=\sum_{j\in N(b)\setminus\{a\}}\boldsymbol{v}_{j}$
-such that $\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4.
-The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14.
-There, colors indicate node-wise stability: Blue indicates $\mu>0$ (stable), red for $\mu<0$ (unstable), black for $\mu=0$.
-Figure 12. Structural closed twins attaining nodewise equilibrium.
-Remark 6.3.
-Here the case (d) is detailed in the following figure. One can clearly see that
-$\boldsymbol{v}_{a}+\boldsymbol{v}_{b}+\boldsymbol{q}=\boldsymbol{0},$
-where $N$ denotes the set of common neighbors of $a$ and $b$, excluding each other.
-Figure 13.
-Geometric illustration of case (d) for false twins at equilibrium.
-The red arrow represents the vector sum $\boldsymbol{q}$, where $N$ is the common neighborhood of nodes $a$ and $b$, excluding each other.
-In this case, $\mu_{a}=\mu_{b}=-1$, and the three vectors $\boldsymbol{v}_{a}$, $\boldsymbol{v}_{b}$, and $\boldsymbol{q}$ form a triangle summing to zero.
-Similarly, structural open twins $a,b$ at equilibrium in $\boldsymbol{\theta}$ are also geometric open twins, since there exists
-$\boldsymbol{q}=\sum_{j\in N(a)}\boldsymbol{v}_{j}=\sum_{j\in N(b)}\boldsymbol{v}_{j}$
-such that $\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}$ and $\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}$ for some $\mu_{a},\mu_{b}\in\mathbb{R}$, according to notations in 4.4.
-The relative positions of nodes $a$ and $b$ on the unit circle, together with their nodewise stability in this case, are illustrated in Figure 14.
-The same as in the above mentioned closed twins case, that colors indicate their stability.
-Figure 14. Structural open twins attaining nodewise equilibrium.
-Our strategy for proving that a given graph admits global synchronization is to characterize the positions of all nodes at stable equilibrium (in fact, more general, at SOSP of $E(\boldsymbol{\theta})$).
-Furthermore, the lemma on structural twins at equilibrium shows that whenever two structural twins both attain stable equilibrium, they must synchronize if they are closed twins.
-It is worth noting, however, that in the case of open twins, synchronization is not guaranteed: it may happen that both strengths $\mu_{i}$ and $\mu_{j}$ vanish, in which case the phasors of $i$ and $j$ remain unconstrained.
-At this point, we are already able to characterize certain new classes of globally synchronizing graphs, according to results in this section.
-Theorem 6.4.
-Complete split graphs, windmill graphs, complete bipartite graphs are globally synchronizing.
-(a) Complete split graphs.
-(b) Windmill graphs $W_{5,6}$.
-(c) Complete bipartite.
+Figure 10. Windmill graphs $W_{5,6}$.
 5.3. Synchronous Pendant Extension
-In this section, we present a scenario in which a group of nodes forms closed g-twins without being structural closed twins, that is called synchronous pendant extension.
-Specifically, consider a set $S$ that satisfies the structural assumption that the additional neighborhood of some subset $S_{1}\subseteq S$ is attached to $S_{1}$ in a leaf-like manner.
-The geometric assumption, on the other hand, is that this additional neighborhood synchronizes with $S_{1}$.
-To prepare for the presentation and proof of Lemma 6.6, we first establish Lemma 6.5, which highlights the following fact: If there exists a synchronized, fully connected set $S$ of nodes with a synchronized pendant attached, then each node in $S$ is aligned with the vector sum of the phasors of its remaining neighbors, namely those not in $S$ nor in the pendant.
-From the perspective of synchronization propagation, this can be interpreted as synchronization spreading from $S$ together with its pendant to the group of all other neighbors.
+In this section, we present a scenario in which a group of nodes satisfy the phasor relations of Lemma 5.1 without being closed twins.
+More explicitly, the main result of this section is Lemma 5.8, which shows that even when a set of nodes are not twins, they may still be forced to synchronize at any second-order stationary point, provided that their distinct neighbors form a synchronous pendant whose combined phasor contribution collapses to a single, non-adversarial direction.
+The next lemma, Lemma 5.7, serves as a preparatory step toward establishing Lemma 5.8.
 Lemma 5.7.
-Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
-$W=Q\cup S\cup P.$
-under the following structural assumptions:
-(i) For all $i\in Q$, we have $N(i)\subseteq S$;
-(ii) The subgraph $G[S]$ is a complete graph;
-(iii)For any $i\in S$, its neighbourhood consists of the set $P$, its neighbours in $Q$, and the rest of the clique $S\setminus\{i\}$.
-Assume further that all nodes in $Q\cup S$ are synchronized, i.e., there exists $\boldsymbol{v}$ such that $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}=\boldsymbol{v},\quad\forall i,j\in Q\cup S$. Then
-$\sum_{i\in P}\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}\quad\text{or}\quad\sum_{i\in P}\boldsymbol{v}_{i}=\boldsymbol{0}.$
-Figure 16.
-Illustration of Lemma 6.5.
-Elements in the figure follow the same conventions as in Figure 18.
-Moreover, a large node represents a set of nodes.
-An edge from a small node $A_{i}$ to a large node $S$ indicates that $A_{i}$ is connected to all nodes in $S$.
-The diagram depicts a configuration where a set of synchronized “false twin-like” nodes, with differing but synchronized neighbors, implies that their shared neighbors (the large white node) sum to a vector pointing in the same direction as the synchronized nodes.
+Let $G=(V,E)$ be a graph, and let $W\subseteq V$ be partitioned into
+$W=Q\uplus S\uplus P,$
+subject to the following structural conditions:
+  1. (i)
+    Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
+  2. (ii)
+    $S$ induces a clique in $G$;
+  3. (iii)
+    For each $i\in S$, its neighborhood is $P\cup N_{Q}(i)\cup(S\setminus\{i\})$, where $N_{Q}(i)=\{\,q\in Q:q\sim i\,\}$.
+Assume further that the configuration $(\boldsymbol{v}_{i})_{i\in V}$ is a second-order stationary point of the Kuramoto energy (4), and that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$.
+Then the phasor sum over $P$ satisfies
+$\sum_{i\in P}\boldsymbol{v}_{i}\;\Uparrow\;\boldsymbol{v}\qquad\text{or}\qquad\sum_{i\in P}\boldsymbol{v}_{i}=\boldsymbol{0}.$
 Proof.
 For any node $i\in S$,
-$N(i)=Q\cup(S\setminus\{i\})\cup P$ (17)
+$N(i)=Q\uplus(S\setminus\{i\})\uplus P$ (15)
 and
@@
 Plug (15) into (16), we obtain
-$\sum_{j\in Q\cup(S\setminus\{i\})\cup P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
-Since $Q\cup S$ synchronize, we have
-$(|Q|+|S|-1)\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
+$\sum_{j\in Q\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
+Since $Q\uplus S$ forms a synchronized group, we have
+$\bigl(|Q|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
 Thus
-$\sum_{j\in P}\boldsymbol{v}_{j}=\left(\mu_{i}-\left(|Q|+|S|-1\right)\right)\boldsymbol{v}_{i},\quad\text{where}\quad\mu_{i}\geq 0.$
-Now, we are left to show $\mu_{i}-\left(|Q|+|S|-1\right)<0$ is not possible.
-Indeed, let $\boldsymbol{v}$ be an $n$-dimensional vector where the $i$-th element is one if $i\in Q\cup S$, and zero otherwise.
-Then
-$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\cup S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (19)
+$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|Q|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
+We now show that $\mu_{i}-(|Q|+|S|-1)<0$ is impossible, because in this case there would exist a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$.
+To see this, let $\boldsymbol{v}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $v_{i}=1$ for $i\in Q\uplus S$ and $v_{i}=0$ otherwise. Then
+$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\uplus S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (17)
 $=\sum_{i\in S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$
@@
 $=\sum_{i\in S}\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}$
-$=|S|\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}$
-where the second equality is due to there is no connection between $Q$ and $P$.
-Clearly, if $\mu_{i}-\left(|Q|+|S|-1\right)<0$, then the above form is negative, and this contradicts to the stability of $\boldsymbol{\theta}$.
-The proof is complete.
+$=|S|\,\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}.$
+where the second equality uses the fact that there are no edges between $Q$ and $P$.
+Clearly, if $\mu_{i}-(|Q|+|S|-1)<0$, then the above quantity is negative, which contradicts the second-order stationarity of $\boldsymbol{\theta}$.
 ∎
-We are now ready to present the case of a synchronous pendant, which constitutes one possible configuration leading to the formation of geometric stable closed twins.
-This case will serve as one of the core lemmas in our proof for threshold graphs.
-Consider a set of nodes $S$ attaining stable equilibrium at $\boldsymbol{\theta}$ which induces a clique [^4] Forming a clique guarantees the “closed” property required for closed twins..
-$S$ can be divided into $S_{1}$ which connects only to a set $P$, and a synchronized group $S_{2}$ which in addition connects to a synchronized pendant set [^5] Specifically, $Q$ can be understood as a generalization of a leaf attached to one node, and it is well known that the presence of a leaf does not affect the stability of the state.
-$Q$.
-We will show that although $S_{1}$ and $S_{2}$ are not structurally identical, this difference, arising from the synchronous pendant, does not interfere with synchronization, in other words, the two groups behave as geometric stable closed twins.
-The proof of Lemma 6.6 is based on Lemma 6.5. If we can decompose the set $P$ there into $\{A\}\cup P$, where $A$ connects to all nodes in $S$, and all nodes in $P$, then we have the following Lemma.
+Figure 11.
+Illustration of Lemma 5.7.
+The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$.
+The large circular node represents the set $P$ of vertices.
+Gray edges indicate synchronization interactions between the corresponding nodes.
 Lemma 5.8.
 Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
-$W=Q\cup S_{1}\cup S_{2}\cup P.$
+$W=Q\uplus S_{1}\uplus S_{2}\uplus P.$
 under the following structural assumptions:
 (i) For all $i\in Q$, we have $N(i)\subseteq S_{1}$;
-(ii) The subgraph $G[S_{1}\cup S_{2}]$ is a complete graph;
-(iii) For any $i\in S_{1}\cup S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\cup S_{2}$.
+(ii) The subgraph $G[S_{1}\uplus S_{2}]$ is a complete graph;
+(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$.
 Formally, if we denote $Q_{i}=N(i)\cap Q$, then
-$N(i)\;=\;P\cup Q_{i}\cup\bigl((S_{1}\cup S_{2})\setminus\{i\}\bigr).$
+$N(i)\;=\;P\uplus Q_{i}\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
 Equivalently,
-$N(i)\setminus\bigl(Q\cup(S_{1}\cup S_{2})\bigr)\;=\;P.$
-Assume further that all nodes in $Q\cup S_{1}$ are synchronized, i.e.,
-$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\cup S_{1}.$
-Then the set $S_{1}\cup S_{2}$ forms a group of geometric stable closed twins, therefore they synchronize.
-Figure 17.
-Illustration of Lemma 6.6.
-Elements follow the same conventions as in previous figures.
-The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes.
-Each orange arc separates a small blue node from a group of nodes it connects to, indicating that the vector of the blue node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
-The figure illustrates how the synchronized block $A\cup S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
+$N(i)\setminus\bigl(Q\uplus(S_{1}\uplus S_{2})\bigr)\;=\;P.$
+Assume further that all nodes in $Q\uplus S_{1}$ are synchronized, i.e.,
+$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\uplus S_{1}.$
+Then the set $S_{1}\uplus S_{2}$ synchronize.
 Proof.
 Observe that the set $S_{1}$ induces a fully-connected subgraph, and $Q$ is a synchronous pendant to $S_{1}$, then according to Lemma 5.8, we have
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
+$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
 Combine this with the fact
@@
 we obtain
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\cup P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
+$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
 Equivalently, we can write
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (20)
+$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (18)
 Another observation is on nodes in $S_{2}$ that
-$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\cup P}\boldsymbol{v}_{k}=\boldsymbol{0}$
+$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}$
 since every node in $S_{2}$ attains stable equilibrium in $\boldsymbol{\theta}$. Equivalently, we can write
-$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\cup P\cup S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (21)
-Combined the facts 20 and 21, we get nodes in $S_{1}\cup S_{2}$ are geometric stable closed twins. ∎
-7. Main proof: the induction argument
-7.1. One more key lemma
-To complete the argument, we establish one more lemma prior to the proof of the main theorem.
-Lemma 7.1 (Synchronization of closed twins with common neighbour).
-Let $S\subseteq V$ be a set of structural closed twins in a graph $G$, and let among their common neighbourhood, there exists a node $A$ such that its neighbourhood is exactly $S$, i.e., $N(A)=S$.
-Assume that all nodes in $S\cup\{A\}$ attain stable equilibrium at configuration $\boldsymbol{\theta}$, that is,
-$\forall i\in S\cup\{A\},\quad\mu_{i}v_{i}=\sum_{j\in N(i)}v_{j}\quad\text{ with }\quad\mu_{i}\geq 0,$
-then the nodes in $S\cup\{A\}$ synchronize at $\boldsymbol{\theta}$.
+$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (19)
+Combined the facts (18) and (19), we get nodes in $S_{1}\uplus S_{2}$ are geometric stable closed twins. ∎
+Clearly, this is another situation in which geometric twins, as defined in Definition 5.3, appear in addition to closed twins at the second-order stationary point.
+Figure 12.
+Illustration of Lemma 5.8, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
+The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$.
+The gray-shaded nodes are synchronized; an edge with a gray background indicates synchronization between the corresponding nodes.
+Each orange arc separates a small green node from a group of nodes it connects to, indicating that the vector of the green node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
+5.4. Twin Attachment
+If a set of nodes all have identical neighborhoods, they form a group of closed twins.
+When an additional node is connected exclusively to this group, the entire collection of nodes, must take the same phase at any second-order stationary point.
+Lemma 5.9.
+Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$.
+If $S\subseteq V$ is a set of closed twins in $G$ and $A\in V$ satisfies $N(A)=S$, then $\theta_{i}=\theta_{j}$ for all $i,j\in S\cup\{A\}$.
+Proof.
+Since $\boldsymbol{\theta}$ is a second-order stationary point of $E_{G}(\boldsymbol{\theta})$ and $S$ forms a set of closed twins, Lemma 5.5 yields
+$\theta_{i}=\theta_{j}\qquad\text{for all }i,j\in S.$
+Because $N(A)=S$, Lemma 4.4 implies
+$\theta_{A}=\theta_{i}\qquad\text{for all }i\in S.$
+Hence all nodes in $S\cup\{A\}$ synchronize. ∎
 Figure 13.
@@
 The conclusion is that all green nodes must synchronize, indicated by the thick edges in the right diagram.
-Proof.
-Since $S$ forms a set of structural stable closed twins, we have $i,j\in S$, we have $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$ and let us denote this common phasor as $\boldsymbol{v}$.
-Moreover, since $A$ only connects to the set $S$ which is synchronizing, and it attains stable equilibrium at $\boldsymbol{\theta}$, we have
-$\mu_{A}\boldsymbol{v}_{A}=\sum_{j\in N(A)}\boldsymbol{v}_{j}=\sum_{j\in S}\boldsymbol{v}_{j}=|S|\boldsymbol{v}$
-Thus $\boldsymbol{v}_{A}=\boldsymbol{v}$. The proof is complete. ∎
-7.2. Illustrative example
-In this section, we present a simple but general enough example to illustrate how to use Lemma 6.6 and 7.1 to establish global synchronization of connected threshold graphs.
-Specifically, to show how synchronization starts from a local substructure and eventually propagates to the entire threshold graph.
-The proof for all connected threshold graphs follows the same idea, with additional work needed for full generalization.
-Let us take a look at the threshold graph corresponding to the sequence
-$0101$
-We use capital letters to represent the nodes corresponding to each bit in the sequence, including the initial starting node.
-A0B1C0d1E0f1G0h1I
-The corresponding threshold graph is shown as follows.
-Figure 19. Threshold graph corresponding to the sequence $0101$, with vertices labeled AA to $I$ in the order of their addition.
-Concluding global synchronization step by step.
-To show that this graph globally synchronizes, we prove that any second order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ on the threshold graph must be synchronous.
-Clearly, any second order stationary point $\boldsymbol{\theta}$ is also a local minimum of $E_{G}(\boldsymbol{\theta})$ or stable equilibrium of the Kuramoto model 2.
-Thus, the Lemmas established in Chapter 4 applies here.
-That is, every node attains stable equilibrium in $\boldsymbol{\theta}$, according to Lemma 4.4.
-We will use Lemmas 6.6 and 7.1 repeatedly in the following illustration.
-Step 1.
-Start with the edge between $H$ and $I$.
-Synchronization between nodes $H$ and $I$ can be obtained from stability condition of $H$ soely.
-Indeed, since $H$ attains stable equilibrium at $\boldsymbol{\theta}$, there exists $\mu_{H}\geq 0$ such that $\mu_{H}\boldsymbol{v}_{H}=\sum_{j\in N(H)}\boldsymbol{v}_{j}$ and we know that $\sum_{j\in N(H)}\boldsymbol{v}_{j}=\boldsymbol{v}_{I}$, we obtain $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$.
-One could also use Lemma to argue $\boldsymbol{v}_{H}=\boldsymbol{v}_{I}$, as this is the simplest case of the lemma.
-Step 2.
-Observe that the pair of node $G$ and $I$ forms geometric stable closed twins at $\boldsymbol{\theta}$.
-Indeed, the only difference between $N[G]$ and $N[I]$ is that $I$ additionally has a synchronizing pendant $H$.
-Thus $\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Now nodes $G,I,H$ are synchronized.
-Step 3. Note that node $F$ connects to $G$ and $I$ which are synchronized as established in Step 2, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
-Step 4.
-Observe that nodes $E,G,I$ induce a fully connected subgraph, and nodes $G$ and $I$ have additional synchronizing pendant, therefore, according to Lemma 6.6, three of them form geometric stable closed twins.
-Thus, $\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Step 4. Observe that node $D$ connects to $E,G,I$ which are synchronzed, thus $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$, according to Lemma.
-Step 5.
-Observe that nodes $C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $E,G,I$ admit additional synchronzing pendant $D,F,H$.
-Consequently $C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
-Step 6.
-Now note that the node $B$ connects to a group of synchronzing nodes $C,E,G,I$.
-Therefore according to Lemma, node $B$ synchronzes with them as well.
-That is, $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Step 7.
-Observe that nodes $A,C,E,G,I$ induce a fully-connected subgraph, and among them, nodes $C,E,G,I$ admit additional synchronzing pendant $B,D,F,H$.
-Consequently $A,C,E,G,I$ form a group of geometric stable closed twins, and therefore they synchronize.
-In conclusion, $\boldsymbol{\theta}$ is a synchronized state, and thus the threshold graph globally synchronizes.
-7.3. Main proof
-Before proceeding, we introduce some notation.
-Any connected threshold graph can be specified by a bit sequence consisting of alternating blocks of $1$’s and 0’s.
-Since the graph is connected, the last bit must be $1$.
-We group the sequence into maximal contiguous blocks of $1$’s and 0’s.
-Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block directly above $U_{i}$ (if any).
-Thus, the sequence takes the form either
+6. Main Proof: the Induction Argument
+6.1. A Prototype Example
+Let us take a look at the threshold graph corresponding to the sequence 01010101.
+The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence, with each bit determining whether the added vertex is isolated (0) or dominating (1).
+Let $G$ be the threshold graph corresponding to the sequence 01010101.
+To prove that $G$ is globally synchronizing, we show that every second-order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ is synchronous; that is, $\theta_{1}=\cdots=\theta_{n}$.
+The following figure illustrates how synchronization must propagate, step by step, in any second-order stationary configuration on this graph.
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+ACEGIHFDB
+Figure 14. Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous.
+Step (1) We begin with the edge between $H$ and $I$.
+Applying Lemma 4.4 to node $H$ shows that, since $\boldsymbol{\theta}$ is a second-order stationary point, we must have $\boldsymbol{v}_{H}\Uparrow\boldsymbol{v}_{I}$.
+Equivalently, nodes $H$ and $I$ synchronize.
+Step (2) Observe that $N(G)=N(I)\cup H$ and $\theta_{H}=\theta_{I}$, applying Lemma by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain $\theta_{G}=\theta_{I}$.
+Intuitively, the only difference between the neighborhoods of $G$ and $H$ is that $H$ has one additional neighbor, namely $I$.
+This extra neighbor does not obstruct the synchronization between $G$ and $I$, because $H$ is already synchronized with $I$ and therefore behaves as a synchronous pendant at this stage.
+Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2).
+Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
+Step (4) Observe that nodes $F$ and $H$ connect only to $G$ and $I$, and the four-node set $\{F,H,G,I\}$ behaves as a synchronizing group.
+Furthermore, ignoring the leaf-like nodes $F$ and $H$ and edges in the induced clique on $\{E,G,I\}$, the two nodes $G$ and $I$ share exactly the same neighbors:
+$N(G)\setminus(\{F\}\cup\{E,I\})\;=\;N(I)\setminus(\{F,H\}\cup\{E,G\})\;=\;N(E)\setminus\{G,I\}\;=\;\{D,B,A,C\}.$
+Therefore, by applying Lemma 5.8 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude that the nodes $E$, $G$, and $I$ form a synchronizing group.
+Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4).
+Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
+Step (6) Observe that nodes $D$, $F$, and $H$ connect only to $E$, $G$, and $I$, and therefore the six-node set $\{D,F,H,E,G,I\}$ behaves as a synchronizing group.
+Furthermore, ignoring the leaf-like nodes $D$, $F$, and $H$ and edges in the induced clique on $\{C,E,G,I\}$, the four nodes $C$, $E$, $G$, and $I$ share exactly the same neighbors, namely $\{A,B\}$.
+Therefore, by applying Lemma 5.8 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude that the nodes $C$, $E$, $G$, and $I$ form a synchronizing group.
+Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6).
+Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
+Step (8) As in Step (7), Lemma 5.9 implies that $A$ synchronizes with $C$, $E$, $G$, and $I$, since these nodes form a synchronizing group and constitute all neighbors of $A$.
+In conclusion, $\boldsymbol{\theta}$ must assign the same angle to every node.
+Equivalently, every second-order stationary point is synchronous, and hence the threshold graph is globally synchronizing.
+6.2. Proof of Theorem 1.2: General Case
+Proof of Theorem 1.2.
+Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, as introduced in Definition 2.1.
+Since the graph is connected, the last bit must be a $1$.
+We decompose the sequence into maximal contiguous blocks of $1$’s and 0’s.
+Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block that lies immediately above $U_{i}$, if such a block exists.
+In this notation, the sequence has one of the two forms
 $U_{1},\,I_{2},\,U_{2},\,\dots,\,I_{k-1},\,U_{k-1},\,I_{k},\,U_{k},$
@@
 $I_{1},\,U_{1},\,I_{2},\,U_{2},\,\dots,\,I_{k-1},\,U_{k-1},\,I_{k},\,U_{k},$
-depending on whether the sequence begins with a 0-block.
-Proof of Theorem 1.2.
-We prove the main theorem 1.2 by induction.
-First, assume $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$.
-Our goal is to show $\theta_{1}=\cdots=\theta_{n}$.
-Let the sequence defining the threshold graph consist of maximal contiguous blocks of 1’s and 0’s, arranged as either:
-$U_{1},I_{1},U_{2},I_{2},\dots,U_{k}\quad\text{or}\quad I_{0},U_{1},I_{1},U_{2},I_{2},\dots,U_{k}.$
-Base case.
-At a stable equilibrium $\boldsymbol{\theta}$, the last block $U_{k}$ must synchronize.
-This follows directly from Lemma 5.2, as the nodes in $U_{k}$ are false twins and thus must synchronize.
-Inductive hypothesis. Assume that the nodes in the last $m$ blocks of 1’s, $U_{k-m+1},\dots,U_{k}$, are synchronized.
-Inductive step.
-Here, we show that the synchronization of $\bigcup_{i=k-m+1}^{k}U_{i}$ propagates to $\bigcup_{i=k-m}^{k}U_{i}$, since the preceding block $U_{k-m}$ must synchronize with $\bigcup_{i=k-m+1}^{k}U_{i}$.
-Indeed, since all nodes in $\bigcup_{i=k-m+1}^{k}U_{i}$ synchronize, any node added as isolated nodes whose neighbourhood is a subset of $\bigcup_{i=k-m+1}^{k}U_{i}$ must synchronize with them as well.
-These nodes are exactly nodes in $\bigcup_{i=k-m+1}^{k}I_{i}$.
-That is, every nodes $i,j\in\bigcup_{i=k-m+1}^{k}U_{i}\cup I_{i}$ satisfy $\boldsymbol{v}_{i}=\boldsymbol{v}_{j}$.
-Observe that $\bigcup_{i=k-m+1}^{k}I_{i}$ are in fact synchronizing pendant attaches to $\bigcup_{i=k-m+1}^{k}U_{i}$, and they are the only difference of neighbourhood between the $U_{k-m}$ and $\bigcup_{i=k-m+1}^{k}U_{i}$, thus according to Lemma 6.6, nodes in $\bigcup_{i=k-m}^{k}U_{i}$ form a set of geometric stable closed twins, and they must synchronize.
-Conclusion.
-All nodes in $\cup_{i=1}^{k}\left(U_{i}\cup I_{i}\right)$ or $I_{0}\cup\left(\cup_{i=1}^{k}(U_{i}\cup I_{i})\right)$ synchronize.
-Namely $\theta_{1}=\cdots=\theta_{n}$, thus any threshold graph is globally synchronizing as long as it is connected.
+depending on whether or not it begins with a block of 0’s. We focus on the first case; the second follows by an entirely analogous argument.
+Given that $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$, we aim to show that $\theta_{1}=\cdots=\theta_{n}$, and we proceed by induction.
+We proceed by induction.
+For the base case, observe that all nodes within a block of $1$’s form a set of closed twins.
+Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Lemma 5.1.
+Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\dots,U_{k}$, are synchronized.
+Lemma 5.9 then implies that the adjacent 0-blocks $I_{k-m+1},\dots,I_{k}$ synchronize with them as well.
+We next show that the preceding $1$-block $U_{k-m}$ also synchronizes with $U_{k-m+1},\dots,U_{k}$.
+Indeed, define
+$Q$ $=I_{k-m+1}\uplus\cdots\uplus I_{k},$
+$S_{1}$ $=U_{k-m+1}\uplus\cdots\uplus U_{k},$
+$S_{2}$ $=U_{k-m},$
+$P$ $=U_{1}\uplus\cdots\uplus U_{k-m-1}\uplus I_{2}\uplus\cdots\uplus I_{k-m}.$
+We next verify that the sets $Q,S_{1},S_{2}$, and $P$ satisfy the assumptions of Lemma 5.8.
+First, for every node $i$ in $Q$, its neighborhood is contained in $S_{1}$; indeed, these nodes are isolated at this stage and connect only to the dominating nodes that are added later.
+Second, the sets $S_{1}$ and $S_{2}$ together induce a clique.
+Third, for any node in $S_{1}\cup S_{2}$, its neighbors consist exactly of all other nodes in $S_{1}\cup S_{2}$, all nodes in $P$, and possibly some nodes in $Q$.
+Hence all conditions of Lemma 5.8 are satisfied, and we conclude that $S_{1}$ and $S_{2}$ must synchronize.
+In particular, the block $U_{k-m}$ synchronizes with $U_{k-m+1}\uplus U_{k-m+2}\uplus\cdots\uplus U_{k}$.
+By induction, all nodes synchronize at $\boldsymbol{\theta}$.
+Therefore, every second-order stationary point is a synchronous state.
+It follows that every connected threshold graph is globally synchronizing.
 ∎
+7. Discussion and Open Problems
+In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$, revealing a new mechanism of global synchronization that has not been articulated before.
+Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes.
+This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.
+A natural question is whether this closure phenomenon extends beyond threshold graphs.
+Specifically, suppose $G$ is an arbitrary globally synchronizing graph.
+If we add a universal node to $G$, is the resulting graph still globally synchronizing?
+We expect the answer to be yes, since attaching a node that is connected to all existing vertices does not appear to introduce new geometric directions for the oscillators beyond the synchronized configuration.
+However, we do not currently know how to prove this, and we leave this problem for future work.
+Moreover, we expect that our results may admit further generalizations.
+Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [Geshkovski2025], and considering broader classes of graphs.
+For instance, one may study systems in which oscillators are coupled through both positive and negative edges.
 Acknowledgements
-We would like to thank Afonso S. Bandeira for an insightful discussion at the early stage of this work.
+We would like to thank Afonso S. Bandeira for insightful discussions.

### v4

@@
 The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph.
-Each edge acts like a spring that pulls the two adjacent oscillators toward each other whenever their phases differ.
-A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions, except for a set of measure zero.
+Each edge acts like a spring that pulls adjacent oscillators toward each other whenever their phases differ.
+A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions.
 This property is tightly linked to the benign nonconvexity of the model’s energy landscape.
-Existing guarantees for global synchronization rely on minimum-degree thresholds, which require the graph to be highly dense.
-In this work, we show that connected threshold graphs, whose density may vary from $2/n$ to $1$, are globally synchronizing.
+Existing guarantees for global synchronization typically require every node to connect to sufficiently many neighbors, which in turn enforces rather homogeneous degree sequences and highly dense graphs.
+In this work, we identify threshold graphs as a class of graphs that lie outside this regime yet nevertheless exhibit global synchronization.
+More precisely, connected threshold graphs realize all admissible edge densities ranging from $2/n$ to $1$, and their degree sequences are maximally left-skewed among all graphs with the same edge density, when ordered in nonincreasing order.
 Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
@@
 Since then, synchronization phenomena have been widely studied across both natural and technological systems.
-For example, the flashing of fireflies [Buck1968, Strogatz1993], the rhythmic beating of heart cells [Winfree1967, Mikhailov2002], synchronization in power grids [dorfler2012synchronization], coordinated motion in robotic networks [OlfatiSaber2006], modern artificial intelligence models such as neural networks [KuramotoNeurons2024], and Transformer architecture [Geshkovski2025, Criscitiello2024].
+For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [WCP+24], and Transformer architecture [GLPR25, CRMB24].
 A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness.
-The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kuramoto1975], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
+The Kuramoto model, introduced by Yoshiki Kuramoto in 1975 [Kur75], provided precisely such a balance, and has since become a central and widely accepted model for analyzing synchronization phenomena.
 The Kuramoto model describes $n$ oscillators on the unit circle, coupled according to a graph $G=(V,E)$ with adjacency matrix $A$. Mathematically, it is a set of ordinary differential equations
@@
 A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized.
-This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [michaels1987sinoatrial], technological applications such as deep brain stimulation [tass2003desynchronizing, nabi2011singleinput] and power systems [chiang1995direct, dorfler2012synchronization].
-For comprehensive surveys of this line of work and its applications, we refer to [strogatz2000, acebron2005, dorfler2014].
+This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [Tas03, NM11] and power systems [CCC95, DB12].
+For comprehensive surveys of this line of work and its applications, we refer to [Str00, ABPV+05, DB14].
 Another line of research focuses on the homogeneous Kuramoto model where all intrinsic frequencies are identical, i.e., $\omega_{i}=\omega$ for all $i\in\{1,\cdots,n\}$.
@@
 This becomes immediate after applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, under which the dynamics reduce to the gradient system
-$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in[n].$ (2)
+$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (2)
 By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$.
@@
 Transforming back to the original coordinates, this implies that the oscillators always achieve frequency synchronization.
-In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [Maistrenko2005].
+In contrast, when intrinsic frequencies are non-identical, the dynamics may exhibit significantly more complex long-term behavior, including chaotic attractors [MPT05].
 In the homogeneous setting, the remaining challenge is to determine whether phase synchronization occurs.
@@
 Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero?
-This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [bandeira2016low, ling2019landscape, ling2023local] and the dynamics of Transformer architectures [Geshkovski2025, geshkovski2023emergence].
+This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin23] and the dynamics of Transformer architectures [GLPR25, GLPR23].
 In this work, we follow this line of research, and beginning with a precise formulation of the problem.
-Problem Setting.
+1.1. Problem Setting
 Definition 1.1 (Globally synchronizing graph).
-We say that a graph $G=(V,E)$ is globally synchronizing if, for almost all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, the solution to (2) satisfies
+We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero, the solution to the Kuramoto model (2) on $G$ satisfies
 $\lim_{t\to\infty}(\theta_{i}(t)-\theta_{j}(t))=0\qquad\text{for all }i,j\in V.$
@@
 Nevertheless, one may form an intuitive if somewhat heuristic expectation: since in a complete graph each oscillator tends to move closer to every other oscillator, the abundance of interactions may indeed drive the system to synchronize from almost all initial configurations.
-Historically, Verwoerd and Mason [verwoerd2007global] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing.
-This conjecture was resolved by Richard Taylor in 2012 [taylor2012there], and global synchronization of complete graphs has since become a standard benchmark result in this line of research.
-For classical results on the Kuramoto model on complete graphs, we refer to [acebron2005, strogatz2000, PikovskyRosenblum2015, Bick2020].
-In the same work [taylor2012there], Taylor also initiated what is now known as the extremal combinatorics version of the problem [bandeira2025randomstrasse101openproblems2024], which sparked substantial interest and subsequent developments.
+Historically, Verwoerd and Mason [VM07] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing.
+This conjecture was resolved by Richard Taylor in 2012 [Tay12], and global synchronization of complete graphs has since become a standard benchmark result in this line of research.
+For classical results on the Kuramoto model on complete graphs, we refer to [ABPV+05, Str00, PR15, BGLM20].
+In the same work [Tay12], Taylor also initiated what is now known as the extremal combinatorics version of the problem [BKMR25], which sparked substantial interest and subsequent developments.
 A more detailed discussion of these advances will be given in Prior Work.
@@
 In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing.
-A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lageman2007] together with the center manifold theorem; see Lemma A.1 in [Geshkovski2025] for an excellent exposition.
-Prior Work.
-Existing guarantees for global synchronization on deterministic graphs focus on the extremal combinatorics version of the problem.
-The aim is to identify minimum-degree thresholds above which a graph is guaranteed to be globally synchronizing.
+A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lag07] together with the center manifold theorem; see Lemma A.1 in [GLPR25] for an excellent exposition.
+1.2. Prior Work
+Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions under which a graph is guaranteed to exhibit global synchronization.
+A prominent line of research formulates this question in an extremal combinatorics version:
 Problem 1.
-Let $\delta(G)$ denote the minimum degree of $G$. Find a constant $\mu_{c}\in[0,1]$ such that
-$\delta(G)\ \geq\ \mu_{c}(n-1)\ \Longrightarrow\ G\ \text{is globally synchronizing},$
-whereas for every $\varepsilon>0$ there exist graphs with $\delta(G)\leq(\mu_{c}-\varepsilon)(n-1)$ that are not globally synchronizing.
-The intuition behind this problem is that when every node is connected to a sufficiently large fraction of the network, the graph becomes dense enough for the collective attractive forces to guarantee convergence to synchrony.
-A remarkable property of these degree-based criteria is that, although they are only sufficient, they apply any graph meeting the threshold.
-This means global synchronization is guaranteed regardless of the graph’s finer structural details.
-Several advances have been made.
-The first result, due to Taylor [taylor2012there] in 2012, established a minimum-degree threshold of 0.9375(n1)0.9375(n-1).
-Subsequently, Ling [ling2019landscape] proved that global synchronization holds whenever the minimum degree exceeds 0.7929(n1)0.7929(n-1), and Lu [lu2021synchronization] further improved this bound to 0.7889(n1)0.7889(n-1).
-The conjecture asserts that the minimum-degree threshold for global synchronization is exactly $3/4$, and below this density, even very dense graphs may fail to synchronize.
-We refer to Conjecture 5 (density threshold for global synchrony) in [bandeira2025randomstrasse101openproblems2024] for a formal statement of this problem.
-Global synchrony on random graphs constitutes another important line of inquiry.
-Ling, Xu, and Bandeira [ling2019landscape] proved that for Erdős–Rényi graphs, global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$.
-This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [kassabov2022global].
-Both works raised the conjecture that synchronization should already occur as soon as the graph is connected, that is, for any $\epsilon>0$ and $p\geq(1+\epsilon)\log(n)/n$.
-This conjecture was recently resolved in [abdalla2024expandergraphsgloballysynchronizing].
-Further improvement has been made in [jain2025random, mcrae2025benign].
-For results on global synchrony of random geometric graphs, see [abdalla2024guarantees].
-Interestingly, recent work has pointed out that the idealized transformer dynamics admits a similar optimization form as the Kuramoto energy function [Geshkovski2025].
-This gives rise to the so-called nonlinear setting of the Kuramoto model, obtained by placing a nonlinear function on top of the inner product $E_{f}(\boldsymbol{\theta})=\-\sum_{1\leq i,j\leq n}A_{ij}f\big(\cos(\theta_{i}-\theta_{j})\big).$ The classical Kuramoto case corresponds to the linear function $f(t)=t$, while in the transformer dynamics one takes $f(t)=e^{\beta t}$ where $\beta>0.$ In this setting, the central question is to determine the values of $\beta$ for which global synchronization occurs, that is, the tokens collapse into a single cluster.
-Equivalently, this means that all token are making uniform predictions.
-A well-known conjecture predicted that global synchronization holds for all $\beta\geq 0$.
-The problem was originally formulated in [Geshkovski2025], with an initial result for $\beta\geq\Omega(1/n)$.
-It was subsequently improved in [Criscitiello2024] to cover the range $\beta\in[0,1]$, and was very recently settled by Polyanskiy, Rigollet, and Yao [polyanskiy2025synchronization].
-For more work connecting the Kuramoto model and Transformer (or attention) dynamics, see [geshkovski2023emergence, karagodin2024clustering, chen2025quantitative, polyanskiy2025synchronization, rigollet2025mean, Geshkovski2024] and the references therein.
-Our Contribution.
-Existing sufficient conditions for global synchronization of deterministic graphs require the graph to be sufficiently dense.
-In this work, we present a class of graphs that does not satisfy this minimum-degree requirement and whose densities range from 0 to $1$, yet globally synchronizes nevertheless.
-Through this result.
-Our main result is the following:
+Let $\delta(G)$ denote the minimum degree of $G$. Find the smallest constant $\mu_{c}\in[0,1]$ such that as long as $\delta(G)\ \geq\ \mu_{c}(n-1)$, the graph is globally synchronizing.
+The intuition for this problem is as follows.
+It is a classical result that complete graphs admit global synchronization [Tay12].
+In this case, each node is connected to all other n1n-1 nodes, corresponding to the maximal value $\mu_{c}=1$.
+A natural question is whether the global synchronization property persists when $\mu_{c}=1$ is relaxed, that is, when each node is connected to only a fraction of the network.
+Several advances have been made in [Tay12, LXB19, LS21, KST21, Can22, YTT21].
+The currently best known bounds on the critical minimum-degree threshold $\mu_{c}$ satisfy $0.6875\leq\mu_{c}\leq 0.75$, where the upper bound is established in [KST21], and the strongest known lower bound follows from [Can22].
+It has been conjectured that the minimum-degree threshold for global synchronization is exactly $\mu_{c}=0.75$, and that below this density, even highly connected graphs may fail to synchronize (see [TSS20] and Conjecture 5 in [BKMR25]).
+A remarkable feature of these minimum degree-based criteria is that, although they are only sufficient, they apply uniformly to all graphs meeting the prescribed threshold.
+In particular, global synchronization is guaranteed based solely on this parameter, without invoking any finer structural properties of the graph.
+Another important observation is that edge density alone is not an appropriate criterion for global synchronization.
+Indeed, Townsend, Stillman, and Strogatz [TSS20] provided examples of dense graphs that nevertheless fail to synchronize globally.
+Nevertheless, imposing a minimum degree condition necessarily forces the graph to be very dense.
+In particular, when $\mu_{c}\geq 75\%$, the total number of edges satisfies
+$|E|\geq\frac{3}{8}\,n(n-1),$
+which scales as $O(n^{2})$ and is therefore comparable to that of a complete graph.
+At the opposite end of the connectivity spectrum from dense networks, how sparse a globally synchronizing graph can be?
+Several existing works provide evidence toward this question.
+First, Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize.
+into a globally synchronizing network.
+It was shown that destabilizing all twisted states, which form a family of non-synchronous equilibria of the ring graph on $n$ nodes, requires adding $O(n\log n)$ edges arranged in a specific manner.
+Numerical experiments further suggest that, for networks of modest size, no non-synchronous stable states persist after such modifications, indicating that certain sparse circulant graphs may exhibit global synchronization.
+Second, results on globally synchronizing Erdős–Rényi graphs further indicate that networks substantially sparser than those satisfying the classical $75\%$ minimum-degree conditions can still exhibit global synchronization.
+Ling, Xu, and Bandeira [LXB19] first proved that global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$.
+This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [KST22].
+Both works raised the conjecture that global synchronization should already occur as soon as the graph is connected, namely, for any $\varepsilon>0$ and $p\geq(1+\varepsilon)\frac{\log n}{n}$.
+This conjecture was recently resolved in [ABK+24], with further refinements obtained in [JMS25, McR25].
+Taken together, the sparsest known families of globally synchronizing graphs beyond the trivial case of trees were supported by evidence at the $O(n\log n)$ edge scale.
+Whether global synchronization can be guaranteed for graphs with significantly fewer edges444beyond the trivial cases tree graphs remained mysterious.
+1.3. Our Contribution
+In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds.
 Theorem 1.2.
 Any connected threshold graph globally synchronizes.
-In the following, we briefly introduce the definition of threshold graphs and the motivation for studying their global synchronization.
-We conclude this section with a discussion of the structural mechanism that underlies the global synchronization of this graph class.
-(1) What Are Threshold Graphs?
 Threshold graphs can be constructed starting from a single vertex by iteratively adding each new vertex as either: an isolated vertex connecting to no existing vertex, or a dominating vertex connecting to all existing vertices.
 This generative process is uniquely encoded by a binary sequence of length n1n-1, where 0 denotes an isolated vertex and 1 a dominating one.
-Note that the binary sequence of any connected threshold graph must end with a 1; otherwise, the graph is not connected since there exists at least one isolated node.
+See Figure 2 for examples.
 A full formal definition and additional characterizations are given in Section 2.
@@
 Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
-(2) Why Threshold Graphs?
-Many theorems begin with initial observations from which suitable conjectures are formulated, and our case is no exception.
-The investigation into global synchronization on threshold graphs began with several simple observations:
-Observation 1. Star graphs and complete graphs are respectively the densest and sparsest connected threshold graphs, and both exhibit global synchronization.
-$\bigl(\underbrace{\texttt{0}\;\texttt{0}\;\cdots\;\texttt{0}}_{\text{all 0}},\;\texttt{1}\bigr)$ $\bigl(\underbrace{\texttt{1}\;\texttt{1}\;\cdots\;\texttt{1}}_{\text{all 1}}\bigr)$
-Star graph $S_{n}$ Complete graph $K_{n}$
-Certainly, this is not sufficient to propose the conjecture that all threshold graphs globally synchronize.
-Indeed, although the two extreme cases synchronize, this does not by itself imply that the entire spectrum of threshold graphs does.
-However, two further observations strengthened our confidence in this conjecture.
-Observation 2.
-One of the many equivalent definitions of threshold graphs requires that their neighborhood-inclusion preorder is total, i.e., the neighborhoods of any two nodes are nested: for any pair $u,v$, either $N(u)\subseteq N[v]$ or $N(v)\subseteq N(u)$.
-This total preorder of vicinities [threshold_graphs_book] suggests that influence may propagate step by step from the most dominant node to the rest without interference, hinting at the possibility of global synchronization.
-Observation 3.
-While far from exhaustive, our numerical experiments did not reveal any non-synchronous stable equilibria on connected threshold graphs.
-Our experiments are twofold: (a) To examine global synchronization precisely in small systems, we employed computational algebraic geometry following [mehta2015algebraic].
-More explicitly, we reformulated the equilibrium equations of the Kuramoto model as a system of polynomial equations and used the Bertini software package [BHSW13] to compute all solutions.
-We then evaluated the Jacobian at each solution to determine its stability.
-(b) For larger size systems (up to 200), using ode45 in Matlab, we initialized a large number of random initial conditions and observed that all trajectories converged to the fully synchronized state.
-(3) Differences Between Our Approach and Existing Techniques.
-Many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma.
-This lemma states that if all oscillators are confined to an open half-circle, then their dynamics will contract toward synchrony.
-Based on this observation, the bootstrap strategy attempts to show that all second-order stationary points must lie within a half-circle.
-If successful, this rules out the existence of non-synchronous state that the dynamics may converge to and thereby implies global synchronization.
+Notably, graphs in this class can have arbitrarily low or high edge density, ranging from $2/n$ to $1$, demonstrating that density alone is not an effective indicator of global synchronization.
+Moreover, their degree sequences are typically highly left-skewed, in sharp contrast to the relatively flat degree sequences arising in prior density-based and random graph models.
+This comparison further indicates that the shape of the degree sequence, by itself, is also not an appropriate indicator of global synchronization.
+We discuss the density and degree sequence of threshold graphs in Section 7.
+In terms of proving the main result, the difficulty arises from the fact that many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma.
+This lemma states that if all oscillators are confined to an open half-circle, then the dynamics contract toward synchrony.
+Building on this observation, the bootstrap strategy seeks to show that all second-order stationary points must lie within a half-circle.
+If successful, this excludes the existence of non-synchronous states to which the dynamics may converge, thereby implying global synchronization.
 However, for threshold graphs, it remains unclear how to apply this strategy.
@@
 Building on this understanding, we employ an inductive strategy: synchronization is first established locally on small substructures, and then shown to propagate across the whole graph along the neighborhood-inclusion hierarchy of threshold graphs.
-(4) Understanding the Mechanism Behind Global Synchrony in Threshold Graphs.
+Understanding the Mechanism Behind Global Synchrony in Threshold Graphs.
 Previous global synchronization results focus on density-based conditions.
-For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [lu2020synchronization].
+For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [LS20].
 In other words, these results require the graph to be sufficiently dense.
 Threshold graphs, however, lie outside this regime.
-Depending on the positions of dominating vertices in the generating sequence, their density can vary from $O(2/n)$ (stars) to $1$ (cliques), and many intermediate densities occur.
+Depending on the positions of dominating vertices in the generating sequence, their density can vary from $2/n$ (stars) to $1$ (cliques), and many intermediate densities occur.
 Nevertheless, we prove in this work that they also exhibit global synchronization.
@@
 Consequently, every threshold graph is globally synchronizing.
-Notation.
+1.4. Structure of the paper
+The preliminaries on threshold graphs and the Kuramoto model are presented in Sections 2 and 3, respectively.
+The basic geometric tools, referred to as phasor geometry or vector geometry, are introduced in Section 4.
+The key lemmas, termed Local Synchronization Primitives, which underpin our main proof, are presented in Section 5.
+The proof of the main result is given in Section 6.
+In Section 7, we discuss the density and degree sequence of threshold graphs.
+Finally, Section 8 concludes the paper with a discussion of open problems.
+1.5. Notation
 Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges.
@@
 2. Preliminaries of Threshold Graphs
-Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across various disciplines such as scheduling theory, psychology, computer science [threshold_graphs_book].
-Threshold graphs have been defined by Chvs and Hammer as follows [ChvatalHammer1973, ChvatalHammer1977].
+Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across a wide range of disciplines.
+They were discovered independently by researchers working in different fields and reported in separate venues.
+The term threshold graph was introduced by Chvátal and Hammer [CH73, CH77] in their study of set-packing problems.
+Around the same time, the same class of graphs was independently identified by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs.
+Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [MP95].
+2.1. Definition and Characterizations
 Definition 2.1.
@@
     add a dominating vertex, adjacent to all existing vertices.
-The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating.
-Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [threshold_graphs_book].
+The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating node.
+Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [MP95].
 Theorem 2.2.
@@
     For any distinct $u,v\in V$, $N(u)\subseteq N[v]\quad\text{or}\quad N(v)\subseteq N[u].$
-Remark 2.3 (Edge densities of threshold graphs).
-The edge density of threshold graphs ranges from $O(1/n)$ to $1$. Indeed, if $(b_{2},\dots,b_{n})$ is the generating sequence, then
-$|E|=\sum_{i:b_{i}=1}(i-1),\qquad\mathrm{dens}(G)=\frac{|E|}{\binom{n}{2}}$.
-By choosing the positions at which $b_{i}=1$, one obtains connected threshold graphs whose densities range from $2/n$ (stars) to $1$ (the complete graph), spanning many intermediate values.
-Moreover, threshold graphs are exactly the graphs whose degree sequences saturate the Erdős–Gallai inequalities in the characteristic Ferrers pattern corresponding to nested neighborhoods [threshold_graphs_book].
 11223344
@@
 Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered.
-To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [Ausiello1977], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs.
-This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [ChvatalHammer1977].
+To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [AMP77], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs.
+This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [CH77].
 3. Preliminaries on Kuramoto Model
@@
 In this section, we introduce the notion of equilibrium and stable equilibrium of the Kuramoto model (2).
-By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [khalil2002nonlinear].
+By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02].
 Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
@@
 $\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (5)
-An equilibrium point is considered stable if small perturbations around it remain small over time [khalil2002nonlinear].
+An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02].
 The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
@@
 The relation mentioned here is summarized in Figure 4.
-Such an equivalence was established in the celebrated work [Absil2006StableEquilibrium].
-SOSP of 4 Local minima of 4 = stable equilibria of 2
+Such an equivalence was established in the celebrated work [AK06].
+SOSP of (4) Local minima of (4) = stable equilibria of (2)
 Figure 4. Relation between different concepts.
@@
 $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}\quad\text{where}\quad\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}).$ (7)
-The proof can be found in [monzon2005global], we present here for completeness.
+where $v_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
+The proof can be found in [MP05], we present here for completeness.
 Proof of Lemma 4.1.
@@
 where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form.
-If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bhatia2007] [HornJohnson1985].
+If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bha07] [HJ85].
 ∎
@@
 ∎
-7. Discussion and Open Problems
-In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$, revealing a new mechanism of global synchronization that has not been articulated before.
+7. Discussion: Beyond Density-Based Guarantees for Global Synchronization
+As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction ($\geq 75\%$) of the remaining vertices.
+In this regime, graphs are necessarily highly dense, and their degree sequences are relatively homogeneous.
+This naturally raises the question:
+Does a graph that exhibits global synchronization necessarily have high edge density or a homogeneous degree sequence (i.e., with relatively small variation in node degrees)?
+In this section, we provide a negative answer to this question [^5] Of course, trees provide trivial examples.
+by exploiting two structural properties of threshold graphs, which were shown in the previous sections to be globally synchronizing.
+More precisely, we show that threshold graphs realize every admissible edge density among connected graphs.
+Moreover, their degree sequences are highly heterogeneous: among all connected graphs with a fixed edge density, threshold graphs attain the maximally left-skewed degree sequence when ordered in nonincreasing order.
+7.1. Threshold Graphs Realize All Densities
+Note that the edge density of a graph $G$ on $n$ vertices takes values in
+$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,\binom{n}{2}\}\,\right\}$. (20)
+and for connected graphs
+$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{n-1,\ldots,\binom{n}{2}\}\,\right\}$. (21)
+If the graph is not connected, to realize any density in $\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,n-2\}\,\right\}$, construct a threshold graph on $n$ vertices whose n1n-1-length sequence contains $1$ at the $m$-th position.
+001
+Figure 15.
+Illustration of the threshold graph construction.
+A single 11 at the third position of the generating sequence adds a dominating vertex, which contributes three incident edges shown as curved arcs.
+The following lemma shows that every admissible edge density in the range (21) is realized by a connected threshold graph on $n$ vertices.
+Lemma 7.1.
+Fix $n\geq 2$. For any integer $m\in[\,n-1,\binom{n}{2}\,$], there exists a connected threshold graph on $n$ vertices with exactly $m$ edges.
+Proof.
+Since the graph is connected, the last vertex added in the construction of a threshold graph must be a dominating vertex, contributing n1n-1 edges.
+To obtain a graph with exactly $m$ edges, the vertices added before the last one must therefore contribute a total of m(n1)m-(n-1) edges.
+Since $m\in\{n-1,\ldots,\binom{n}{2}\}$, it follows that
+$m-(n-1)\in\left\{0,1,\ldots,\binom{n}{2}-(n-1)\right\}=\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
+Edges are created only when dominating vertices are added during the construction.
+We fix the following convention: starting from an initial vertex, we index the construction sequence by counting only the dominating vertices added thereafter.
+Under this convention, a dominating vertex added at position $p$ contributes exactly $p$ new edges, as it connects to all previously existing vertices.
+Hence the possible edge contributions of dominating vertices, except for the last one added in the construction sequence, take values in $\{1,2,\ldots,n-2\}$.
+Therefore, the problem reduces to determining at which positions the dominating vertices except for the last one are added so that they collectively contribute m(n1)m-(n-1) edges.
+More precisely, it suffices to prove the following combinatorial statement:
+For any integer
+$k\in\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\},$
+there exists a subset
+$A\subseteq\{1,2,\ldots,n-2\}\quad\textit{such that}\quad\sum_{a\in A}a=k.$
+We prove this claim by induction on $n$.
+Base case. For $n=3$, we have $\{1,2,\ldots,n-2\}=\{1\}$ and $k\in\{0,1\}$. If $k=0$, take $A=\emptyset$; if $k=1$, take $A=\{1\}$.
+Induction hypothesis. Assume that for any
+$k\in\left\{0,1,2,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
+there exists a subset
+$A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime}_{k}}a=k.$
+Induction step. We now consider any
+$k\in\left\{0,1,2,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
+Case 1. If
+$k\leq\frac{(n-3)(n-2)}{2},$
+then by the induction hypothesis there exists $A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}$ such that $\sum_{a\in A^{\prime}_{k}}a=k$. Setting $A:=A^{\prime}_{k}$ yields the desired subset.
+Case 2. If
+$k>\frac{(n-3)(n-2)}{2},$
+then using only values in $\{1,2,\ldots,n-3\}$ cannot achieve the sum $k$, since the maximal possible sum is
+$\sum_{i=1}^{n-3}i=\frac{(n-2)(n-3)}{2},$
+which is strictly smaller than $k$. Hence the element n2n-2 must be included in $A$.
+It then remains to find a subset
+$A^{\prime\prime}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime\prime}}a=k-(n-2).$
+Since
+$k-(n-2)\in\left\{\frac{(n-2)(n-5)}{2}+1,\ldots,\frac{(n-2)(n-3)}{2}\right\}\subseteq\left\{0,1,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
+the induction hypothesis guarantees the existence of such a subset; in particular, we may take
+$A^{\prime\prime}=A^{\prime}_{\,k-(n-2)}.$
+Setting $A:=A^{\prime\prime}\cup\{n-2\}$ then yields $\sum_{a\in A}a=k$.
+This completes the induction and hence the proof. ∎
+7.2. Degree Sequence of Threshold Graphs
+One way to compare the “shape” of the degree sequences of two graphs with the same edge density is through majorization, which provides a partial order quantifying how spread out or concentrated the degrees are.
+More specifically, let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
+$\mathbf{d}(G)=(d_{1},\dots,d_{n}),\qquad\mathbf{d}(G^{\prime})=(d^{\prime}_{1},\dots,d^{\prime}_{n})$
+denote their degree sequences arranged in nonincreasing order. We say that $\mathbf{d}(G)$ majorizes $\mathbf{d}(G^{\prime})$, and write $\mathbf{d}(G)\succ\mathbf{d}(G^{\prime})$, if
+$\sum_{i=1}^{k}d_{i}\geq\sum_{i=1}^{k}d^{\prime}_{i}\quad\text{for all }k=1,\dots,n-1,$
+and
+$\sum_{i=1}^{n}d_{i}=\sum_{i=1}^{n}d^{\prime}_{i}.$
+The following is a classical result on degree sequence (called a threshold sequence) of threshold graphs.
+Theorem 7.2 ([PS89]).
+A degree sequence $\mathbf{d}$ is a threshold sequence if and only if it is not strictly majorized by any other graphical degree sequence.
+As shown in the previous section, for any admissible edge density there exists at least one (not necessarily unique) threshold graph realizing that density.
+Moreover, Theorem 7.2 tells that, among all graphs realizing a given edge density, threshold graphs are exactly those whose degree sequences are maximal with respect to majorization.
+In this sense, threshold graphs are always the most left-skewed among graphs with a fixed edge density, since their degree sequences are maximal under majorization, meaning that the degree mass is as concentrated as possible on high-degree vertices.
+8. Summary and Open Problems
+In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$.
 Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes.
@@
 Moreover, we expect that our results may admit further generalizations.
-Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [Geshkovski2025], and considering broader classes of graphs.
+Possible directions include extending the analysis to oscillators evolving on higher-dimensional spheres, incorporating nonlinear interaction terms as in [GLPR25], and considering broader classes of graphs.
 For instance, one may study systems in which oscillators are coupled through both positive and negative edges.
 Acknowledgements
-We would like to thank Afonso S. Bandeira for insightful discussions.
+We would like to thank Afonso S. Bandeira for insightful discussions and Kevin Lucca for pointing out relevant references on the known bounds of the density threshold.

### v5

@@
 Abstract.
-The Kuramoto model describes phase oscillators on the unit circle whose interactions are encoded by a graph.
-Each edge acts like a spring that pulls adjacent oscillators toward each other whenever their phases differ.
-A central question is to determine which graphs are globally synchronizing, meaning that trajectories of the Kuramoto dynamics converge to the fully synchronized state from almost all initial conditions.
-This property is tightly linked to the benign nonconvexity of the model’s energy landscape.
-Existing guarantees for global synchronization typically require every node to connect to sufficiently many neighbors, which in turn enforces rather homogeneous degree sequences and highly dense graphs.
-In this work, we identify threshold graphs as a class of graphs that lie outside this regime yet nevertheless exhibit global synchronization.
-More precisely, connected threshold graphs realize all admissible edge densities ranging from $2/n$ to $1$, and their degree sequences are maximally left-skewed among all graphs with the same edge density, when ordered in nonincreasing order.
-Our proof relies on a phasor-geometric analysis of the stationary points of the associated energy landscape.
+The Kuramoto model can be formulated as a gradient flow on a nonconvex energy landscape of the form $E(\boldsymbol{\theta}):=\frac{1}{2}\sum_{1\leq i,j\leq n}A_{ij}\bigl(1-\cos(\theta_{i}-\theta_{j})\bigr).$ A fundamental question is to identify graph structures for which this landscape is benign, in the sense that every second-order stationary point corresponds to a fully synchronized state.
+This property guarantees that all trajectories of the Kuramoto model converge to a fully synchronized state except for a measure-zero set of initial conditions, a phenomenon known as global synchronization.
+Existing guarantees typically require that each node be connected to a sufficiently large fraction of the other nodes, enforcing high graph density.
+In this work, we show that threshold graphs lie well outside this regime while still exhibiting global synchronization.
+In particular, threshold graphs realize arbitrary edge densities and have degree sequences that are extremal in the sense of majorization.
+Our analysis is based on a phasor–geometric characterization of stationary points that exploits the structural and geometric symmetries induced by threshold graphs.
 1. Introduction
@@
 Since then, synchronization phenomena have been widely studied across both natural and technological systems.
-For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [WCP+24], and Transformer architecture [GLPR25, CRMB24].
+For example, the flashing of fireflies [BB68, SS93], the rhythmic beating of heart cells [Win67, MS06], synchronization in power grids [DB12], coordinated motion in robotic networks [OS06], modern artificial intelligence models such as neural networks [MLGW24], and Transformer architecture [GLPR25, CRMB24].
 A major obstacle in the study of synchronization was the absence of a model that simultaneously offered analytical tractability and sufficient dynamical richness.
@@
 The Kuramoto model describes $n$ oscillators on the unit circle, coupled according to a graph $G=(V,E)$ with adjacency matrix $A$. Mathematically, it is a set of ordinary differential equations
-$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$ (1)
+$\frac{d\theta_{i}}{dt}=\omega_{i}+\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V,$
 where $\theta_{i}:\mathbb{R}\to\mathbb{S}^{1}$ denotes the phase of oscillator $i$.
@@
 A fundamental question in this setting is to determine the critical coupling strength $K_{c}$, such that for $K>K_{c}$, the oscillators converge to a common frequency, a phenomenon known as frequency synchronization [^1] That is, all derivatives $\tfrac{d\theta_{i}(t)}{dt}$ become identical as $t\to\infty$., and for $K<K_{c}$, they remain desynchronized.
-This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [Tas03, NM11] and power systems [CCC95, DB12].
+This problem is of central interest across many scientific domains, ranging from biological synchronization, rhythmic phenomena [MMJ87], technological applications such as deep brain stimulation [NM11] and power systems [CCC95, DB12].
 For comprehensive surveys of this line of work and its applications, we refer to [Str00, ABPV+05, DB14].
@@
 This becomes immediate after applying a co-rotating frame $\theta_{i}(t)\leftarrow\theta_{i}(t)-\omega t$, under which the dynamics reduce to the gradient system
-$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (2)
+$\frac{d\theta_{i}}{dt}=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i}),\qquad i\in V.$ (1)
 By LaSalle’s invariance principle, every trajectory converges to the invariant set of equilibria satisfying $\dot{\theta}_{i}=0$ for all $i$.
@@
 Over the past two decades, a central question that has emerged is the following: under what graph structures does the homogeneous Kuramoto system achieve phase synchronization for all initial conditions, except for a set of measure zero?
-This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin23] and the dynamics of Transformer architectures [GLPR25, GLPR23].
+This problem is not only mathematically appealing but also closely connected to several recent developments across different research areas, including the Burer–Monteiro approach for semidefinite programming [BBV16, LXB19, Lin25] and the dynamics of Transformer architectures [GLPR25, GLPR23].
 In this work, we follow this line of research, and beginning with a precise formulation of the problem.
@@
 Definition 1.1 (Globally synchronizing graph).
-We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$ except for a set of measure zero, the solution to the Kuramoto model (2) on $G$ satisfies
+We say that a graph $G=(V,E)$ is globally synchronizing if, for all initial conditions $\boldsymbol{\theta}(0)\in\mathbb{R}^{n}$, except for a set of measure zero, the solution to the Kuramoto model (1) on $G$ satisfies
 $\lim_{t\to\infty}(\theta_{i}(t)-\theta_{j}(t))=0\qquad\text{for all }i,j\in V.$
-The key question is now stated as:
-Which graphs are globally synchronizing?
+This naturally raises the following question:
+Key question. Which graphs are globally synchronizing?
 To build some intuition for the model (1) and the definition (1.1), one may imagine placing $n$ oscillators on the unit circle and connecting them by edges specified by the adjacency matrix $A$.
@@
 Nevertheless, one may form an intuitive if somewhat heuristic expectation: since in a complete graph each oscillator tends to move closer to every other oscillator, the abundance of interactions may indeed drive the system to synchronize from almost all initial configurations.
-Historically, Verwoerd and Mason [VM07] conjectured in 2007 that every complete graph $K_{n}$ ($n\geq 2$) is globally synchronizing.
-This conjecture was resolved by Richard Taylor in 2012 [Tay12], and global synchronization of complete graphs has since become a standard benchmark result in this line of research.
-For classical results on the Kuramoto model on complete graphs, we refer to [ABPV+05, Str00, PR15, BGLM20].
-In the same work [Tay12], Taylor also initiated what is now known as the extremal combinatorics version of the problem [BKMR25], which sparked substantial interest and subsequent developments.
-A more detailed discussion of these advances will be given in Prior Work.
-One standard approach to this problem is to study the corresponding energy landscape.
-Note that the Kuramoto model (2) can be formulated as a gradient flow of a certain energy function.
-Precisely, the system of ordinary differential equations (2) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$.
+1.2. Nonconvex Optimization Landscape
+From an optimization viewpoint, global synchronization can be implied by the absence of spurious second-order stationary points of the nonconvex energy function.
+Precisely, the Kuramoto model (1) defines a gradient dynamical system (or called gradient flow [^2] We call a dynamical system a gradient flow on $E(x)$ if it has the form $\frac{dx(t)}{dt}=-\nabla E(x)$.
 Its trajectories evolve along the steepest descent direction of $E$.) with energy function $E_{G}(\boldsymbol{\theta})$, satisfying
-$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$ (3)
+$\frac{d\boldsymbol{\theta}}{dt}=-\nabla E_{G}(\boldsymbol{\theta})$
 where $E_{G}(\boldsymbol{\theta}):(\mathbb{S}^{1})^{|V|}\rightarrow\mathbb{R}$ with graph $G=(V,E)$ is defined as
 $E_{G}(\boldsymbol{\theta})$= $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\cos(\theta_{i}-\theta_{j})\big)$ (2)
-$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle v_{i},v_{j}\rangle\big).$
-Here, $v_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
+$=$ $\frac{1}{2}\sum_{1\leq i,j\leq n}\boldsymbol{A}_{ij}\big(1-\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle\big).$
+Here, $\mathbf{v}_{i}$ is the unit vector $(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}$.
 Clearly, this function is nonconvex.
 Moreover, understanding the geometry of (2) provides direct insight into the long-term dynamics of (1).
-In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minimua) correspond to non-synchronous stable equilibria.
+In particular, the synchronous states correspond exactly to the global minima of this energy function, and those minima that are local but not global (called spurious local minima) correspond to non-synchronous stable equilibria.
 If such spurious local minima exist, the gradient dynamics may converge to them whenever the initialization lies within their basin of attraction, in which case the graph $G$ is not globally synchronizing.
-In fact, if one can show that the energy function (4) admits no spurious second-order stationary point (in such case we say the function is benign), then the graph $G$ is globally synchronizing.
+In contrast, if the energy function (2) admits no spurious second-order stationary points, in the sense that the only equilibria at which the Hessian $\nabla^{2}E_{G}(\boldsymbol{\theta})$ is positive semidefinite are the fully synchronized states, then the graph $G$ is globally synchronizing.
 A rigorous proof of this statement can be established using the Łojasiewicz gradient inequality [Lag07] together with the center manifold theorem; see Lemma A.1 in [GLPR25] for an excellent exposition.
 1.3. Prior Work
-Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions under which a graph is guaranteed to exhibit global synchronization.
-A prominent line of research formulates this question in an extremal combinatorics version:
+Previous works on globally synchronizing graphs primarily focus on identifying sufficient conditions that guarantee global synchronization.
+A dominant line of research formulates this question from an extremal combinatorics perspective:
 Problem 1.
 Let $\delta(G)$ denote the minimum degree of $G$. Find the smallest constant $\mu_{c}\in[0,1]$ such that as long as $\delta(G)\ \geq\ \mu_{c}(n-1)$, the graph is globally synchronizing.
-The intuition for this problem is as follows.
-It is a classical result that complete graphs admit global synchronization [Tay12].
-In this case, each node is connected to all other n1n-1 nodes, corresponding to the maximal value $\mu_{c}=1$.
-A natural question is whether the global synchronization property persists when $\mu_{c}=1$ is relaxed, that is, when each node is connected to only a fraction of the network.
-Several advances have been made in [Tay12, LXB19, LS21, KST21, Can22, YTT21].
+Several advances have been made in [Tay12, LXB19, LS20, KST21, Can22, YTT21].
 The currently best known bounds on the critical minimum-degree threshold $\mu_{c}$ satisfy $0.6875\leq\mu_{c}\leq 0.75$, where the upper bound is established in [KST21], and the strongest known lower bound follows from [Can22].
 It has been conjectured that the minimum-degree threshold for global synchronization is exactly $\mu_{c}=0.75$, and that below this density, even highly connected graphs may fail to synchronize (see [TSS20] and Conjecture 5 in [BKMR25]).
-A remarkable feature of these minimum degree-based criteria is that, although they are only sufficient, they apply uniformly to all graphs meeting the prescribed threshold.
-In particular, global synchronization is guaranteed based solely on this parameter, without invoking any finer structural properties of the graph.
-Another important observation is that edge density alone is not an appropriate criterion for global synchronization.
-Indeed, Townsend, Stillman, and Strogatz [TSS20] provided examples of dense graphs that nevertheless fail to synchronize globally.
-Nevertheless, imposing a minimum degree condition necessarily forces the graph to be very dense.
-In particular, when $\mu_{c}\geq 75\%$, the total number of edges satisfies
-$|E|\geq\frac{3}{8}\,n(n-1),$
-which scales as $O(n^{2})$ and is therefore comparable to that of a complete graph.
-At the opposite end of the connectivity spectrum from dense networks, how sparse a globally synchronizing graph can be?
-Several existing works provide evidence toward this question.
-First, Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize.
+Sufficiently large threshold for minimal degree enforces high density.
+Indeed, when $\mu_{c}\geq 75\%$, the total number of edges satisfies $|E|\geq\frac{3}{8}\,n(n-1)$, which scales as $O(n^{2})$.
+However, dense graphs do not necessarily global synchronize, as indicated by the examples in [TSS20].
+How sparse a globally synchronizing graph can be?
+Townsend, Stillman, and Strogatz [TSS20] approached this problem by asking how few edges are required to transform a ring graph [^3] A ring graph of size $n\geq 5$ does not globally synchronize.
 into a globally synchronizing network.
-It was shown that destabilizing all twisted states, which form a family of non-synchronous equilibria of the ring graph on $n$ nodes, requires adding $O(n\log n)$ edges arranged in a specific manner.
-Numerical experiments further suggest that, for networks of modest size, no non-synchronous stable states persist after such modifications, indicating that certain sparse circulant graphs may exhibit global synchronization.
-Second, results on globally synchronizing Erdős–Rényi graphs further indicate that networks substantially sparser than those satisfying the classical $75\%$ minimum-degree conditions can still exhibit global synchronization.
-Ling, Xu, and Bandeira [LXB19] first proved that global synchronization occurs with high probability as $n\to\infty$ provided that the edge probability satisfies $p\gg\log(n)/n^{1/3}$.
-This bound was later improved by Kassabov, Strogatz, and Townsend to $p\gg\log^{2}(n)/n$ [KST22].
-Both works raised the conjecture that global synchronization should already occur as soon as the graph is connected, namely, for any $\varepsilon>0$ and $p\geq(1+\varepsilon)\frac{\log n}{n}$.
-This conjecture was recently resolved in [ABK+24], with further refinements obtained in [JMS25, McR25].
-Taken together, the sparsest known families of globally synchronizing graphs beyond the trivial case of trees were supported by evidence at the $O(n\log n)$ edge scale.
-Whether global synchronization can be guaranteed for graphs with significantly fewer edges444beyond the trivial cases tree graphs remained mysterious.
+They showed that destabilizing all non-synchronizing twisted states requires adding $O(n\log n)$ edges arranged in a specific manner, indicating that adding $O(n\log n)$ edges to a ring graph in a strategic way may lead to global synchronization.
+Moreover, a line of work on global synchronization over random graphs [LXB19, KST22, ABK+26, JMS25, McR25] provides further evidence that sparse networks can globally synchronize.
+A natural question is to determine the critical connectivity threshold for Erdős–Rényi graphs to exhibit global synchronization with high probability.
+The best known bound asserts that, for any $\varepsilon>0$, global synchronization holds with high probability when $p\geq(1+\varepsilon)\frac{\log n}{n},$ as shown in [ABK+26], with further refinements in [JMS25, McR25].
+However, this regime still corresponds to graphs with $O(n\log n)$ edges.
 1.4. Our Contribution
-In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds.
+In this work, we identify a class of graphs that admit global synchronization without requiring the minimum degree to exceed prescribed thresholds, and can have any edge density as will be shown in Section 7.
 Theorem 1.2.
@@
 Figure 2.
-Four connected threshold graphs on 1919 vertices, from the sparsest ($S_{18}$) to the densest ($K_{19}$).
-Node labels indicate the construction order: vertex $1$ is the initial white vertex, and all subsequent labels follow the binary code of the construction.
-White and black circles denote isolated and dominating vertices (bit 0 and bit 11).
+Four connected threshold graphs on 1919 vertices, from the sparsest (star graph) to the densest (complete graph).
+Node labels indicate the vertex addition order in the construction, with vertex $1$ as the initial vertex.
+White and black circles denote isolated and dominating vertices (bit 0 and bit 11), and the initial vertex $1$ is colored in white.
 Codes from left to right: 000000000000000001, 000100000000010001, 010001101101010001, $111111111$.
-Notably, graphs in this class can have arbitrarily low or high edge density, ranging from $2/n$ to $1$, demonstrating that density alone is not an effective indicator of global synchronization.
-Moreover, their degree sequences are typically highly left-skewed, in sharp contrast to the relatively flat degree sequences arising in prior density-based and random graph models.
-This comparison further indicates that the shape of the degree sequence, by itself, is also not an appropriate indicator of global synchronization.
-We discuss the density and degree sequence of threshold graphs in Section 7.
 In terms of proving the main result, the difficulty arises from the fact that many existing proofs of global synchronization rely, either explicitly or implicitly, on the so-called half-circle lemma.
@@
 It typically depends on sufficient connectivity in the network thus naturally suited to dense graphs, where the large number of edges ensures that nodes inside the half-circle collectively exert sufficient force to pull outside nodes back into alignment.
-In contrast, threshold graphs span the full spectrum from extremely sparse examples such as stars to dense cliques, and their highly inhomogeneous, layered structure complicates the application of this strategy.
-As a result, existing techniques offer limited insight into whether and why these graphs synchronize globally.
+In contrast, threshold graphs span the full spectrum from extremely sparse examples, such as stars, to extremely dense ones, such as complete graphs, making it unclear how such techniques can be applied in this setting.
 In this work, we bypass the half-circle approach by reducing the high-dimensional geometric question of whether the energy landscape is benign to a planar vector-geometric condition.
@@
 Previous global synchronization results focus on density-based conditions.
-For instance, any graph whose minimum degree exceeds 0.7889(n1)0.7889(n-1) is globally synchronizing [LS20].
+For instance, any graph whose minimum degree exceeds 0.75(n1)0.75(n-1) is globally synchronizing [KST21].
 In other words, these results require the graph to be sufficiently dense.
 Threshold graphs, however, lie outside this regime.
-Depending on the positions of dominating vertices in the generating sequence, their density can vary from $2/n$ (stars) to $1$ (cliques), and many intermediate densities occur.
 Nevertheless, we prove in this work that they also exhibit global synchronization.
@@
 1.6. Notation
-Throughout this work, we consider only connected simple graphs, i.e., graphs without self-loops or weighted edges.
-Following the convention of dynamical systems, we refer to the variable $\theta=(\theta_{1},\cdots,\theta_{n})$ of the Kuramoto model (2) as a state, and call the solution $\boldsymbol{\theta}(t)$ with respect to the initial condition $\boldsymbol{\theta}(0)$ in the Kuramoto model (2) a trajectory starting from $\boldsymbol{\theta}(0)$.
-Moreover, $\mathbb{S}^{1}:=\{x\in\mathbb{R}^{2}:\|x\|=1\}$ denotes the unit circle.
-And $\boldsymbol{x}\Uparrow\boldsymbol{y}$ means that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear: $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$.
-We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\nsim v$ otherwise.
+Throughout this work, all graphs are assumed to be connected and simple, i.e., without self-loops or weighted edges.
+We write $u\sim v$ to indicate that vertices $u$ and $v$ are adjacent in a graph, and $u\not\sim v$ otherwise.
 For a vertex $i$ and a subset $Q$ of vertices, we denote
-$N_{Q}(i)=\{\,q\in Q:q\sim i\,\},$
-the set of neighbors of $i$ that lie in $Q$.
+$N_{Q}(i):=\{q\in Q:q\sim i\},$
+[…]
+We denote by $\mathbb{S}^{1}:=\{\boldsymbol{x}\in\mathbb{R}^{2}:\|\boldsymbol{x}\|=1\}$ the unit circle.
+For vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{2}$, we write $\boldsymbol{x}\Uparrow\boldsymbol{y}$ to indicate that $\boldsymbol{x}$ and $\boldsymbol{y}$ are nonzero and positively collinear, i.e., $\boldsymbol{x}=\lambda\boldsymbol{y}$ for some $\lambda>0$.
+Following the convention in dynamical systems, we refer to the variable $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{n})$ of the Kuramoto model (1.2) as the state, and call the solution $\boldsymbol{\theta}(t)$ with initial condition $\boldsymbol{\theta}(0)$ a trajectory starting from $\boldsymbol{\theta}(0)$.
 2. Preliminaries of Threshold Graphs
 Threshold graphs are a well-studied class of graphs that play a central role in graph theory and have found applications across a wide range of disciplines.
-They were discovered independently by researchers working in different fields and reported in separate venues.
-The term threshold graph was introduced by Chvátal and Hammer [CH73, CH77] in their study of set-packing problems.
-Around the same time, the same class of graphs was independently identified by Henderson and Zalcstein in the context of parallel processing, where they were referred to as PV-chunk definable graphs.
-Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [MP95].
+The term threshold graph was introduced by Chvátal and Hammer [CH77b] in their study of set-packing problems.
+Since then, threshold graphs have appeared in numerous settings, including resource allocation, cyclic scheduling, manpower planning, complexity theory, and the study of graphical degree sequences [CH77a].
 2.1. Definition and Characterizations
@@
 The resulting graph on vertices $\{1,\dots,n\}$ is uniquely encoded by the binary sequence $(b_{2},\dots,b_{n})$, where $b_{i}=0$ if vertex $i$ is added as isolated and $b_{i}=1$ if it is added as dominating node.
-Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independant interest to recall that they admit several equivalent characterizations [MP95].
+Although in this work we rely only on the construction process of threshold graphs to analyze their global synchronization, it is of independent interest to recall that they admit several equivalent characterizations [CH77b].
 Theorem 2.2.
@@
   4. (iv)
-    $G$ is $\{P_{4},C_{4},2K_{2}\}$-free.
-  5. (vi)
+    $G$ is $\{P_{4},C_{4},2K_{2}\}$-free (see Fig. 3).
+  5. (v)
     For any distinct $u,v\in V$, $N(u)\subseteq N[v]\quad\text{or}\quad N(v)\subseteq N[u].$
@@
 Interestingly, threshold graphs appeared in the synchronization literature long before their role in the Kuramoto model was considered.
-To the best of our knowledge, the earliest instance dates back to 1977, with the seminal work of Ausiello, Messina, and Protasi [AMP77], who provided a graph-theoretic characterization of the PVc class of synchronizing primitives and introduced the notion of PVc-definable graphs.
-This line of work emerged almost simultaneously with the independant introduction of threshold graphs by Chvátal and Hammer in their study of aggregation of inequalities in integer programming [CH77].
-3. Preliminaries on Kuramoto Model
-In this section, we introduce two families of concepts for the Kuramoto model: one arising from dynamical systems and the other from nonconvex optimization.
-Presenting these notions side by side is important, as many of the local synchronization lemmas developed later rely on interpreting stability to infer the relative positions of phasors at second-order stationary points of the energy landscape.
-3.1. Equilibrium and Local Stability
-In this section, we introduce the notion of equilibrium and stable equilibrium of the Kuramoto model (2).
-By the definition, equilibrium points of a dynamical system are the points where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02].
-Setting the right-hand side of (2) to zero provides the characterization of an equilibrium point in the homogeneous Kuramoto model on the graph $G$.
-Definition 3.1 (Equilibrium of the Kuramoto model (2)).
+In particular, as early as 1977, Ausiello, Messina, and Protasi [AMP77] introduced a class of graphs (later recognized as threshold graphs) in the study of synchronization primitives in distributed computation.
+Their work provided a graph-theoretic characterization of so-called PVc-definable graphs, which encode combinatorial conditions under which simple local update rules lead to global agreement.
+3. Preliminaries on the Kuramoto Model
+3.1. Equilibria and Local Stability
+By the definition, equilibrium points of a dynamical system are the states where the system remains at rest, meaning that the first-order derivative of the state vanishes [Kha02].
+Thus setting the right-hand side of (1) to zero provides the characterization of an equilibrium point of the Kuramoto model.
+Definition 3.1 (Equilibria of (1)).
 We call $\boldsymbol{\theta}\in\mathbb{R}^{n}$ an equilibrium of the Kuramoto model (1) if it satisfies
-$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (5)
-An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02].
-The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
-Proposition 3.2 (Characterization of stable equilibrium).
-An equilibrium point $\boldsymbol{\theta}$ is stable if the Jacobian matrix $\boldsymbol{\mathbf{J}}$, defined as:
+$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\quad\forall i\in[n].$ (3)
+An equilibrium point is said to be stable if sufficiently small perturbations around it remain small for all future time [Kha02].
+To assess the stability of a given equilibrium point, one typically studies the linearization of the system at that point, which is characterized by the Jacobian matrix.
+Proposition 3.2 (Stable equilibrium of (1)).
+An equilibrium point $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model (1) is stable if the Jacobian matrix $\boldsymbol{\mathbf{J}}$, defined as:
 $\boldsymbol{\mathbf{J}}_{ij}=\begin{cases}-\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i=j,\\[8.0pt] \boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i\neq j,\end{cases}$
 is negative semi-definite.
-3.2. Stationary Point and Local Optimality
-Definition 3.3 (Stationary point of the energy function (4)).
-We call $\theta\in\mathbb{R}^{n}$ a stationary point of the energy function (4) if it satisfies $\nabla E(\theta)=0$, namely
-$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (6)
-Obviously, from Proposition 3.1 and Definition 3.3, a configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a stationary point of the energy function (4) if and only if it is an equilibrium of the Kuramoto model (2).
+3.2. Stationary Points and Local Optimality
+Definition 3.3 (First-order stationary points of (2)).
+We call $\boldsymbol{\theta}\in\mathbb{R}^{n}$ a first-order stationary point of the energy function (2) if it satisfies $\nabla E(\boldsymbol{\theta})=0$, namely
+$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\quad\forall i\in[n].$ (4)
+Proposition 3.4 (Hessian $\boldsymbol{H}$ of (2)).
+For the Kuramoto energy (2), the Hessian has entries
+$\boldsymbol{H}_{ij}:=(\nabla^{2}E(\boldsymbol{\theta}))_{ij}=\begin{cases}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i=j,\\[8.0pt] -\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i\neq j,\end{cases}$
+Definition 3.5 (Second-order stationary points of (2)).
+A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is called a second-order stationary point (SOSP) of (2) if $\nabla E(\boldsymbol{\theta})=0$, and the Hessian $\boldsymbol{H}$ is positive semidefinite.
+3.3. Correspondence Between the Two Sets of Concepts
+Proposition 3.6.
+A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is an equilibrium of the Kuramoto model (1) if and only if it is a first-order stationary point of the Kuramoto energy (2).
+Proof.
+This is clear, since both are characterized by
+$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\qquad\forall i\in[n].$
+Moreover, since (2) is real analytic, an equilibrium of (1) is stable if and only if it is a local minimum of (2), modulo the global rotation symmetry.
+Such an equivalence was established in the celebrated work [AK06].
 Definition 3.7 (Local minimum of the energy function).
-A configuration $\theta\in\mathbb{R}^{n}$ is called a local minimum of the energy function $E$ if there exists $\varepsilon>0$ such that
-$E(\theta)\leq E(\theta^{\prime}),\quad\forall\theta^{\prime}\in\mathbb{R}^{n}\ \text{with}\ \|\theta^{\prime}-\theta\|<\varepsilon.$
-Proposition 3.5 (First- and second-order necessary conditions).
-If a configuration $\theta$ is a local minimum of $E$, then
-  1. (1)
-    $\nabla E(\theta)=0$ (first-order condition), and
-  2. (2)
-    the Hessian $\nabla^{2}E(\theta)$ is positive semidefinite.
-A point satisfying the above two conditions is also called a second-order stationary point (SOSP) of (4).
-3.3. Correspondence between Stability and Optimality
-In this section, we briefly outline the relationship between the equilibrium stability of the dynamical system (2) and the local optimality of the energy function (4).
-Understanding this correspondence provides valuable intuition.
-In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order stationary point.
-At the first-order level, the two notions coincide.
-Indeed, a configuration $\boldsymbol{\theta}$ is a first-order stationary point of the energy function (4) if and only if it is an equilibrium of the dynamical system (2) on the graph $G$ with adjacency matrix $A$, since both are characterized by
-$\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i})=0\qquad\text{for all }i\in V.$
-One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape.
-It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4).
-The relation mentioned here is summarized in Figure 4.
-Such an equivalence was established in the celebrated work [AK06].
+A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is called a local minimum of the energy function $E(\boldsymbol{\theta})$ if there exists $\varepsilon>0$ such that
+$E(\boldsymbol{\theta})\leq E(\boldsymbol{\theta}^{\prime}),\quad\forall\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{n}\ \text{with}\ \|\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}\|<\varepsilon.$
+Therefore, the relations between stable equilibria, second-order stationary points and local minima can be summarized in Figure 4.
 SOSP of (2) Local minima of (2) = stable equilibria of (1)
@@
 4. Vector Geometry of Equilibria
-In this section, we recast the algebraic first- and second-order stationary conditions in Section 3 into geometric ones.
-This geometric viewpoint replaces the nonlinear equations (6) and the Hessian constraints by combinatorial relations among the vectors associated with the nodes.
-They form the starting point of our analysis.
-Additionally, stronger geometric constraints will be derived in later sections and combined with the results here to rule out all non-synchronous second-order stationary points of the energy function $E_{G}(\boldsymbol{\theta})$ where $G$ is any connected threshold graph.
+In this section, we recast the algebraic conditions introduced in Section 3 into geometric ones, namely combinatorial relations among vectors associated with the nodes.
+By doing so, the analysis of global synchronization on threshold graphs reduces to reasoning purely about geometric relations between the unit vectors corresponding to the angles in $\boldsymbol{\theta}$.
 4.1. Vector-labeled Graph
-The basic object underlying this geometric perspective is the vector-labeled graph corresponding to angle configuration of $\boldsymbol{\theta}$.
-Defining such object allows us reasoning about whether a given state $\boldsymbol{\theta}$ is an equilibrium, and whether an equilibrium is stable, i.e., can we move the angles so as to decrease the energy.
-Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
+Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
 $\boldsymbol{v}_{i}:=(\cos\theta_{i},\sin\theta_{i})\in\mathbb{R}^{2}.$
-This is called a phasor, following the standard convention in the Kuramoto model.
+This is called a phasor, following the terminology of Yoshiki Kuramoto.
 The object $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$, called vector-labeled graph encoding $\boldsymbol{\theta}$ combinatorially and geometrically.
@@
 4.2. Alignment and Equilibrium
-Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms.
+Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\}_{i\in V})$ associated with the state $\boldsymbol{\theta}$, we reformulate the equilibrium condition (3) (equivalently (4)) in geometric terms.
 Lemma 4.1 (Equilibrium condition rephrased geometrically).
@@
 $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}\quad\text{where}\quad\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}).$ (5)
-where $v_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
+where $\boldsymbol{v}_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
 The proof can be found in [MP05], we present here for completeness.
 Proof of Lemma 4.1.
-Let us firstly derive an equivalent form of (7). Rewrite all vectors $\boldsymbol{v}_{i}$ in (7) in terms of their Cartesian components $(\cos\theta_{i},\sin\theta_{i})$, yielding
+The condition (5) is equivalently written as
+$\sum_{j\in N(i)}e^{\mathrm{i}\theta_{j}}=\mu_{i}e^{\mathrm{i}\theta_{i}}.$ (6)
+Indeed, rewrite all vectors $\boldsymbol{v}_{i}$ in (5) in terms of their Cartesian components, yielding
 $\sum_{j\in N(i)}(\cos\theta_{j},\sin\theta_{j})=\mu_{i}(\cos\theta_{i},\sin\theta_{i}).$
-Equivalently, this can be written as
+Equivalently,
 $\left\{\begin{aligned} \sum_{j\in N(i)}\cos\theta_{j}&=\mu_{i}\cos\theta_{i},\\ \sum_{j\in N(i)}\sin\theta_{j}&=\mu_{i}\sin\theta_{i}.\end{aligned}\right$. (7)
-This is equivalent to verifying the equality between two complex numbers:
+The condition (7) is equivalent to the equality between the following two complex numbers:
 $\sum_{j\in N(i)}\cos\theta_{j}+I\sum_{j\in N(i)}\sin\theta_{j}=\mu_{i}\cos\theta_{i}+I\mu_{i}\sin\theta_{i},$
-which can be compactly rewritten as
-$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}.$ (9)
-using the Euler formula $e^{I\theta_{i}}=\cos\theta_{i}+I\sin\theta_{i}$.
-Now that we have derived the equivalence between conditions (7) and (9), our goal now reduces to proving the equivalence between (9) and (4), namely
-$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}\Longleftrightarrow\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0.$
-This indeed holds.
-To prove the implication from the right-hand side to the left-hand side, we construct such a $\mu_{i}$ that satisfies the left-hand side.
-Dividing both sides of (9) by $e^{I\theta_{i}}$ (noting that $e^{I\theta_{i}}\neq 0$), we obtain
-$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\mu_{i}.$
-Such a $\mu_{i}$ exists and is in fact given by
-$\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}),$
-at an equilibrium point $\boldsymbol{\theta}$, as we will show below. Expanding the fraction, we obtain
-$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$ $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$ (10)
-$=\sum_{j\in N(i)}(\cos(\theta_{j}-\theta_{i})+I\sin(\theta_{j}-\theta_{i})).$
-Since $\boldsymbol{\theta}$ is an equilibrium, the imaginary part must vanish, leading to
-$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
-Thus, we conclude
-$\sum_{j\in N(i)}e^{I\theta_{j}}=\sum_{j\in N(i)}{\cos(\theta_{j}-\theta_{i})}\cdot e^{I\theta_{i}}.$
-Conversely, to prove the left-hand side implies the right-hand side, we show that if there exists a constant $\mu_{i}$ such that
-$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}},$
-then $\boldsymbol{\theta}$ is an equilibrium point. This is equivalent to showing that
+which can be compactly rewritten as (6) using the Euler formula $e^{I\theta_{i}}=\cos\theta_{i}+I\sin\theta_{i}$.
+Now that we have derived the equivalence between conditions (5) and (6), our goal reduces to proving the equivalence between (6) and (3), namely
+$\forall i\in V,\qquad\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}\Longleftrightarrow\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0.$
+The implication from the left-hand side to the right-hand side is established by showing that
 $I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
-Indeed, we compute
-$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$ (11)
+This can be verified by a direct computation:
+$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$
 $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$
@@
 $=\mu_{i}.$
-Thus, the proof is complete. ∎
+To prove the other direction, we construct for every node $i\in V$, such a $\mu_{i}$ that satisfies the left-hand side.
+Dividing both sides of (6) by $e^{I\theta_{i}}$ (noting that $e^{I\theta_{i}}\neq 0$), we obtain
+$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\mu_{i}.$
+Compute
+$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$ $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$
+$=\sum_{j\in N(i)}(\cos(\theta_{j}-\theta_{i})+I\sin(\theta_{j}-\theta_{i}))$
+$=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$
+where the last equality is due to the fact $\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0$. Thus
+$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}$
+holds for every $i\in V$ by letting $\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$. ∎
 Remark 4.2 (Normalization and local coherence).
@@
 Due to the unit circle geometry $\left|\boldsymbol{v}_{i}\right|=1$ of the Kuramoto model, such a scaling force does not induce any motion while a tangential component can cause rotation.
+4.3. Angles and Second-Order Stationarity
+Given an equilibrium $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model on graph $G=(V,E)$, if there exists a node $i\in V$ whose corresponding $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point of the Kuramoto energy (2).
+This is because a small perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
+Lemma 4.4.
+If $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a second-order stationary point of (2), then for every $i\in V$ there exists $\mu_{i}\geq 0$ such that
+$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
+In particular, either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ (i.e., $\mu_{i}=0$), or $\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0$.
+Proof of Lemma 4.4.
+By Proposition 3.4, the diagonal entry of the Hessian satisfies
+$\boldsymbol{H}_{ii}=\sum_{j\in N(i)}\cos(\theta_{i}-\theta_{j}).$
+Since $\theta$ is an equilibrium, we have
+$\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0,$
+which implies that
+$\sum_{j\in N(i)}e^{\mathrm{i}(\theta_{j}-\theta_{i})}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
+By Lemma 4.1, this real quantity equals $\mu_{i}$ at first-order stationary point $\boldsymbol{\theta}$.
+Hence $\boldsymbol{H}_{ii}=\mu_{i}$.
+If $\mu_{i}<0$, then $\boldsymbol{H}_{ii}<0$, contradicting the positive semidefiniteness of $\boldsymbol{H}$.
+∎
 (A)(B)(C)(D)
 Figure 6.
-Geometric intuition for the first- and second-order stationary conditions.
 The solid arrow represents $\boldsymbol{v}_{i}$, and the dashed arrow represents $\sum_{j\in N(i)}\boldsymbol{v}_{j}$.
-For each node $i$, the four panels show all possible configurations satisfying the first-order condition.
-Among them, cases (A) and (B) violate the second-order condition because the aggregated phasor lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
-4.3. Angle and Stability
-Given an equilibrium $\boldsymbol{\theta}$, if there exists a node $i\in V$ whose corresponding force satisfies $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point.
-Indeed, a perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
-Lemma 4.4.
-If $\boldsymbol{\theta}$ is a second-order stationary point of (4), then either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ or
-$\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0.$
-Proof of Lemma 4.4.
-The $i$-th diagonal entry of $\mathrm{Hess}\,E(\theta)$ can be written as
-$\mathrm{Hess}\,E(\boldsymbol{\theta})_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
-where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form.
-If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bha07] [HJ85].
-∎
-Remark 4.5 (Mechanical insight into node at stable equilibrium).
-The equilibrium condition in Lemma 4.4 for node $i\in V$ has a mechanical interpretation.
-The sum of the phasors representing its neighbors can be visualized as a set of unit-length rods with directions.
-We place one phasor’s starting point at the origin, then connect the remaining phasors head-to-tail.
-The head-to-tail connected rods can freely adjust their orientations, but the endpoint of the last phasor must remain on the right-end unrestricted sliding track, to ensure node $i$ being stable there.
-Figure 7.
-With $\boldsymbol{v}_{i}$ fixed at angle 0, the endpoint of the aggregated phasor $\sum_{j}A_{ij}\boldsymbol{v}_{j}$ is constrained to the right-end sliding track.
-The rods may freely adjust their orientations, but the endpoint of the last rod must lie in the feasible sliding interval.
-The solid vectors show the unit-length phasors of the neighbors of node ii.
+For each node $i$, the four panels show all possible relations between $\boldsymbol{v}_{i}$ and $\sum_{j\in N(i)}\boldsymbol{v}_{j}$.
+Among them, case (A) violates the first-order condition, while case (B) violates the second-order stationary condition, since the aggregated phasor $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
 The geometric conditions of second-order stationary points presented here are only the most basic one.
@@
 Lemma 5.1.
-Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}\subset\mathbb{R}^{2}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
+Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
 $\boldsymbol{v}_{b}+\boldsymbol{q}=\mu_{a}\boldsymbol{v}_{a}\qquad\text{and}\qquad\boldsymbol{v}_{a}+\boldsymbol{q}=\mu_{b}\boldsymbol{v}_{b}.$ (8)
@@
 $(\mu_{a}+1)\boldsymbol{v}_{a}=(\mu_{b}+1)\boldsymbol{v}_{b}.$ (9)
-Case 1: $\mu_{a}=\mu_{b}=-1$. In this case, we have
-$\boldsymbol{q}=\boldsymbol{v}_{a}+\boldsymbol{v}_{b}$
-Case 2: $\mu_{a}\neq-1$ and $\mu_{b}\neq-1$. In this case, from (13), it must hold that
+Case 1: $\mu_{a}\neq-1$ and $\mu_{b}\neq-1$. In this case, from (9), it must hold that
 $\boldsymbol{v}_{a}\parallel\boldsymbol{v}_{b}.$
-This leads to two subcases: (1) If $\mu_{a}=\mu_{b}\neq-1$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) If $\mu_{a}+1=-(\mu_{b}+1)$, then equivalently, we have $\mu_{a}+\mu_{b}=-2$.
-This implies $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$.
-∎
+This leads to two subcases:
+(1.1) If $\mu_{a}=\mu_{b}\neq-1$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$;
+(1.2) If $\mu_{a}+1=-(\mu_{b}+1)$, then equivalently, we have $\mu_{a}+\mu_{b}=-2$. This implies $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$.
+Case 2: $\mu_{a}=\mu_{b}=-1$. In this case, we have
+$\boldsymbol{q}+\boldsymbol{v}_{a}+\boldsymbol{v}_{b}=\boldsymbol{0}.$
 $\mu_{a}$$\mu_{b}$
 Figure 7.
-The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; and (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
+The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
 Corollary 5.2.
@@
 Proof.
-With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
-Definition 5.3 (Geometric twins).
-At state $\boldsymbol{\theta}$ of Kuramoto model on $G$, nodes $i$ and $j$ are called geometric twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the compatibility relations of Lemma 5.1.
+With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1.1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
+We introduce a geometric analogue of the notion of twins in graph theory, tailored to the Kuramoto model.
+Definition 5.3 (Geometrically closed twins).
+Motivated by the notion of closed twins in graph theory, we say that, at a state $\boldsymbol{\theta}$ of the Kuramoto model on $G$, nodes $i$ and $j$ are called geometrically closed twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the relations of Lemma 5.1.
+Definition 5.4 (Geometrically stable closed twins).
+We say that, at a state $\boldsymbol{\theta}$ of the Kuramoto model on $G$, nodes $i$ and $j$ are called geometrically stable closed twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the relations of Lemma 5.1 and, in addition, $\mu_{i},\mu_{j}\geq 0$.
 5.2. Phasor Geometry of Closed Twins
@@
 Figure 8. The two black nodes form a pair of closed twins.
-Open twins, defined by $N(i)=N(j)$, will not be needed in this work.
+Open twins, corresponding to $N(i)=N(j)$, will not be needed in this work.
 Corollary 5.6.
-Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$. If nodes $a$ and $b$ form a pair of closed twins, then $\theta_{a}=\theta_{b}$.
+Let $\boldsymbol{\theta}\in\mathbb{R}^{n}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$.
+If nodes $a$ and $b$ form a pair of closed twins (satisfying (10)), then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$.
 Proof.
@@
 $\begin{cases}\displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{b}=\mu_{a}\boldsymbol{v}_{a},\\[4.0pt] \displaystyle\sum_{j\in M}\boldsymbol{v}_{j}+\boldsymbol{v}_{a}=\mu_{b}\boldsymbol{v}_{b}.\end{cases}$
-By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$, and therefore $\theta_{a}=\theta_{b}$. ∎
-Clearly, this is a case where geometric twins defined in (5.3) formed. Moreover, Corollary 5.5 implies immediately that windmill graphs are globally synchronizing.
+By Corollary 5.2 and the fact that $\mu_{a},\mu_{b}\geq 0$, we conclude that $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
+Clearly, this is a case where geometric twins defined in (5.4) formed.
+Moreover, Corollary 5.6 implies immediately that windmill graphs are globally synchronizing.
 Theorem 5.7.
@@
 Proof.
-Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$, and let $G$ be a windmill graph.
-Within each clique of $G$, all non-central nodes have identical closed neighborhoods and therefore form closed twins.
-By Corollary 5.5, all nodes in each clique must share the same phase.
-Contracting each clique into a single node yields a star graph, which is globally synchronizing.
-Hence all nodes of $G$ have the same phase.
+Let $G$ be a windmill graph $W_{p,m}$ with central vertex $c$. Then there exists a partition of the vertex set
+$V\setminus\{c\}\;=\;\biguplus_{\ell=1}^{m}S_{\ell}$
+such that each induced subgraph $G[S_{\ell}]$ is a clique, and every vertex in $S_{\ell}$ is adjacent to $c$.
+For each $\ell$, all vertices in $S_{\ell}$ have identical closed neighborhoods and therefore form a class of closed twins.
+By Corollary 5.6, all vertices in each $S_{\ell}$ must share the same phase.
+According to Lemma (4.4), we have for $\forall i\in S_{l}$, that
+$\boldsymbol{v}_{i}\Uparrow\boldsymbol{v}_{c}+\sum_{j\in S_{l},j\neq i}\boldsymbol{v}_{j}\qquad\text{or}\qquad\boldsymbol{v}_{c}+\sum_{j\in S_{l},j\neq i}\boldsymbol{v}_{j}=\boldsymbol{0}.$
+We distinguish two cases.
+Case 1.
+$\boldsymbol{v}_{c}\Uparrow\boldsymbol{v}_{i}+\sum_{j\in S_{\ell},\,j\neq i}\boldsymbol{v}_{j}$.
+It is then clear that either $\boldsymbol{v}_{c}=\boldsymbol{v}_{i}$ or $\boldsymbol{v}_{c}=-\boldsymbol{v}_{i}$ holds.
+If $\boldsymbol{v}_{c}=-\boldsymbol{v}_{i}$, then $\boldsymbol{\theta}$ is not a second-order stationary point.
+Indeed, let $\boldsymbol{x}\in\mathbb{R}^{n}$ be the vector whose $i$-th component equals $1$ if $i\in S_{l}$ and 0 otherwise.
+Then
+$\boldsymbol{x}^{\top}\boldsymbol{H}\boldsymbol{x}$ $=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
+$=\sum_{i\in S_{l}}\sum_{j\in V\setminus S_{l}}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})$
+$=\sum_{i\in S_{l}}\cos(\theta_{i}-\theta_{c})<0,$
+which contradicts the positive semidefiniteness of $\boldsymbol{H}$.
+Case 2.
+$\boldsymbol{v}_{c}+\sum_{j\in S_{\ell},\,j\neq i}\boldsymbol{v}_{j}=\boldsymbol{0}$.
+Then necessarily $|S_{l}|=2$.
+In this case, the phasors of the two nodes in $S_{l}$ are opposite to $\boldsymbol{v}_{c}$.
+By an argument analogous to the one above, the matrix $\boldsymbol{H}$ is not positive semidefinite.
 ∎
@@
 5.3. Synchronous Pendant Extension
-In this section, we present a scenario in which a group of nodes satisfy the phasor relations of Lemma 5.1 without being closed twins.
-More explicitly, the main result of this section is Lemma 5.8, which shows that even when a set of nodes are not twins, they may still be forced to synchronize at any second-order stationary point, provided that their distinct neighbors form a synchronous pendant whose combined phasor contribution collapses to a single, non-adversarial direction.
-The next lemma, Lemma 5.7, serves as a preparatory step toward establishing Lemma 5.8.
+In this section, we present Lemma 5.9 which describes a scenario where a group of nodes forms geometrically stable closed twins 5.4 without being closed twins in the graph-theoretic sense.
+Before that, we establish Lemma 5.8, which serves as a preparatory step.
 Lemma 5.8.
-Let $G=(V,E)$ be a graph, and let $W\subseteq V$ be partitioned into
+Let $G=(V,E)$ be a graph, and let $W\subseteq V$ admit a partition
 $W=Q\uplus S\uplus P,$
-subject to the following structural conditions:
-  1. (i)
-    Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
-  2. (ii)
-    $S$ induces a clique in $G$;
-  3. (iii)
-    For each $i\in S$, its neighborhood is $P\cup N_{Q}(i)\cup(S\setminus\{i\})$, where $N_{Q}(i)=\{\,q\in Q:q\sim i\,\}$.
-Assume further that the configuration $(\boldsymbol{v}_{i})_{i\in V}$ is a second-order stationary point of the Kuramoto energy (4), and that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$.
+with the following structural conditions:
+(1) Every node in $Q$ has all its neighbors inside $S$, i.e., $N(i)\subseteq S$ for all $i\in Q$;
+(2) The vertex set $S\subseteq V(G)$ induces a clique in $G$;
+(3) For each $i\in S$, its neighborhood is $P\uplus N_{Q}(i)\uplus(S\setminus\{i\})$, where
+$N_{Q}(i)=\{\,q\in Q:q\sim i\,\}.$
+Assume further that $\{\boldsymbol{v}_{i}\}_{i\in V}$ is the vector configuration corresponding to any second-order stationary point $\boldsymbol{\theta}$ of the Kuramoto energy (2), and that the induced subgraph on $Q\uplus S$ is synchronized at $\boldsymbol{\theta}$, namely $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in Q\uplus S$, for some common phasor $\boldsymbol{v}$.
 Then the phasor sum over $P$ satisfies
@@
 For any node $i\in S$,
-$N(i)=Q\uplus(S\setminus\{i\})\uplus P$ (15)
+$N(i)=N_{Q}(i)\uplus(S\setminus\{i\})\uplus P$ (11)
 and
-$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$ (16)
+$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$ (12)
 Plug (11) into (12), we obtain
-$\sum_{j\in Q\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
+$\sum_{j\in N_{Q}(i)\uplus(S\setminus\{i\})\uplus P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0$
 Since $Q\uplus S$ forms a synchronized group, we have
-$\bigl(|Q|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
+$\bigl(|N_{Q}(i)|+|S|-1\bigr)\,\boldsymbol{v}_{i}+\sum_{j\in P}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$ (13)
 Thus
-$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|Q|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
-We now show that $\mu_{i}-(|Q|+|S|-1)<0$ is impossible, because in this case there would exist a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$.
-To see this, let $\boldsymbol{v}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $v_{i}=1$ for $i\in Q\uplus S$ and $v_{i}=0$ otherwise. Then
-$\boldsymbol{v}^{T}H\boldsymbol{v}$ $=\sum_{i\in Q\uplus S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$ (17)
+$\sum_{j\in P}\boldsymbol{v}_{j}=\bigl(\mu_{i}-(|N_{Q}(i)|+|S|-1)\bigr)\boldsymbol{v}_{i},\qquad\mu_{i}\geq 0.$
+In fact, $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$ does not hold, because in this case there exists a direction along which the energy decreases in a neighborhood of $\boldsymbol{\theta}$.
+To see this, let $\boldsymbol{x}$ be the $n$-dimensional indicator vector of $Q\uplus S$, i.e., $\boldsymbol{x}_{i}=1$ for $i\in Q\uplus S$ and $\boldsymbol{x}_{i}=0$ otherwise.
+Since $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$, we have $\langle\sum_{j\in P}\boldsymbol{v}_{j},\boldsymbol{v}_{i}\rangle<0$.
+Thus
+$\boldsymbol{x}^{T}H\boldsymbol{x}$ $=\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$ (14)
+$=\sum_{i\in Q\uplus S}\sum_{j\in V\setminus(Q\uplus S)}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
+$=\sum_{i\in S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})(x_{i}-x_{j})^{2}$
 $=\sum_{i\in S}\sum_{j\in P}\cos(\theta_{i}-\theta_{j})$
@@
 $=|S|\,\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}.$
-where the second equality uses the fact that there are no edges between $Q$ and $P$.
-Clearly, if $\mu_{i}-(|Q|+|S|-1)<0$, then the above quantity is negative, which contradicts the second-order stationarity of $\boldsymbol{\theta}$.
-∎
+Clearly, if $\mu_{i}-(|N_{Q}(i)|+|S|-1)<0$, then (13) implies that
+$\boldsymbol{v}_{i}\cdot\sum_{j\in P}\boldsymbol{v}_{j}<0.$
+Consequently, the quantity (14) is negative, which contradicts the second-order stationarity of the state $\boldsymbol{\theta}$. ∎
 Figure 10.
@@
 The large circular node represents the set $P$ of vertices.
-Gray edges indicate synchronization interactions between the corresponding nodes.
+Gray-shaded edges between two nodes indicate that the vectors corresponding to these two nodes are synchronized.
+The structural relations among the sets $Q$, $S$, and $P$, together with the synchronization of $Q\uplus S$, imply a synchronization in phase between $Q\uplus S$ and the phasor sum over $P$.
 Lemma 5.9.
-Let $G=(V,E)$ be a graph, and suppose there exists a subset of nodes $W\subseteq V$ partitioned into four disjoint subsets:
-$W=Q\uplus S_{1}\uplus S_{2}\uplus P.$
-under the following structural assumptions:
-(i) For all $i\in Q$, we have $N(i)\subseteq S_{1}$;
-(ii) The subgraph $G[S_{1}\uplus S_{2}]$ is a complete graph;
-(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$.
-Formally, if we denote $Q_{i}=N(i)\cap Q$, then
-$N(i)\;=\;P\uplus Q_{i}\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
-Equivalently,
-$N(i)\setminus\bigl(Q\uplus(S_{1}\uplus S_{2})\bigr)\;=\;P.$
-Assume further that all nodes in $Q\uplus S_{1}$ are synchronized, i.e.,
-$\boldsymbol{v}_{i}=\boldsymbol{v}_{j},\quad\forall i,j\in Q\uplus S_{1}.$
-Then the set $S_{1}\uplus S_{2}$ synchronize.
+Let $G=(V,E)$ be a graph, and let $W\subseteq V$ admit a partition
+$W=Q\uplus S_{1}\uplus S_{2}\uplus P$
+with the following structural conditions:
+(i) Every node in $Q$ has all its neighbors inside $S_{1}$, i.e., $N(i)\subseteq S_{1}$ for all $i\in Q$;
+(ii) The vertex set $S_{1}\uplus S_{2}\subseteq V(G)$ induces a clique in $G$;
+For each $i\in S_{1}\uplus S_{2}$, its neighborhood is $P\uplus N_{Q}(i)\uplus(S\setminus\{i\})$, where
+$N_{Q}(i)=\{\,q\in Q:q\sim i\,\}.$
+(iii) For any $i\in S_{1}\uplus S_{2}$, its neighbourhood can be decomposed into three parts: the set $P$, its neighbours inside $Q$, and the rest of the clique $S_{1}\uplus S_{2}$. That is
+$N(i)\;=\;P\uplus N_{Q}(i)\uplus\bigl((S_{1}\uplus S_{2})\setminus\{i\}\bigr).$
+Assume further that $\{\boldsymbol{v}_{i}\}_{i\in V}$ corresponds to a second-order stationary point $\boldsymbol{\theta}$ of the Kuramoto energy (2), at which the induced subgraph on $Q\uplus S_{1}$ is synchronized, i.e.,
+$\boldsymbol{v}_{i}=\boldsymbol{v},\quad\forall i\in Q\uplus S_{1},$
+for some common phasor $\boldsymbol{v}$. Then the set $S_{1}\uplus S_{2}$ synchronize.
 Proof.
-Observe that the set $S_{1}$ induces a fully-connected subgraph, and $Q$ is a synchronous pendant to $S_{1}$, then according to Lemma 5.8, we have
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}.$
-Combine this with the fact
-$\boldsymbol{v}_{i}=\boldsymbol{v}_{j}\quad\forall i\in S_{1},j\in S_{2},$
-we obtain
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{j}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{j}\quad\text{or}\quad\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}+\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$
-Equivalently, we can write
-$\forall i\in S_{1},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (18)
-Another observation is on nodes in $S_{2}$ that
-$\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{2}\setminus\{i\}}\boldsymbol{v}_{k}+\sum_{j\in S_{1}\uplus P}\boldsymbol{v}_{k}=\boldsymbol{0}$
-since every node in $S_{2}$ attains stable equilibrium in $\boldsymbol{\theta}$. Equivalently, we can write
+Step A. Observe that the set $S_{1}$ induces a clique, and $Q\uplus S_{1}$ forms a synchronizing group. Hence by Lemma 5.8, for every $i\in S_{1}$ there exists $\mu_{i}\geq 0$ such that
+$\sum_{k\in S_{2}\uplus P}\boldsymbol{v}_{k}=\mu_{i}\,\boldsymbol{v}_{i}.$ (15)
+Step B. Since $S_{1}$ is a synchronizing group, we have
+$\sum_{k\in S_{1}\setminus\{i\}}\boldsymbol{v}_{k}=(|S_{1}|-1)\boldsymbol{v}_{i}.$
+Adding this term to (15) yields
+$\sum_{k\in S_{2}\uplus P\uplus(S_{1}\setminus\{i\})}\boldsymbol{v}_{k}=\bigl(\mu_{i}+|S_{1}|-1\bigr)\boldsymbol{v}_{i}.$
+where $\mu_{i}+|S_{1}|-1>0$. In other words,
+$\forall i\in S_{1},\qquad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{2}\uplus P\uplus(S_{1}\setminus\{i\})}\boldsymbol{v}_{k}.$ (16)
+Step C. According to Lemma 4.4, nodes in the set $S_{2}$ satisfy
 $\forall i\in S_{2},\quad\boldsymbol{v}_{i}\Uparrow\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}\quad\text{or}\quad\sum_{k\in S_{1}\uplus P\uplus S_{2}\setminus\{i\}}\boldsymbol{v}_{k}=\boldsymbol{0}.$ (17)
-Combined the facts (18) and (19), we get nodes in $S_{1}\uplus S_{2}$ are geometric stable closed twins. ∎
-Clearly, this is another situation in which geometric twins, as defined in Definition 5.3, appear in addition to closed twins at the second-order stationary point.
+since $\boldsymbol{\theta}$ is a second-order stationary point.
+Combining (16) and (17), we deduce that all nodes in $S_{1}\uplus S_{2}$ are geometrically stable twins in the sense of Definition 5.4.
+It then follows from Corollary 5.2 that $S_{1}\uplus S_{2}$ forms a synchronizing group.
+∎
 Figure 11.
-Illustration of Lemma 5.8, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
+Illustration of the proof of Lemma 5.9, illustrating how the synchronized block $Q\uplus S_{1}$ extends to include $S_{2}$ through propagation enabled by the specific local structure.
 The green vertices correspond to nodes for which $\mu_{i}\boldsymbol{v}_{i}=\sum_{j\in N(i)}\boldsymbol{v}_{j}$ holds with $\mu_{i}\geq 0$.
@@
 Each orange arc separates a small green node from a group of nodes it connects to, indicating that the vector of the green node is in the same direction as the sum of the vectors of the nodes on the opposite side of the arc.
-5.4. Twin Attachment
-If a set of nodes all have identical neighborhoods, they form a group of closed twins.
-When an additional node is connected exclusively to this group, the entire collection of nodes, must take the same phase at any second-order stationary point.
-Lemma 5.9.
-Let $\boldsymbol{\theta}$ be a second-order stationary point of $E_{G}(\boldsymbol{\theta})$.
-If $S\subseteq V$ is a set of closed twins in $G$ and $A\in V$ satisfies $N(A)=S$, then $\theta_{i}=\theta_{j}$ for all $i,j\in S\cup\{A\}$.
-Proof.
-Since $\boldsymbol{\theta}$ is a second-order stationary point of $E_{G}(\boldsymbol{\theta})$ and $S$ forms a set of closed twins, Lemma 5.5 yields
-$\theta_{i}=\theta_{j}\qquad\text{for all }i,j\in S.$
-Because $N(A)=S$, Lemma 4.4 implies
-$\theta_{A}=\theta_{i}\qquad\text{for all }i\in S.$
-Hence all nodes in $S\cup\{A\}$ synchronize. ∎
-Figure 13.
-Illustration of Lemma 5.9.
-All green nodes attain stable equilibrium.
-The set $S$ (middle row) forms a set of false twins, and node AA (top) has neighborhood exactly $S$.
-Thin lines represent the edges in the original graph.
-The conclusion is that all green nodes must synchronize, indicated by the thick edges in the right diagram.
 6. Main Proof: the Induction Argument
@@
 Let us take a look at the threshold graph corresponding to the sequence 01010101.
-The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence, with each bit determining whether the added vertex is isolated (0) or dominating (1).
+The first vertex is denoted by $A$, and we label the vertices $A$ through $I$ in the order in which they appear in the sequence.
 Let $G$ be the threshold graph corresponding to the sequence 01010101.
 To prove that $G$ is globally synchronizing, we show that every second-order stationary point $\boldsymbol{\theta}$ of $E_{G}(\boldsymbol{\theta})$ is synchronous; that is, $\theta_{1}=\cdots=\theta_{n}$.
-The following figure illustrates how synchronization must propagate, step by step, in any second-order stationary configuration on this graph.
 ACEGIHFDB
@@
 ACEGIHFDB
-Figure 14. Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous.
+Figure 12.
+Synchronization propagates in eight steps, forcing every second-order stationary point to be synchronous.
+Gray-shaded edges represent synchronizing relations that have been established at each step.
 Step (1) We begin with the edge between $H$ and $I$.
 Applying Lemma 4.4 to node $H$ shows that, since $\boldsymbol{\theta}$ is a second-order stationary point, we must have $\boldsymbol{v}_{H}\Uparrow\boldsymbol{v}_{I}$.
-Equivalently, nodes $H$ and $I$ synchronize.
-Step (2) Observe that $N(G)=N(I)\cup H$ and $\theta_{H}=\theta_{I}$, applying Lemma by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain $\theta_{G}=\theta_{I}$.
-Intuitively, the only difference between the neighborhoods of $G$ and $H$ is that $H$ has one additional neighbor, namely $I$.
+Equivalently,
+$\theta_{H}=\theta_{I}.$
+Step (2) Observe that $N(G)=N(I)\cup\{H\}$ and $\theta_{H}=\theta_{I}$, applying Lemma 5.9 by letting $Q=\{H\}$, $S_{1}=\{I\}$, $S_{2}=\{G\}$, $P=\{F,D,B,A,C,E\}$, we obtain
+$\theta_{G}=\theta_{I}.$
+Intuitively, the only difference between the neighborhoods of $G$ and $I$ is that $I$ has one additional neighbor, namely $H$.
 This extra neighbor does not obstruct the synchronization between $G$ and $I$, because $H$ is already synchronized with $I$ and therefore behaves as a synchronous pendant at this stage.
-Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2).
-Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{F}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Step (4) Observe that nodes $F$ and $H$ connect only to $G$ and $I$, and the four-node set $\{F,H,G,I\}$ behaves as a synchronizing group.
-Furthermore, ignoring the leaf-like nodes $F$ and $H$ and edges in the induced clique on $\{E,G,I\}$, the two nodes $G$ and $I$ share exactly the same neighbors:
-$N(G)\setminus(\{F\}\cup\{E,I\})\;=\;N(I)\setminus(\{F,H\}\cup\{E,G\})\;=\;N(E)\setminus\{G,I\}\;=\;\{D,B,A,C\}.$
-Therefore, by applying Lemma 5.8 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude that the nodes $E$, $G$, and $I$ form a synchronizing group.
-Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4).
-Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{D}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Step (6) Observe that nodes $D$, $F$, and $H$ connect only to $E$, $G$, and $I$, and therefore the six-node set $\{D,F,H,E,G,I\}$ behaves as a synchronizing group.
-Furthermore, ignoring the leaf-like nodes $D$, $F$, and $H$ and edges in the induced clique on $\{C,E,G,I\}$, the four nodes $C$, $E$, $G$, and $I$ share exactly the same neighbors, namely $\{A,B\}$.
-Therefore, by applying Lemma 5.8 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude that the nodes $C$, $E$, $G$, and $I$ form a synchronizing group.
-Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6).
-Therefore, by Lemma 5.9, we have $\boldsymbol{v}_{B}=\boldsymbol{v}_{C}=\boldsymbol{v}_{E}=\boldsymbol{v}_{G}=\boldsymbol{v}_{I}$.
-Step (8) As in Step (7), Lemma 5.9 implies that $A$ synchronizes with $C$, $E$, $G$, and $I$, since these nodes form a synchronizing group and constitute all neighbors of $A$.
-In conclusion, $\boldsymbol{\theta}$ must assign the same angle to every node.
-Equivalently, every second-order stationary point is synchronous, and hence the threshold graph is globally synchronizing.
+Step (3) Note that node $F$ connects only to $G$ and $I$, which are synchronized as established in Step (2). Therefore, by Lemma 4.4, we have
+$\theta_{F}=\theta_{G}=\theta_{I}.$
+Step (4) Applying Lemma 5.9 with $Q=\{F,H\}$, $S_{1}=\{G,I\}$, $S_{2}=\{E\}$, and $P=\{D,B,A,C\}$, we conclude
+$\theta_{E}=\theta_{G}=\theta_{I}.$
+Step (5) Note that node $D$ connects to $E$, $G$, and $I$, which are synchronized as established in Step (4). Therefore, by Lemma 4.4, we have
+$\theta_{D}=\theta_{E}=\theta_{G}=\theta_{I}.$
+Step (6) Applying Lemma 5.9 with $Q=\{D,F,H\}$, $S_{1}=\{E,G,I\}$, $S_{2}=\{C\}$, and $P=\{A,B\}$, we conclude
+$\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}.$
+Step (7) Note that node $B$ connects to $C$, $E$, $G$, and $I$, which are synchronized as established in Step (6). Therefore, by Lemma 4.4, we have
+$\theta_{B}=\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}.$
+Step (8) As in Step (7), Lemma 4.4 implies that
+$\theta_{A}=\theta_{C}=\theta_{E}=\theta_{G}=\theta_{I}$
+since these nodes form a synchronizing group and constitute all neighbors of $A$.
+In conclusion, every second-order stationary point $\boldsymbol{\theta}$ is a fully synchronized state. Hence the threshold graph corresponding to 01010101 is globally synchronizing.
 6.2. Proof of Theorem 1.2: General Case
 Proof of Theorem 1.2.
-Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, as introduced in Definition 2.1.
-Since the graph is connected, the last bit must be a $1$.
-We decompose the sequence into maximal contiguous blocks of $1$’s and 0’s.
-Let $U_{1},U_{2},\dots,U_{k}$ denote the blocks of $1$’s (indexed from bottom to top), and let $I_{i}$ denote the 0-block that lies immediately above $U_{i}$, if such a block exists.
-In this notation, the sequence has one of the two forms
+Any connected threshold graph can be described by a bit sequence of $1$’s and 0’s, where the last bit must be a $1$; otherwise, the threshold graph is not connected.
+For simplicity, we decompose the sequence into maximal contiguous blocks of $1$’s and 0’s:
 $U_{1},\,I_{2},\,U_{2},\,\dots,\,I_{k-1},\,U_{k-1},\,I_{k},\,U_{k},$
@@
 $I_{1},\,U_{1},\,I_{2},\,U_{2},\,\dots,\,I_{k-1},\,U_{k-1},\,I_{k},\,U_{k},$
-depending on whether or not it begins with a block of 0’s. We focus on the first case; the second follows by an entirely analogous argument.
-Given that $\boldsymbol{\theta}$ is a second-order stationary point of $E(\boldsymbol{\theta})$, we aim to show that $\theta_{1}=\cdots=\theta_{n}$, and we proceed by induction.
-We proceed by induction.
+where $U_{i}$ denotes a subsequence of $1$’s and $I_{i}$ denotes a subsequence of 0’s.
+Without loss of generality, we focus on threshold graphs $G=(V,E)$ of the first sequence type; the second case follows by an entirely analogous argument.
+Let $\boldsymbol{\theta}$ be a second-order stationary point of the energy function $E_{G}(\boldsymbol{\theta})$ of the homogeneous Kuramoto model on $G$.
+We prove that $\boldsymbol{\theta}$ is a fully synchronized state by induction on $m$, where $m$ denotes the number of $1$-blocks counted from the end of the threshold sequence.
 For the base case, observe that all nodes within a block of $1$’s form a set of closed twins.
-Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Lemma 5.1.
-Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\dots,U_{k}$, are synchronized.
-Lemma 5.9 then implies that the adjacent 0-blocks $I_{k-m+1},\dots,I_{k}$ synchronize with them as well.
-We next show that the preceding $1$-block $U_{k-m}$ also synchronizes with $U_{k-m+1},\dots,U_{k}$.
-Indeed, define
+Hence, at any second-order stationary point $\boldsymbol{\theta}$, the nodes in the last block $U_{k}$ must synchronize, as follows directly from Corollary 5.6.
+Assume as the inductive hypothesis that the nodes in the last $m$ blocks of $1$’s, $U_{k-m+1},\ldots,U_{k}$, are synchronized, i.e., there exists a common phasor $\boldsymbol{v}$ such that $\boldsymbol{v}_{i}=\boldsymbol{v}$ for all $i\in U_{k-m+1}\uplus\cdots\uplus U_{k}$.
+Lemma 4.4 then implies that the nodes in the adjacent 0-blocks $I_{k-m+1}\uplus\cdots\uplus I_{k}$ also satisfy $\boldsymbol{v}_{i}=\boldsymbol{v}$.
+Define
 $Q$ $=I_{k-m+1}\uplus\cdots\uplus I_{k},$
@@
 $P$ $=U_{1}\uplus\cdots\uplus U_{k-m-1}\uplus I_{2}\uplus\cdots\uplus I_{k-m}.$
-We next verify that the sets $Q,S_{1},S_{2}$, and $P$ satisfy the assumptions of Lemma 5.8.
-First, for every node $i$ in $Q$, its neighborhood is contained in $S_{1}$; indeed, these nodes are isolated at this stage and connect only to the dominating nodes that are added later.
-Second, the sets $S_{1}$ and $S_{2}$ together induce a clique.
-Third, for any node in $S_{1}\cup S_{2}$, its neighbors consist exactly of all other nodes in $S_{1}\cup S_{2}$, all nodes in $P$, and possibly some nodes in $Q$.
-Hence all conditions of Lemma 5.8 are satisfied, and we conclude that $S_{1}$ and $S_{2}$ must synchronize.
-In particular, the block $U_{k-m}$ synchronizes with $U_{k-m+1}\uplus U_{k-m+2}\uplus\cdots\uplus U_{k}$.
-By induction, all nodes synchronize at $\boldsymbol{\theta}$.
-Therefore, every second-order stationary point is a synchronous state.
-It follows that every connected threshold graph is globally synchronizing.
-∎
-7. Discussion: Beyond Density-Based Guarantees for Global Synchronization
-As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction ($\geq 75\%$) of the remaining vertices.
+We now verify that the assumptions of Lemma 5.9 are satisfied by the above decomposition, which follows directly from the threshold construction.
+By construction, the vertices in $Q$ correspond to the most recent 0-blocks preceding $S_{1}$.
+Hence, every vertex in $Q$ is adjacent only to vertices in $S_{1}$.
+Moreover, since $S_{1}$ and $S_{2}$ consist of consecutive dominating blocks, the induced subgraph on $S_{1}\cup S_{2}$ is a clique.
+Finally, for each $i\in S_{1}$, its neighborhood can be decomposed as
+$N(i)=P\;\uplus\;\bigl((S_{1}\cup S_{2})\setminus\{i\}\bigr)\;\uplus\;N_{Q}(i),$
+where $P$ denotes the set of vertices added earlier in the construction.
+Hence, all the assumptions of Lemma 5.9 are satisfied, and we conclude that the sets $S_{1}$ and $S_{2}$ synchronize.
+In other words, the last $m+1$ blocks of $1$’s $U_{k-m},U_{k-m+1},\cdots,U_{k}$ synchronize.
+Therefore, every second-order stationary point is a synchronous state, and it follows that $G$ is globally synchronizing. ∎
+7. Discussion: Beyond Density-Based Guarantees
+As noted in the introduction, previous guarantees for global synchronization have largely focused on minimum-degree conditions: that is, requiring every node to be connected to more than a fixed fraction of the remaining vertices.
 In this regime, graphs are necessarily highly dense, and their degree sequences are relatively homogeneous.
 This naturally raises the question:
-Does a graph that exhibits global synchronization necessarily have high edge density or a homogeneous degree sequence (i.e., with relatively small variation in node degrees)?
+Does a graph that exhibits global synchronization necessarily have high edge density or a nearly regular degree sequence (i.e., with small degree variability)?
 In this section, we provide a negative answer to this question [^4] Of course, trees provide trivial examples.
-by exploiting two structural properties of threshold graphs, which were shown in the previous sections to be globally synchronizing.
-More precisely, we show that threshold graphs realize every admissible edge density among connected graphs.
-Moreover, their degree sequences are highly heterogeneous: among all connected graphs with a fixed edge density, threshold graphs attain the maximally left-skewed degree sequence when ordered in nonincreasing order.
+by showing that threshold graphs realize every admissible edge density, and that their degree sequences are extremal in the sense that they majorize any other degree sequence.
 7.1. Threshold Graphs Realize All Densities
-Note that the edge density of a graph $G$ on $n$ vertices takes values in
+The edge density of a graph $G$ on $n$ vertices takes values in
 $\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,\binom{n}{2}\}\,\right\}$. (18)
-and for connected graphs
-$\mathrm{density}(G)\in\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{n-1,\ldots,\binom{n}{2}\}\,\right\}$. (21)
-If the graph is not connected, to realize any density in $\left\{\,\frac{m}{\binom{n}{2}}\;\middle|\;m\in\{0,\ldots,n-2\}\,\right\}$, construct a threshold graph on $n$ vertices whose n1n-1-length sequence contains $1$ at the $m$-th position.
-001
-Figure 15.
-Illustration of the threshold graph construction.
-A single 11 at the third position of the generating sequence adds a dominating vertex, which contributes three incident edges shown as curved arcs.
-The following lemma shows that every admissible edge density in the range (21) is realized by a connected threshold graph on $n$ vertices.
+Moreover, for connected graphs, the edge density takes values corresponding to
+$m\in\{n-1,\ldots,\binom{n}{2}\}$.
+The following lemma shows that every admissible edge density can be realized by a threshold graph.
 Lemma 7.1.
-Fix $n\geq 2$. For any integer $m\in[\,n-1,\binom{n}{2}\,$], there exists a connected threshold graph on $n$ vertices with exactly $m$ edges.
+For any pair of integers $n>0$ and $0\leq m\leq\binom{n}{2}$, there is a threshold graph with $n$ vertices and $m$ edges.
+If $m\geq n-1$, there is a connected threshold graph with $n$ vertices and $m$ edges.
 Proof.
-Since the graph is connected, the last vertex added in the construction of a threshold graph must be a dominating vertex, contributing n1n-1 edges.
-To obtain a graph with exactly $m$ edges, the vertices added before the last one must therefore contribute a total of m(n1)m-(n-1) edges.
-Since $m\in\{n-1,\ldots,\binom{n}{2}\}$, it follows that
-$m-(n-1)\in\left\{0,1,\ldots,\binom{n}{2}-(n-1)\right\}=\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
-Edges are created only when dominating vertices are added during the construction.
-We fix the following convention: starting from an initial vertex, we index the construction sequence by counting only the dominating vertices added thereafter.
-Under this convention, a dominating vertex added at position $p$ contributes exactly $p$ new edges, as it connects to all previously existing vertices.
-Hence the possible edge contributions of dominating vertices, except for the last one added in the construction sequence, take values in $\{1,2,\ldots,n-2\}$.
-Therefore, the problem reduces to determining at which positions the dominating vertices except for the last one are added so that they collectively contribute m(n1)m-(n-1) edges.
-More precisely, it suffices to prove the following combinatorial statement:
-For any integer
-$k\in\left\{0,1,\ldots,\frac{(n-2)(n-1)}{2}\right\},$
-there exists a subset
-$A\subseteq\{1,2,\ldots,n-2\}\quad\textit{such that}\quad\sum_{a\in A}a=k.$
-We prove this claim by induction on $n$.
-Base case. For $n=3$, we have $\{1,2,\ldots,n-2\}=\{1\}$ and $k\in\{0,1\}$. If $k=0$, take $A=\emptyset$; if $k=1$, take $A=\{1\}$.
-Induction hypothesis. Assume that for any
-$k\in\left\{0,1,2,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
-there exists a subset
-$A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime}_{k}}a=k.$
-Induction step. We now consider any
-$k\in\left\{0,1,2,\ldots,\frac{(n-2)(n-1)}{2}\right\}.$
-Case 1. If
-$k\leq\frac{(n-3)(n-2)}{2},$
-then by the induction hypothesis there exists $A^{\prime}_{k}\subseteq\{1,2,\ldots,n-3\}$ such that $\sum_{a\in A^{\prime}_{k}}a=k$. Setting $A:=A^{\prime}_{k}$ yields the desired subset.
-Case 2. If
-$k>\frac{(n-3)(n-2)}{2},$
-then using only values in $\{1,2,\ldots,n-3\}$ cannot achieve the sum $k$, since the maximal possible sum is
-$\sum_{i=1}^{n-3}i=\frac{(n-2)(n-3)}{2},$
-which is strictly smaller than $k$. Hence the element n2n-2 must be included in $A$.
-It then remains to find a subset
-$A^{\prime\prime}\subseteq\{1,2,\ldots,n-3\}\quad\text{such that}\quad\sum_{a\in A^{\prime\prime}}a=k-(n-2).$
-Since
-$k-(n-2)\in\left\{\frac{(n-2)(n-5)}{2}+1,\ldots,\frac{(n-2)(n-3)}{2}\right\}\subseteq\left\{0,1,\ldots,\frac{(n-3)(n-2)}{2}\right\},$
-the induction hypothesis guarantees the existence of such a subset; in particular, we may take
-$A^{\prime\prime}=A^{\prime}_{\,k-(n-2)}.$
-Setting $A:=A^{\prime\prime}\cup\{n-2\}$ then yields $\sum_{a\in A}a=k$.
-This completes the induction and hence the proof. ∎
+The proof is by induction over $n$. For the base case, note that the only graph with $n=1$ vertex has $0=m=\binom{n}{2}$ edges and is connected by definition. For $n>1$, we consider two cases:
+  • If m<n1m<n-1, the induction hypothesis implies that there is a threshold graph with $n-1\geq 1$ vertices and $m<n-1\leq\binom{n-1}{2}$ edges to which an isolated vertex can be added to obtain a threshold graph with $n$ vertices and $m$ edges.
+  • Otherwise, the induction hypothesis implies that there is a threshold graph with $m-(n-1)\leq\binom{n-1}{2}$ edges to which a universal vertex can be added to obtain a connected threshold graph with $n$ vertices and $m$ edges.
 7.2. Degree Sequence of Threshold Graphs
-One way to compare the “shape” of the degree sequences of two graphs with the same edge density is through majorization, which provides a partial order quantifying how spread out or concentrated the degrees are.
-More specifically, let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
+One way to compare how spread out the degree sequences of two graphs are is through majorization. Let $G$ and $G^{\prime}$ be two graphs with the same edge density, and let
 $\mathbf{d}(G)=(d_{1},\dots,d_{n}),\qquad\mathbf{d}(G^{\prime})=(d^{\prime}_{1},\dots,d^{\prime}_{n})$
@@
 A degree sequence $\mathbf{d}$ is a threshold sequence if and only if it is not strictly majorized by any other graphical degree sequence.
-As shown in the previous section, for any admissible edge density there exists at least one (not necessarily unique) threshold graph realizing that density.
-Moreover, Theorem 7.2 tells that, among all graphs realizing a given edge density, threshold graphs are exactly those whose degree sequences are maximal with respect to majorization.
-In this sense, threshold graphs are always the most left-skewed among graphs with a fixed edge density, since their degree sequences are maximal under majorization, meaning that the degree mass is as concentrated as possible on high-degree vertices.
+This theorem formalizes the intuition that threshold graphs are extremal, in the sense that their degree mass is concentrated as much as possible on high-degree vertices.
 8. Summary and Open Problems
-In this work, we present a class of globally synchronizing graphs — threshold graphs — whose density vary from $2/n$ to $1$.
-Our proof crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the addition of isolated nodes and universal nodes.
-This closure property is essential to our inductive argument, as it allows us to show that global synchronization is preserved when starting from a threshold graph and adding either an isolated node or a universal node.
+In this work, we identify threshold graphs as a class of interaction networks that are globally synchronizing, despite spanning the full range of admissible edge densities from $2/n$ to $1$.
+This significantly extends existing sufficient conditions for global synchronization, which are largely restricted to graphs with high minimum degree or uniformly dense connectivity.
+Our analysis crucially exploits the constructive nature of threshold graphs: they form the smallest graph class closed under the successive addition of isolated and universal nodes.
+This closure property enables an inductive approach, whereby global synchronization is shown to be preserved when starting from a smaller synchronizing threshold graph and adding either an isolated node or a universal node.
+From a geometric perspective, this reveals a local-to-global propagation mechanism, whereby synchronization constraints imposed on a small subset of nodes progressively force synchrony across the entire network.
 A natural question is whether this closure phenomenon extends beyond threshold graphs.